A Bibliography of Publications in Scripta Mathematica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 March 2024 Version 1.07

Title word cross-reference

# [Blo53a]. #161 [Moe47].

1, 2, 3, 4, ⋯ [Uhl53a]. 10^{m} [Uhl53b, Uhl53a]. 16 × 16 [Bra36]. 1^{n} + 2^{n} + 3^{n} + ⋯ + (m − 1)^{n} = m^{n} [Mos53]. 2 [ABG73, Uhl49b]. 2^{n} ± 1 [Kra52]. 2x^{2} + 1 ≡ 0 (mod p^{2}) [Glo53c]. 2x^{2} + 1 ≡ 0 (mod p^{4}) [Glo53c]. 4000 < p < 6000 [Glo55]. 449! [Uhl55]. 57 [FG54]. 751! [Uhl55]. 8 [Hor55]. 9^{9^{9^{9}}} [Sha58]. a [Bee50]. a^{10} + a^{5} + 1 [Jon39]. A^{2} + 2B^{2} = C^{2} + D^{2} [Whi51]. a^{n−1} ≡ (mod n) [Bee50]. χ^{2} [Uhl52b]. E_{6} [Bro73]. G_{2} [Bro73]. M [Sin67]. n [Bee50]. n = 1, 3, 5, 7 [Moe52]. p [Fra73]. π [Gri60]. px^{2} − y^{2} − z^{2} = 0 [Dav66]. q [Gou63]. r [Mir49]. ^{3} \sqrt{2} [Uhl52b]. ^{3} \sqrt{3} [Uhl52b]. ^{3} \sqrt{4} [Uhl52b]. ^{3} \sqrt{9} [Uhl52b]. sin 1^{∘} [Aab54]. \sqrt{2} [Blo55]. u^{2} + v^{2} = w^{2} + 1 [Ven58]. φ [Coh57, Coh58]. x^{2} − kxy + y^{2} = 1 [Bro39]. x^{3} + y^{3} = u^{3} + v^{3} = N [Bri60]. x^{4} + 1 ≡ 0(modp^{3}) [Glo50]. x^{4} + 1 ≡ 0 (mod p^{2}) [Glo55]. x^{8} − 1 ≡ 0 (mod p) [Glo49a]. x_{1}^{n} + x_{2}^{n} + x_{3}^{n} + x_{4}^{n} + x_{5}^{n} = y_{1}^{n} + y_{2}^{n} + y_{3}^{n} + y_{4}^{n} + y_{5}^{n} [Moe52]. y^{2} = ax^{3} + bx^{2} + cx + d [Mor64, Mor65].

-function [Coh57, Coh58]. -rank [ABG73]. -series [Gou63]. -th [Uhl53b]. -valued [Sin67].

10th [Pol32]. 150th [Mal63]. 16th [Uhl53d]. 17th [Uhl53d]. 1850 [Kar54].

A. [Ano73a, Ano73b, Her73, Kar46, Lor32, Lor34]. Abaci [Eis51]. abacus [Loc32b, Ven54a, Ven54b]. Abel [AF59]. Abraham [Dav36, Wal34]. Abrahim [AM59]. Abram [Arc35d]. absolute [AF59]. Abstractions [Lap52, Lap53]. Academy [GA35]. accomptant [Kar35]. according [Cox36, Neu59]. Addition [Hor53, Hor55]. admit [Fin63]. adoption [Sim36]. Adrian [Ano73a, Ano73b]. Advanced [Arc32f]. affine [Coh64, Coh65, Cox55]. after [Str49]. age [Boy54a, Boy54b, Boy55a, Boy51a, Boy51b]. ago [Smi36, Smi38]. al [AM59, Aab54, Ken64]. Al-Kāshı [Aab54]. al-Khayyami [AM59]. Al-Khwārizmī [Ken64]. Albert [Ano73a, Ano73b, Her73]. Alexandrian [Boy54a, Boy54b, Boy55a]. Alfred [Fra38, Ham48a, Ham48b]. Algebra [Mac44, Smi36, Smi38, Ber50, Hes73, McC73a, Oeh73, Sim32a, Sim34d]. Algebraic [Kar44, BF38a, BF38b, HD35a, HD35b, JK73, Sel73]. algebras [Alb63, Bro73, Fra73, JK73, Kap73, McC73b, Nem73, Osb73, Sch73a]. Ali [iAa60]. allied [Van61]. alphabetical [Ira52]. alternative [Lig63]. altitude [Gol64]. altitudes [Cou48]. Ambiguities [Laf47]. America [Kar54, Kar41, Kar47]. American [Sim34d, Dor50, Fle43, Sim32a, Sim35a, Sim36, Smi34, Yal32]. Amicable [Esc46, Gar57, Gar58, Pou48]. Amir [iAa60]. Among [Sim45, AF59]. anachronism [Loc32a]. analog [Car63]. analogue [Gou63]. analysis [Glo53a]. Analyst [Gar32]. Analytic [Boc64, Boy54a, Boy54b, Boy55a, Boy53a, Boy54c, Boy55b, De 48, Lor32, Sie63]. analytical [Low50]. Anania [Vog35]. anatomy [McC38, McC40]. ancient [Bru38a, Bru38b, Bru38c, Loc32b, Neu56, Neu57, Sei59]. angle [Kas45]. angles [KH50, KH51]. anniversary [Mal59, Mal63]. annulus [RW60, Ric70]. Anonymous [Jon47, Cla34]. antecedent [Uhl56]. Apollonian [Kas43]. Apollonius [Bea55, Neu59]. application [KM61, KC36]. appreciation [Bel58]. approximation [RA52]. approximations [Uhl52b, Win57, Win58]. Arabian [Uhl49a]. Arabic [Ira52]. arbelos [Ban55]. arbitrarily [Har57, Har58]. arches [Win57, Win58]. Archimedes [GH60, Sho34a, Sho34b]. area [Thé52a, Thé52b]. areas [Whi43a, Whi43b]. Arithmetic [Kar33, Kar34, Leh73, Sim32c, Van61]. arithmeticae [GA35]. Arithmetical [Coh63, Bel47a, Bra38, Coh60, Del63, Kar70]. arithmetics [Kar35, Sie60]. arithmetization [Gin38a]. arithmomachia [Ric43a, Ric43b, Ric43c]. Art [AC35, Sta38, Loc33, Whi32]. Arthur [Wat39]. articles [Str49]. asides [Cou47]. Aspects [Bli67, Bel56, vdB67]. associated [Fel40, Spe63, Tho73a]. associates [Gil64a, Gil65]. associative [Alb63, Div73, Kle73, Sch73a]. astronomical [Tho55]. astronomy [Fra60, Hea38, Neu56, Neu57, Tho57, Tho58]. asymptotic [Miu73]. aud [Fel40]. Author [Jon47, Bra49]. autograph [Sim45]. averages [Tho73a]. axially [Gil64a, Gil65]. axiomatic [Zlo60]. axioms [Sza64a, Sza64b].

Bīrūnī [Ken59]. Babbage [Arc35b]. Babylonian [Neu34a]. Ball [Cox52b]. Banach [Alt60, Ste63]. based [BF38a, BF38b]. basic [Car63]. Bastille [Sch32a]. Bauer [Car64, Car65]. bearings [Key32a]. beauty [Mor36]. became [De 48]. beginnings [Gol73, Sza64a, Sza64b]. below [Sch73b]. ben [Dav36, Gin32a, Gin32b]. Benjamin [Key34e, Gin34]. Berkeley [Arc35c]. Bernoulli [Car56, Car57, Car61, Goo70, Kel57, Kel58, Van61]. Bernstein [Laf47]. Bertrand [Arc45a]. Berwald [Pin64, Pin65]. between [BB61, Can51, Div73, Gin54a, Uhl55, Wei35]. Bianchini [Tho50a, Tho50b, Tho53, Tho55]. Bibliographia [Arc32a, Arc32b, Arc34a, Arc34b, Arc34c, Arc35a, Arc35b]. Bibliographical [Arc48, Kar46, Arc45b, Arc46a, Arc46b, Kau67]. Bibliographies [Arc32c, Arc32g]. Bibliography [Arc35c, Kar41, Kar54]. Bieberbach [CS60, Gar35, God35]. Bierens [Sch55]. big [Leh33]. bimagic [Leo54]. binary [Dav66]. Binomial [Glo48a, Gin38b]. Bio [Arc48, Arc45b, Arc46a, Arc46b]. Bio-Bibliographical [Arc48, Arc45b, Arc46a, Arc46b]. Biography [Arc36a, Arc36b, Arc36c, Arc36d, Hil34]. birth [Mal63]. Bohr [Boc64]. Boissière [Ric46]. Bolyai [Bel38a, Bel38b]. Boncompagni [Arc32d]. Bonse [Kan56]. Book [Bli67, Ste50, vdB67]. Books [Sel55, Sim34a, Bra38, Kar32, Str49]. Boolean [Ber50]. Borel [Coh73]. both [Kel57, Kel58]. bound [Thoxx]. bounded [Sch73b]. bounds [JT45]. Boyd [WC60]. boyhood [Tal32]. brief [Kar47, Uhl52a]. Brouwerian [Lar40]. Bullettino [Arc32d]. Bureau [Low49]. Busard [Mur64]. Busts [Arc48, Arc45b, Arc46a, Arc46b].

calculation [Loc33]. calculations [Res43]. Calculus [Men56a, Men56b, BM53, CW47]. calendar [Ken64, Res43]. Cambridge [Ham48a, Wat39]. camera [Uhl53c]. Canada [Bra52]. Canonical [RW60, iH73]. Cantor [Arc34h, Hil34, Rus34, Smi32b]. Card [Gar48]. Cardan [Men39]. Carroll [Arc32g]. Cartesian [Boy47, Boy55b]. Cassius [Key47]. Castillon [Cou54a, Cou54b]. catacaustica [Emc52]. Cauchy [Lor32]. Cavalieri [Boy41]. Cayley [Mos57, Rap59, Wat39]. centenarians [Kar33, Kar34]. center [AS48, Ste48]. Central [Gol56]. Centroidal [KJWC34]. centuries [Smi35a, Tho57, Tho58]. century [Dav36, Eis59, Eis60, Gin32h, Ken51, Pol32, Sim34a, Sim35a]. certain [Alb67, Ber57a, Ber58a, Ber61, Kra45, Low50]. Cesàro [Lor57, Lor58]. Cevian [Cou57]. chain [Cou57]. chains [Man61]. Changes [Smi35a, Bli67, vdB67]. chapter [Raj49]. characteristic [Che73, Fra73].

Charles [Arc35b, Eis51, Eis59, Eis60, Key35a, Key41, Neu34b]. check [Kar46]. checkers [Lan54]. Chinese [Bli67, vdB67]. choice [Zlo60]. Christopher [Hux60a, Sim34b]. circle [AS48, Bra47, Hea35c, Kla54, Ste48]. circles [KCW43, Man61]. circular [Mor39, RW60]. clarifying [Emc52]. Clasen [Glo60a]. class [Coh64, Coh65, Del63, PS56, PS57, Wym73, Zei73]. classes [Div73, Gil56, Gil57]. clue [Key32b, Key47]. codes [Tho73a]. coefficient [CS60, SS50]. coefficients [Gin38b, Jar54]. coin [Gro45]. coins [Yal32]. Colin [Boy66]. Collected [Key47]. Collection [Syn45, Sch34, Sim45]. College [Rol19]. color [Fra39a, Fra39b, Rap59]. coloring [Cox57, Cox58, Tut61]. Columbia [Sch34]. Combinatorial [vdB67, Bli67]. comets [Sch58]. Commandino [Ros70]. Common [Gar49]. commutative [Kle73, Sch73a]. compact [Boc64]. Comparison [AM57, AM58a]. Complete [Dre53]. complex [iH73, Lub59]. composite [Bee50]. computation [Gol73, Uhl50a, Low49]. computer [Ken51]. Concentric [Hea35a]. concept [Boy46, MB32a, MB32b]. concepts [Sze60]. concerning [Edr56a, Edr56b, End56a, End56b, Pal53b, Wei55]. conditions [Moe53]. Configuration [Tri55]. conformal [RW60]. Congruence [Car54, Glo49a, Glo50, Glo55]. congruences [Car64, Car65, Glo53c]. Conic [AM58b]. conjecture [CS60, Edr56a, Edr56b, End56a, End56b, Ric70]. consecutive [Ven57]. consecutively [Uhl53b, Uhl53a]. constant [Uhl50b]. constants [Uhl51, Uhl56]. Constructions [Ste50, AS48, Ste48]. constructive [Lig64]. containing [Leh73]. Continued [Leh73, Fra61]. continuity [Rei73]. continuum [Miu73]. contradictions [MB32b]. Contribution [Fle43, CW47, Mal64]. contributions [Boy53a, Boy54c, Loc33]. controversies [Fra47]. convergence [Low56, Low57, Tai65, Tai68]. convergents [Blo55]. convex [Dav59, Lev70, Win57, Win58]. convolution [PS56, PS57]. coordinatization [Sch73a]. Copernican [Kar43]. copies [Bra52]. copying [Uhl53c]. cord [Sei59]. Corpus [Ste51]. coset [iH73]. cosmic [Key34a]. cosmology [Key36a]. Cotes [Hux63]. Couette [Miu73]. Counterpoints [Fel40]. couples [Pou48]. Courses [Ric34]. covering [KCW43]. Cramer [Boy66]. crisis [Fra51]. critical [MB32a]. Cryptanalysis [And52]. cryptography [Men39]. cubes [Gin54a]. cubic [Fel40, Mor63]. cuboids [Kra45]. cuento [Gin32d]. Cultural [Key33a]. Culture [Fle43, Key32b, Key47]. Curiosa [Ann56, Juz39, Lan51, Lan53, Moe47]. curiosity [Sch50c]. Curiosum [Blo53a]. Curious [Moe47, Hea35b]. curve [Emc52]. Curves [Ber54, Ber57b, Ber58b, Ber59a, Bar48a, Bar48b, Fel40, KD49, Mor63]. Curvilinear [Thé53]. cuts [Lig63]. cyclic [Ali53].

D. [Kar46, Mir48, Pal53b]. Damaso [Key36b]. dans [Sie67]. data [Uhl52b]. date [Mal63]. dates [Bra55]. David [Ree45, Ruf41, Sch55, Sim45]. day [Ett56]. debt [Mac44]. December [Ham48a]. decimal [Gin32c, Sch59]. Dedekind [Lig63]. deductive [Sza64a, Sza64b]. Dee [Ros70]. definability [Gri73b]. definition [Ber50, Lig64]. definitions [Sza64a, Sza64b]. degenerate [iH73]. del [Lew32]. demonstrativa [Sac35, Emc35a, Emc35b, Emc35c]. density [Edr57, Edr58, End57]. derivatives [Edr56a, Edr56b, End56a, End56b, Hig40]. Desargues [Cou54c, Cou54d, Ivi43, Ivi47]. Descartes [Boy52, Boy53b]. Descriptions [Sel55]. design [Boy49]. destiny [Uhl51]. Determinants [Sta53]. determination [Aab54, Ken59]. developable [Gol52]. development [Boy46, CW47, Kar45]. developments [Van61]. deviation [RU66]. diagonal [MZ63]. Diagonalverfahren [Rus34]. dictionary [Sim35b]. Differential [Gor73, Arc32c, Ber57a, Ber58a, Ber61, CW47, Fin63, Rei73, Sch73b]. differentiation [Spe63]. difficulties [AM59, iAa60]. Digit [Sta53, Uhl53b, Uhl53a]. digits [Kap53, MM50, Sha58]. dimensional [Kap73, Mal64, Sel73]. Dinner [Ano37]. diophantine [Arc33, Arc34i, Bro39, Glo49b, Glo53b, Glo48b, Glo52a, Glo52b, Glo60b, Mor64, Mor65, Mos53, Pal53a, Whi51]. direct [Löw46]. directed [Har57, Har58]. disc [Lan51]. discarded [Boy41]. disciple [Ruf41]. Discordant [Rio54, Tou53b]. discovery [Gan32]. discrete [PS56, PS57]. Discussion [AM59, iAa60, Las38]. disputation [Wei35]. Disquisitiones [GA35]. dissolution [MB32b]. distribution [MM50]. distributions [Gil64a, Gil65]. Divisibility [Van55, Bre55, Grü55, Kra48]. division [Alb63]. divisor [Rao70]. doctrinal [Key34c, Key34d]. document [Smi34]. dodecahedron [Hil40]. Dr. [Sze60]. Droz [Goo50]. dual [Ber50, Mor39]. dual-symmetric [Ber50]. Duality [Div73]. duplicity [Hun38a, Hun38b]. Dynamic [Bow32, Bel56]. dynamics [Bat44].

Early [Boy45, Boy53a, Boy54c, Dor50, Gin32c, Sim35a, Syn44, Syn45, Yal32]. East [AM56]. Easy [Lan53]. eccentric [Neu59]. eclipse [Ken51]. economist [Key36b]. edited [Mur64]. editions [Ros59]. editor [Kar35]. Edmond [Hux59, Hux60b]. education [Ree38]. eighteenth [Sim35a]. Einstein [Tal32]. elasticity [Mal64]. Electronic [Lap52, Lap53]. Elementary [JT45, Coh60, Low55, Mur44, Smi35a]. elements [Ber50, Cou51a, Cou51b, Ven54c]. elusive [Kar35]. emblem [Yal32]. Emmy [Wey35]. Encapsulated [Rol19]. Engelhardt [Ric39]. engineer [Lor47b]. England [Sim34a]. English [Sim35b, Str49]. entire [Edr57, Edr58, End57]. Epicurus [Key36a, Las38]. epicyclic [Neu59]. equal [Man61, Whi43a, Whi43b]. equation [Boy46, Bri60, Bro39, Dor48, Glo48b, Mor64, Mor65, Mos53, Thoxx, Ven58, Whi51]. Equations [Glo52b, Arc33, Arc34i, Ber57a, Ber58a, Ber61, Boy45, Fin63, Glo49b, Glo52a, Glo53b, Glo60a, Glo60b, Gut47, Hod66]. equiareal [Gin54b, Rai54]. Equipment [Gar52]. equivalence [Neu59]. Equivariant [iH73]. Erdos [Coh63, Del57, Del58]. Ernest [Lor47a]. Escott [Leh47, Mir48, Pal53b, Rob50]. essay [Cla34]. essays [Key47]. essential [Sch73b]. estimate [Coh63]. Euclid [AM59, Eas66, Smi35b, WW60, iAa60]. Euclidis [Mur64]. Eudoxean [Har70]. Eudoxus [Fra61, Neu53, Neu54].

Eugene [Ree45, Sim45]. Euler [Boy50a, Boy50b, Car61, Coh57, Coh58, De 47, Goo70, Lan35a, Lan35b]. every [Bee50, Ett56]. every-day [Ett56]. Exact [Uhl53b, Uhl53a, Uhl56, Uhl49b, Uhl55]. excluded [Fra51]. experiences [Uhl50a]. exponentials [Gin45]. exponents [BF38a, BF38b]. Extension [Kan56, Car64, Car65, De 47, Fur54, Sch50a, Sch50b]. extensions [Wym73]. exterior [Spe63]. Extremal [RU66]. eyes [Eis51]. Ezra [AM57, AM58a].

factorial [Uhl49a]. factorials [Kra48, Uhl55]. factorisations [Glo48a]. factorization [Kra52, Wei55]. factors [Jon39]. Fair [Rol19]. Farny [Goo50]. Father [Bel38a, Bel38b]. February [Ham48a]. Federico [Key36b]. female [Coo51, Kra57, Kra58]. Fermat [Bel49, Boy47, Boy52, Boy53b, CW47]. Fibonacci [Bea50, Blo53a, Blo53b, Can51, Gin54a, Kel57, Kel58, Rai48, Rai54, Rao55, Rei54, Sch50c]. field [Hod66]. fields [Che73, Wym73]. fifteenth [Ken51]. figure [Uhl51, Uhl52b, WC60]. figures [Thé53]. Finite [ABG73, Hod66]. finitely [Gil73]. first [MM50, Sim34c, Uhl54]. Fisher [Kar35]. fixed [Alt60, AS48, CP73, Ste48]. flow [Miu73]. fluxions [Cla34, Sim36]. folding [Tri55]. form [BF73, Dor48, iH73]. formal [Spe63]. forms [Bel47a, Jon73, Mor36]. formula [Boy43, Car52, Car53, Goo70, Gou63, Leh51, Rei54]. formulae [Blo53b, Rao55]. formulas [KM61, Kel57, Kel58]. foundation [Fra47, Sza64a, Sza64b]. foundations [Fra32, Fra33]. four [Fra39a, Fra39b, Fre48, Smi36]. fours [Cox52b]. fourteenth [Gin32h, Tho57, Tho58]. fourth [Can51, CS60]. fractions [Fra61, Leh73]. Fraenkel [Hil34]. Franklin [Bra36]. free [Ber50]. freethinker [Bra49]. French [Boy60, GA35]. Freshmen [Kar39]. Friends [Ano37]. Full [Uhl54]. Fun [Gro50, Gro54, Gro56, Gro59]. function [Boy46, Coh57, Coh58, Dav59, Gup53, Key34c, Key34d, Key35b, Leh51, Thé52a, Thé52b]. functionalistic [Laf41]. functions [BF38a, BF38b, Coh60, De 45, De 47, Del63, Edr57, Edr58, End57, Gil64b, Kau67, Lan63, Rao70, SS50, Sin67, Wei55, Zei73]. fundamental [BF38a, BF38b].

Game [Kap49, Cox53, Fre48, Leh33, Ric46, Saw49]. Gaon [Pol32]. Gap [Edr57, Edr58, End57]. Gaspard [Smi32a]. Gauss [Dun54, Gin38a, GA35]. genera [Jon73]. General [KD47, Ber59a]. generalization [Gar53, Goo70, Gup53, Kle73]. Generalizations [Coh57, Coh58]. Generalized [Alb64, Gil64a, Gil65, LA38]. generate [Gil64b]. generated [Tri55]. generating [Ber57a, Ber58a, Ber61, Low55, Zei73]. generic [McC73a, Oeh73]. Generically [JK73]. Genevra [Eis50]. Geodesics [Thé55]. geodesy [Sch32a]. geometriam [Mur64]. Geometric [Gri60, CS60]. Geometrical [Ste48, Ste50, AS48, Gin38b, Hux60a, Mor36]. geometry [Arc32c, Boy47, Boy53a, Boy54a, Boy54b, Boy54c, Boy55a, Boy55b, Cou51a, Cou51b, De 48, Emc50, GH60, HD35a, HD35b, KD49, Loc41, Lor32, Sei59, Thé49a, Thé49b, Thé52b, Thé53, Thé55, Thé56a, Thé56b, Men56a, Men56b]. Georg [Hil34]. George [Arc35c, Arc35d, Kar35]. Gerard [Ivi43]. Germain [Ano53]. German [Sim34d, Bra38, Sim32a]. Gill [Neu34b]. Ginsburg [Bel58, Boy58]. Giovanni [Tho50a, Tho50b, Tho53, Tho55]. Girolamo [Emc35a, Emc35b, Emc35c]. given [AS48, Ste48, Tou53b]. glance [Key35a]. glass [Wil60]. global [Hen55, Hen56]. Goldbach [Arc35e, Arc35f]. golden [Ban55, Bea50, Cox53, Boy51a, Boy51b]. graphical [Boy45, Ken59]. graphs [Har57, Har58, Tut61]. Great [Bra47, Key35b]. Greek [Hea38, Sei59]. Greenwood [Sim32c, Sim34c]. Gregorian [Bra55]. Gregory [RA51]. group [Cox36, Dav66, Wym73]. Groups [KC36, Ahl64, ABG73, Boc64, Jur59a, Jur59b, KJWC34, Rap59, Sel73].

H [Mir48, Mur64, Pal53b]. Haan [Sch55]. Hahn [KM61]. Halley [Hux59, Hux60b]. Hamburger [Arc32e]. Hamilton [Syn45, Bat44, Mac44, Rap59, Smi44, Syn44, Syn45]. Hamiltonian [Rei73]. Harmonic [Man61, Gil64b]. Hayyim [Dav36]. Hebrew [Kel32]. Henry [WW60]. Hermitian [iH73]. Hexagrams [Bli67, vdB67]. high [Uhl49b]. higher [Car56, Car57, Kau67]. Hilbert [Wym73]. him [Whi32]. Hindu [Raj49, RA51, RA52, Ven54a, Ven54b]. hints [Uhl50a]. Hippopede [Neu53, Neu54]. historian [Lor47b]. historical [Bra40, MB32a]. history [Dij50, Emc52, Gin32c, Gri73a, Hig40, Lor32, Ric34, Sch58, Uhl52a, Key38]. Hitherto [Jon47, Gin34]. Hitler [Fos43]. Hollis [Sim34c]. homogeneous [De 47, Fre53, Wei55]. homomorphisms [San73]. honeycomb [Pol40]. horn [Kas45, KH50, KH51]. Hunting [Leh33]. Hyperbolic [Cox56]. hypergeometric [Zei73]. hyperprojectivity [Gri73b].

Ibn [AM59, AM57, AM58a]. icosahedral [Cox36]. icosahedron [Hil40]. ideas [Key35a]. Identities [Moe53, And70, Glo51, Kla55]. identity [And70, Jon47, Moe52]. II [Arc32b, Arc35f, Arc46b, Ber57b, Ber58b, Boy51b, Boy54b, Bru38b, Coh63, Cou54d, Emc35b, Fra49, HD35b, Jur59b, Men56b, Smi38, Sza64b, Thé56b, Tho73b]. III [Bru38c, Emc35c, MB32b, TT73b]. illustrated [Tho57, Tho58]. illustrating [SS50]. images [San73]. Imaginary [Cou51a, Cou51b]. immense [Uhl52a]. immersion [iH73]. immortal [Mal63]. Implicative [Nem73]. Implicit [Gri73b]. implies [And70]. Inclusion [AF59]. incomplète [Sie67]. independence [Dre53]. Indian [Gan32]. indirect [Cau68, Löw46]. induction [Sie67]. Inequalities [Way46, Low55, Ost63]. inequality [Dre50]. inference [Fry34]. Infinite [Kap73, Lev70, BCKM55]. Infinite-dimensional [Kap73]. infinitesimals [Boy41]. infinity [Key36a, MB32a, MB32b, Van56]. influence [Bat44]. Institute [Arc32f]. instrument [Tho55]. integers [Boy43, Mir49]. integral [BB61, CP73, Dav66]. integration [Coh73]. Interesting [Juz39, Kap55, Smi34]. interests [Bat38]. Interlocked [Cox52a]. interpretation [Gin38b, Laf41]. intersecting [Lev60]. introduction [Sze60]. Inversion [PS56, PS57, Bar48a, Bar48b]. inversions [Way46]. investigations [Uhl52a]. involutorial [Alb63]. involving [Blo53b, Kap53, Kel57, Kel58, Rao55, Rei54]. irrational [AM68, Gan32]. Isaac [Lan36, Sim32c, Sim34c]. isolated [Fin63]. isotope [McC73a]. Italian [Tho53]. Iterated [Gin45]. iteration [Aab54, Low56, Low57]. IX [Arc36e].

J. [Lor36a, Lor36b, Lor36c]. Jacob [AS48, Gin32a, Gin32b]. Japan [Mik32]. Jefferson [Ano32, Smi32c]. Jekuthiel [Bel58, Boy58]. Jewish [Fra60, Ken64, Res43]. Johann [Bel38a, Bel38b]. John [Gri73a, Ros70, Wei35]. Jonquières [Lor47a]. Jordan [JK73, McC73a, McC73b, Osb73]. Joseph [Lew32].

Kāshı [Aab54]. Karkhi [AM57, AM58a]. Keyser [Key47, Las38]. Khayyām [AM59]. Khayyâm [Smi35b]. Khayyam [AM63, AM68, Ano63]. Khayyami [AM59]. Khwārizmī [Ken64]. kind [Car61]. Kleinian [Ahl64]. Knots [Gil56, Gil57]. known [Saw50].

L [Lor32, Mur64]. Laboratory [Low49, Boy49]. Lacroix [Boy47]. Lao [Eis50]. Laplace [Tai65, Tai68]. large [Kal73]. last [Sha58, Smi35a]. latest [Uhl52a]. Latin [Kar47]. Lattice [MZ63, Gro50, Gro54, Gro56, Gro59]. lattices [Nem73]. law [LFA32]. least [RU66]. lectures [WW60]. legend [Key38]. Lehmer [Mir48, Pal53b]. Lehmus [Hen55, Hen56, Lub59]. Leibniz [Loc33, Pau54]. Lemoyne [Goo52]. Léonard [Lan35a, Lan35b]. Leonardo [Wei35]. letter [Gin34]. letters [Sim45]. Levi [Dav36]. Lewis [Arc32g]. Liber [Eis51]. library [Bra40, Sim34a, Kar32]. Lie [Bro73, Fra73, Kap73]. life [Ett56, Kar33, Kar34, Lan35a, Lan35b, Neu34b, Syn44, Syn45]. limit [Lig64]. limits [Boy41]. line [Dor48, Eve59]. linear [Dav59, Glo60a, Rei73]. linearity [Dav59]. linearized [Miu73]. lines [Rap59]. Liouville [Bel47a, Lor36a, Lor36b, Lor36c]. list [Kar46, Str49]. Lists [Kar70]. literature [Gol32]. local [Ken59]. locally [Lev70]. Locating [Sch59]. location [Uhl53b, Uhl53a]. Loewy [Fra38]. logarithmic [Tai65, Tai68]. logarithms [Bar40, Gri73a, Uhl51]. logic [Gou43, Hun38a, Hun38b]. Logica [Sac35, Emc35a, Emc35b, Emc35c]. Logicism [LL64, LL65]. logics [Sin67]. long [Kar33, Kar34]. looking [Wil60]. lore [Kel32]. lower [Thoxx]. Lubelski [Car65]. Lubleski [Car64]. Lucas [Kla55]. Ludwig [Dun35, Pin64, Pin65]. lunar [Ken51].

MA [Ham48a]. Machir [Gin32a, Gin32b]. MacLaurin [Boy66]. Magic [Gar51, Moe47, Bra36, Gut47, Gut48, Hea35a, Hea35b, Hea35c, Hea36, Hor53, Hor55, McC38, McC40, McC41]. magical [Sch32b]. Magini [Gin32f]. magnitudes [AM68]. Makers [Arc48, Arc45b, Arc46a, Arc46b]. making [Löw46]. Malfatti [Arc32h]. Mamoru [Mik32]. man [Boy63, Eis59, Eis60]. manner [Coh73]. Mannheim [Lor34]. Manuel [McC41]. manuscript [Bra34, Ric43a, Ric43b, Ric43c, Ric39, Sch32a, Sch34]. manuscripts [Bra40, Sch32b, Tho50a, Tho50b, Tho53, Tho57, Tho58]. Many [Uhl51, Uhl52b, Gil73]. Many-figure [Uhl51, Uhl52b]. Map [Cox57, Cox58]. Map-coloring [Cox57, Cox58]. maps [RW60]. march [Bru38a, Bru38b, Bru38c]. Mary [Ric41]. master [Mik32]. Mathematica [Ano37, FKCR38, Fle43, Gin32e, Gin37, Key33a, Rol19, Sla33]. Mathematical [Arc32e, Arc34d, Arc34e, Arc34f, Arc34g, Arc36e, Arc36f, Arc36g, Arc36h, Arc45b, Arc46a, Arc46b, Arc48, Cou47, Gar48, Gar49, Gar52, Sim35b, Bat38, Bra40, Fry34, Hux59, Hux60b, Kar32, Kar41, Kar45, Kar54, Key38, Ros59, Ruf41, Sim32b, Smi35a, Ste51, Str49, Uhl53c, Sze60]. mathematician [Boy63, Dav36, Fry64, Gin32h, Lor34, Nev38, Key36b]. Mathematicians [Boy60, Coo51, Kra57, Kra58]. Mathematicis [Arc35b, Arc32a, Arc32b, Arc34a, Arc34b, Arc34c, Arc35a]. Mathematics [Fos43, Kar39, Kar47, Key32b, Key33b, Key34a, Key34b, Key47, Key49, LFA32, Wil60, AC35, Bru38a, Bru38b, Bru38c, Dor50, Ett56, Fra32, Fra33, Fra43a, Fra43b, Fra43c, Fra47, Fra49, Fra60, Hea38, KH50, Key32a, Key32c, Laf41, Lar40, Mal63, Mik32, Neu34a, Raj49, Ree35, Ree38, Ric34, Sim34a, Sim41, Smi32c, Smi34, Sta38, Sza64a, Sza64b, Tho57, Tho58, Arc32f]. matrices [Jur59a, Jur59b, Leo54, Ost63]. matrix [Hod66, TT73a, Tho73b, TT73b]. Maupertuis [Sch58]. Maurolico [Ros59]. Maxima [Gol32]. maximal [Bro73]. Mean [Tai65, Tai68, Uhl49a]. Meaning [Key49, Key32c]. means [Tai65, Tai68]. Measure [Coh73]. measures [Boc64]. mechanical [Loc33]. medical [Kel32]. medieval [Gol64, Neu56, Neu57]. Medigo [Lew32]. memory [Pin64, Pin65, Syn45]. menage [Gil57, Car54, Gil56]. ménages [KR46]. Menelaus [Fur54, Gin32a, Gin32b, Sch34]. Mengenlehre [Bri34a, Bri34b]. meridian [Ken59]. Meromorphic [Che73]. Mersenne [Arc35g, Uhl52a]. method [Aab54, Gar35, Glo60a, God35, Kan56, Low55, Pol32, Sim36]. methods [AF59, AM57, AM58a, Fra43a, Fra43b, Fra43c, Fra49, Lar41]. Mexico [Bra52]. Michigan [Kar32]. middle [Fra51]. might [Cou56]. Miller [Arc35d]. Mimori [Mik32]. mind [Key34a]. mindful [Whi32]. minima [Gol32]. minimal [Glo55]. minister [Sim34a]. minor [Sch50c]. minus [EK46]. Miscellaneous [Boy66, Uhl50a]. mixed [Hig40]. mixtilinear [Thé53]. Möbius [Gup53]. Models [SS50]. modern [Bat44, Bru38a, Bru38b, Bru38c, Fra32, Fra33, Fra43a, Fra43b, Fra43c, Fra49, Ree38]. modification [WC60]. modules [Gor73]. Móez [iAa60]. Moivre [Wal34]. Monge [Smi32a]. Monuments [Arc48, Arc45b, Arc46a, Arc46b]. Moritz [Smi32b]. Moschopoulos [McC41]. motion [Neu59]. motions [Har70]. Multigrade [Glo53a, Glo51, Gut47]. Multigrades [Ven54c]. multinomial [Car63]. Multiperfect [Bro54]. multiple [Moe52]. multiplication [Hor53, Hor55, Ira52, Pol32]. multiplicative [Del63]. multiply [FG54, Gar53]. multipoint [KC36]. Music [AM56]. Musical [Bar40]. Muskhelishvili [Mal64].

Napier [Gri73a, Mor39]. National [Low49]. Natural [Fra40, Hux63, Uhl53b, Uhl53a]. nature [Key34c, Key34d]. near [Miu73]. near-continuum [Miu73]. neglected [Raj49]. Neo [FF56a, Kas47]. Neo-Pythagorean [FF56a, Kas47]. Neopythagorean [FF56b]. Nests [Whi53]. Netherlands [Dij50]. Newton [Boy50a, Boy50b, Bra52, Lan36, Ruf41]. Nicole [Mur64]. Nieuwland [Sch50a, Sch50b]. Nights [Uhl49a]. nilpotence [McC73b]. Nine [Sta53, Uhl55]. Nine-Digit [Sta53]. nineteenth [Eis59, Eis60]. No [BM53, Fin63]. Noether [Wey35]. nombres [Sie67]. non [Dav59, iH73, Kas43, Lev60]. non-Apollonian [Kas43]. non-degenerate [iH73]. non-intersecting [Lev60]. non-linearity [Dav59]. nonlinear [Thoxx]. norm [CP73, McC73a]. normal [Dor48]. North [Ham48a, Ham48b]. Note [Car52, Car53, Coh64, Coh65, Dun54, Glo49b, Jon39, Kas43, Las38, Lev60, Low55, Neu34b, Sim34d, Ano32, Car56, Car57, Car61, Gil73, God35, Hig40, Ivi43, Ivi47, Kau67, Leh63, Ros67, Uhl52c]. Notes [Arc48, Glo52a, Glo52b, Glo53b, Hil40, LA38, Ric39, Sze60, Arc45b, Arc46a, Arc46b, Boy66, Coh63, Cox52b, MB32a, Ric46]. Nullifying [Jar54]. Number [Kap55, Ali53, Bra52, Fre48, Gra54, Gri60, JT45, Leh51, Lev60, Mor60, Thoxx, Van55]. Numbers [AM56, Kha55, Arc35g, Bea50, Bee50, Blo53b, Bro54, Car56, Car57, Che73, Esc46, Fra40, FG54, Gar53, Gin54a, Grü55, Kel57, Kel58, Leh33, LA38, Lub59, McC57, McC58, Pou48, Rao55, Rei54, Tou53a, Uhl52a, Uhl53b, Uhl53a, Uhl53d, Uhl54, Van61, Ven57, Ven54c]. Numerics [Kel32].

obeying [Sch73c]. Obituary [Ham48b, Ree45]. Objects [Gar49]. Observation [Juz39]. obvious [Lan53]. Odd [McC57, McC58]. old [Gut48, Kel32]. Omar [AM59, AM59, AM63, Ano63, Fra61, Smi35b]. only [Gil73]. onto [RW60]. operational [Rus34]. operations [Dre50]. operator [Thoxx]. Operators [Ber57a, Ber58a, Ber61, Gil64b, Sch73c, BB61, Kal73, Sch73b]. order [Car56, Car57, Dre50, Hor55, Kau67, Way46]. ordinals [Fra40]. Oresme [Mur64]. Origin [Kar39]. orthogonal [CP73]. other [Key47, Uhl51]. Oware [Saw49]. Oxford [Sim35b].

packing [Kas43]. Paintings [Arc48, Arc45b, Arc46a, Arc46b]. pairs [Gar57, Gar58, Hod66, Lev60]. Palermo [Wei35]. palindromic [Ven54c]. panelled [Hea36]. paper [AM63, Ano63]. Papers [Syn45]. parabola [Uhl50b, Uhl52c]. Pareto [Key36b]. Paris [Pau54, Tho50a, Tho50b]. partial [Ber57a, Ber58a, Fin63, Hig40, Sch73b]. partition [Cox36, Goo70, Leh51]. parts [Mor39]. Pascal [Boy43, Boy63]. paths [MZ63]. Paul [Lor47b]. Peg [Sei59]. Peirce [Eis51, Eis59, Eis60, Gin34, Key35a, Key41]. Pell [Ven58]. Pennsylvania [Bra38]. pentagons [Dou60]. perfect [FG54, Gar53, McC57, McC58, Tou53a, Uhl53d, Uhl54]. Permutations [Tou53b, Gil56, Gil57, Rio54]. Personal [Tal32]. personality [Sho34a, Sho34b]. perspective [AM58b]. Peruvian [Loc32b]. Philo [Eve59]. philosophy [Hux63, Lar40]. Phyllotaxis [Cox53]. physical [KD49]. physicist [Bri34a, Bri34b]. pi [RA52]. piecewise [Dav59]. piecewise-linear [Dav59]. Pierre [Sch58]. Pieter [Bra49, Sch50a, Sch50b]. pioneer [Key41]. Pisa [Wei35]. place [Kar45, Ree38]. plane [Cox55, Emc50, KCW43, Miu73]. planes [San73]. planetary [Har70, Neu56, Neu57]. Plateau [Dou38]. Platonicum [Ste51]. please [BM53]. plight [Cou56]. plus [EK46]. poem [Whi32]. point [Alt60, Gin32c, Kal73, Sch59]. points [Emc52, Gro50, Gro54, Gro56, Gro59, Kla54, Mor63]. pointwise [Tai65, Tai68]. polar [KD47]. policy [Gin32e]. politician [Smi32a]. Polya [End56b, Edr56a, Edr56b, End56a]. polygenic [De 45]. polyglot [TT40]. Polygons [Ber59b, KJWC34, KC36]. polyhedra [Cox39, Cox54]. polynomial [Alb67, And70, Boy45, Lan63, Oeh73]. Polynomials [Lan63, Car54, Car61, KM61, RU66]. popular [Kar35]. portrait [Wat39]. Portraits [Arc48, Arc45b, Arc46a, Arc46b]. positive [Jur59a, Jur59b]. possible [Wat39]. Post [Boy55b]. Post-Cartesian [Boy55b]. postulate [Arc45a]. postulated [Ber50]. Postulational [Lar41]. potentials [Gil64a, Gil65]. power [Sch73a]. power-associative [Sch73a]. powers [Boy43, Can51, MM50, Uhl49b]. pre [Zlo60]. pre-axiomatic [Zlo60]. Predecessors [Gin32f]. presentation [Mur44]. prime [Bee50, Gra54, Tou53a]. primes [JT45, Kar70, Uhl52a, Van56]. primitive [Leh63]. Prince [Sch50a, Sch50b]. principal [Rei73]. Principia [Bra52]. principle [Fra51, Zlo60]. printed [Kar41, Kar54]. probability [Bat38, Düt53, Gri60]. problem [Arc32h, Arc40, Bea55, Bel47a, BF73, Cou54a, Cou54b, Dou38, Fra39a, Fra39b, Gin38a, Gro45, Hen55, Hen56, Lan54, Leh47, Mir48, Mos57, Pal53b, Rob50, Sch50a, Sch50b]. problème [KR46]. Problems [Fra43a, Fra43b, Fra43c, Fra49, Kap53, Cox57, Cox58, Emc50, Fra32, Fra33, Low50, RU66, Sie60, Tut61, Van55]. processes [Low56, Low57]. prodigy [Key38]. products [TT73a, Tho73b, TT73b]. professor [Sim34c, Bel58]. programming [Bel56]. progress [Kar43]. progression [Kar70]. progressions [Leh73]. projection [Lev70]. Projective [Sie63, Arc32c, iH73, San73]. proof [Cau68, CS60, RA51]. proofs [Löw46]. properties [Ber57a, Ber58a, Ber61, Blo55, Car54]. Property [Kap55, Lev70, Rei73, Uhl50b]. Proportion [Boy46]. prosthaphaeresis [Lew32]. pseudo [San73]. pseudogroups [Spe63]. publication [Sim32b]. Publications [Sla33, Arc32e, Kar47, Sim41]. published [Cla34, Dij50, Kar46, Str49]. pure [Cou51a, Cou51b]. pursuit [Ber54, Ber57b, Ber58b, Ber59a, Ber59b]. puzzle [Lan51, Saw50]. Pythagoras [Key39]. Pythagorean [Bel47b, FF56a, Kas47, Kha55, LA38, Rai48, Ric46, Rob50, Whi46, Whi53].

Quadratic [Arc33, Arc34i, Bel47a, JK73, Jon73, McC73b]. Quaestiones [Mur64]. Quasi [Jon73]. Quasi-genera [Jon73]. quaternions [Mur44]. queens [Gin38a]. question [Key34a].

R [iAa60]. rôle [Key34d]. rabbinical [Gol32]. radical [Div73]. radicals [Osb73]. Ramanujan [And70]. Ramsgate [Ham48a]. random [Alb64]. rank [ABG73]. Rare [Emc50, Kar32]. Rational [Mor63, Whi43a, Whi43b, Fry34, Gin54b, Key34c, Key34d, Kla54, Kra45]. reaction [Bri34a, Bri34b]. readyreckoner [Sim35a]. reappraisal [Ste51]. reasoning [Cou56]. Rechenbüchlein [Bra34]. reckoning [Gol64]. recollections [Tal32]. Reconstructing [Har70]. reconstruction [Sho34a, Sho34b]. Recreational [Thé49a, Thé49b, Thé52b, Thé53, Thé55, Thé56a, Thé56b]. region [SS50]. regular [Cox39, Hil40, Van56]. Regularity [Ric70]. relating [Smi34]. relation [Key35b]. relations [AF59, BB61]. relationship [Can51, Gin54a]. relative [Alt60]. remark [Mir48]. Remarkable [Glo51]. representation [Emc52]. Representations [Osb73, Mir49]. Reseña [Med40]. Restrictedness [Fra73]. reversal [Kap53]. Review [Mur64]. Reviews [Sel55]. Revolution [Boy60]. right [Whi43a, Whi43b]. rings [Cox52a, Div73, Gil73, Gor73, Kle73, Ros67]. Rittenhouse [Ruf41]. road [Loc41]. Robert [Arc34h]. Roger [Hux63]. Rogers [And70]. role [Ett56, Key34c, Key36a]. roots [GH60, Leh63, Tab38]. Rouse [Cox52b]. routes [Lev60]. Rowan [Smi44, Syn44, Syn45]. royal [Loc41]. rule [Boy66]. Ruler [Ste50, AS48, Ste48]. rules [Mor39]. Rupert [Sch50a, Sch50b]. Russell [Med40].

S [Eis51, Eis59, Eis60, Arc36i, Rol19]. Sûtras [Gan32]. Saadia [Pol32]. Saccheri [Emc35a, Emc35b, Emc35c, Smi35b]. sad [Mal59]. sailor [Lor47a]. Sanders [Key35a, Key41]. Savile [WW60]. scale [Mir49, Tab38]. Schlesinger [Dun35]. Schlicht [SS50]. School [Arc32f]. schools [Ree35, Sim36]. Schröder [Laf47]. Schwarz [Dre50]. Schwenkfelder [Bra40]. Science [Gou43, Key33b, Bru38a, Bru38b, Bru38c, De 48, Dij50, Eis59, Eis60, Key34b, Sza64a, Sza64b]. sciences [Kar45, GA35]. scientific [Gol73]. scientist [Lor47a]. Scripta [Ano37, FKCR38, Fle43, Gin32e, Gin37, Key33a, Rol19, Sla33]. second [Bri34a, Bri34b, Car61]. section [Bea50, Cox53]. sections [AM58b]. selected [Str49]. Semantics [Key33b, Key34b]. semi [Jur59a, Jur59b, Nem73]. semi-groups [Jur59a, Jur59b]. semi-lattices [Nem73]. semisimple [Div73]. Sequel [Uhl52c]. sequence [Lig64, Uhl53b, Uhl53a]. Sequences [CP73, EK46]. series [BCKM55, Can51, Gou63, Rai48, RA51, Tai65, Tai68]. set [Dav59, Zlo60]. sets [Sie63]. sevens [Bel43, Bel44, Bel45a, Bel45b, Bel46]. seventeen [Uhl54]. seventeenth [Sim34a]. sexagesimal [Ira52]. Sheffer [Sin67]. Shiragooni [Vog35]. Shirak [Vog35]. Sicily [Bru38a, Bru38b, Bru38c]. sides [Thé52a, Thé52b]. Sierpiński [Fry64]. Significance [Key33a]. Simons [Eis50]. simple [ABG73, Lig63]. simplicity [Fra73]. Simson [Arc34h]. sine [Kau67, Win57, Win58]. singular [TT73a, Tho73b, TT73b]. singularities [Fin63]. Sir [Syn45, Smi44, Syn44, WW60]. Six [Coo51, Cox54, Kra57, Kra58]. Sixes [Bel43, Bel44, Bel45a, Bel45b, Bel46]. skeleton [Uhl56]. skew [Cox39, Dou60]. slit [RW60]. Smith [Sim45, Key34e, Ree45, Sch34]. Society [Ano37, Arc32e]. sociologist [Key36b]. solar [Gol64]. soldier [Lor34]. Solitaire [Kap49]. Solution [BF73, Glo50, Glo53c, Bro39, Glo48b, Glo60a, Mir48, Miu73]. solutions [Ber57a, Ber58a, Ber61, Boy45, Dav66, Fin63, Glo55, Rei73, Thoxx]. Solvability [McC73b]. Some [Bel56, Blo55, Emc52, Hod66, Leo54, Rao55, Sin67, BB61, BCKM55, Key35a, Kla55, Ost63, Rao70, Sie60, Tho73a]. Somerville [Ric41]. son [Bel38a, Bel38b]. Sophie [Ano53]. space [BF73, Dor48, iH73, Kas43]. space-form [BF73]. spaces [Alt60, Lev70]. Spanish [Gin32g]. specia [Moe53]. Special [Gar52, Ber50, Ros67]. spectrum [Kal73, Sch73b]. sphere [Cox36]. spheres [Cox52a]. spherical [Gin32a, Gin32b, HD35a, HD35b]. Spinoza [Düt53]. spiral [Tab38]. sponges [Cox39]. Square [GH60, Moe47, Bra36, Hea35b, Hea36, Hor55, Tab38, Tri55]. squares [Arc40, Gut47, Gut48, Hea35a, Hor53, McC38, McC40, McC41, Ven57]. squaring [Gol52]. Standards [Low49]. statement [Rus34]. States [Bra52, Ric34]. statistics [Bat38]. Stefan [Ste63]. Steiner [Hen55, Hen56, AS48]. steps [Boy46, MZ63]. straight [Dor48]. strange [Cou54c, Cou54d]. structure [Pol40]. Study [Arc32f, BF38a, BF38b, Tho57, Tho58]. subalgebras [Bro73]. subrings [Gil73]. substitutions [Dav66]. successive [Edr56a, Edr56b, End56a, End56b]. suggested [WC60]. Śulva [Gan32]. Śulva-Sûtras [Gan32]. sum [Fre53]. summation [AF59, Rao55, Rei54]. summations [Blo53a]. sums [Boy43, Lev70, Ven57, Wei55]. sundial [Yal32]. super [Mur64]. Supplement [Kar41, Lor57, Lor58, Kar54, Sim35b]. surface [HD35a, HD35b]. surfaces [Gol52, Mor63]. Survey [De 45, KD49, Bra52, Hen55, Hen56, Kar47]. symbol [Gin32g]. Symbolic [Blo53b, Gou43]. symmetric [Alb64, Ber50, Fre53, Gil64a, Gil65]. Symmetrical [Tut61]. symmetry [Bow32]. Synchronism [Loc32a]. synonyms [Key35b]. system [Glo60a, Har70, Ira52]. systems [Alb67, Glo49b, KD49, Pal53a, Rei73]. Szekeres [Coh63]. Szily [Car52, Car53, Gou63].

Table [Arc48, Glo55, Uhl49b, Arc36i, Arc45b, Arc46a, Arc46b, Gol64, Ira52]. Tables [Sel55, Arc34d, Arc34e, Arc34f, Arc34g, Arc36e, Arc36f, Arc36g, Arc36h, Uhl56]. Talk [FKCR38]. Talmud [Sze60]. Tannery [Lor47b]. Tarry [Leh47, Mir48, Pal53b, Rob50]. Taught [Kar39]. teacher [Bra49, Mik32]. teaching [Ree35]. ten [Kar70, Sha58]. term [Gin32d]. terms [Can51, Sim35b, Smi35a]. ternary [Hes73, Nem73]. Terquem [Hen55, Hen56]. tesselations [Gol56]. Testing [Bre55]. tetrahedron [Cou48]. text [Ric46]. textbook [Bra49]. textbooks [Kar34, Kar33]. th [Dav36, Uhl53b, Uhl53a]. Thébault [Gui61]. their [Ber57a, Ber58a, Ber61, Kar47]. theorem [Alt60, Arc35e, Arc35f, Car63, Cou54c, Cou54d, De 47, Del57, Del58, Dou60, Fur54, Goo50, Goo52, Gra54, Her73, Ivi47, Laf47, Lor57, Lor58, Lub59, Mir49, Pal53b, Sch73a, Sch73c, Wei55]. theorems [Bel47a, Edr57, Edr58, End57]. théorie [Sie67]. theories [Neu56, Neu57]. theory [BF38a, BF38b, Bel56, Düt53, Fry34, Kar43, Kas45, KD47, Leh33, Mor60, Mur44, Ost63, Van55, Van61, Zlo60]. there [Van56]. Third [Kar54, SS50]. thirteenth [Tho57, Tho58]. Thomas [Ano32, Smi32c]. thought [Bat44, Key34c, Key34d]. thousand [Gin32g, Smi36]. Three [Key35b, Ber57a, Ber58a, Ber61, Boy46, Cox52b, Gil64b, Kar35, Smi35a, Tho57, Tho58]. three-variables [Gil64b]. Time [Nev38, Gan32, Gol64]. Topology [Gar51]. Town [FKCR38]. traceable [Har57, Har58]. Tranl [iAa60]. Transformation [Bar48a, Bar48b, Key35b, Sza64a, Sza64b]. transformations [Coh64, Coh65, KC36]. transforms [PS56, PS57]. transitive [Dav66]. translated [Ric46, Str49]. translation [Ste50]. transmission [Neu56, Neu57]. treatise [McC41]. treatment [HD35a, HD35b, Low50]. Trees [HM70]. triangle [Thé52a, Thé52b, Thé56a, Thé56b]. Triangles [Kha55, Cox56, FF56a, FF56b, Gin54b, Kas47, Rai48, Rai54, Rob50, Whi43a, Whi43b, Whi53]. Triangular [Kha55, Leh51]. Triangulations [Dun54]. Tricks [Gar48, Gar49, Gar52]. trigonometry [Bra47, Gin32a, Gin32b, Kar45, Kar46]. Triplets [Gin54b]. trisection [Gar35, God35, KH51]. Tudor [Eas66]. twelve [Gro45]. twelve-coin [Gro45]. Two [Mal64, Sch32b, ABG73, Ros67, Sel73, Tou53b]. two-dimensional [Sel73]. type [Bro73, Pal53a, Spe63]. types [Ros67].

Undergraduate [Sim41, Wat39]. Uniform [Cox54, Miu73]. unitary [Rao70]. United [Bra52, Ric34]. Universal [Gut47]. University [Kar32, Sch34]. unknown [Gin32h, Kas34]. unpublished [Cox52b, Dor50, Gin34, Ric39]. unsolved [Sie60]. use [Gin32d]. used [Pol32].

V [Arc34a]. Valentin [Ric39]. valued [Sin67]. values [TT73a, Tho73b, TT73b, Uhl49b, Uhl51, Uhl54, Uhl56]. variables [Ber57a, Ber58a, Ber61, Gil64b]. variation [Bea55, Bel47b]. variations [Whi46]. various [Low56, Low57]. vector [Che73, Low50]. vector-analytical [Low50]. vectors [CP73]. Venema [Bra49]. version [Gin32a, Gin32b]. VI [Arc34b]. Victor [Gui61]. VII [Arc34c]. VIII [Arc35a]. Vilfredo [Key36b]. Voltaire [KH50]. Vorlesungen [Arc34h].

Wacław [Fry64]. walking [TT40]. walks [Alb64]. Wallis [Bel49]. Waring [Arc40]. Weighted [Tho73a, Uhl49a]. weighted-mean [Uhl49a]. well [Saw50]. well-known [Saw50]. West [AM56]. Weyl [Sch73c]. which [And70, Bee50, Gil64b]. Whitehead [Ham48a, Ham48b]. whose [Fin63, Gil64a, Gil65]. William [Smi44, Syn44, Syn45, Key34e]. Wolfgang [Bel38a, Bel38b]. Wolfram [Arc36i]. Work [Jon47, Bat44, Bra55, Dor50, Gin32a, Gin32b, Hux59, Hux60a, Hux60b, Lor36a, Lor36b, Lor36c, Mor60, Syn44, Syn45]. Works [Dij50, Key47, Kar41, Kar44, Kar46, Kar54, Ros59]. world [Bru38a, Bru38b, Bru38c, Rol19]. Wren [Hux60a, Sim34b]. writer [Kar35]. written [Uhl53b, Uhl53a]. Wythoff [Cox53].

X [Arc35b, Arc36f]. XI [Arc36g]. XII [Arc36h]. XIII [Dav36]. XIII-th [Dav36].

year [Uhl51]. years [Dij50, Smi36, Smi38]. Yeshiva [Rol19]. youth [Tal32].

Zermelo [Kan56]. zeros [Edr56a, Edr56b, End56a, End56b]. Zuckermann [Sze60].

References

Aaboe:1954:AKI

[Aab54] Asger Aaboe. Al-Kāshǐ’s iteration method for the determination of sin 1^{∘}. Scripta Mathematica, 20(??):24–29, ???? 1954. ISSN

         0036-9713. URL https://www.jphogendijk.nl/samarkand/Kashi-
         Aaboe.pdf.

Alperin:1973:FSG

[ABG73] J. L. Alperin, Richard Brauer, and Daniel Gorenstein. Finite simple groups of 2-rank two. Scripta Mathematica, 29 (3–4):191–214, ???? 1973. ISSN 0036-9713. URL https:

         //citeseerx.ist.psu.edu/document?repid=rep1&type=pdf&doi=
         13cb57389503673a211ebc0b104bbc5d2a070235.

Altshiller-Court:1935:AM

[AC35] Nathan Altshiller-Court. Art and mathematics. Scripta Mathematica, 3(??):103–111, ???? 1935. ISSN 0036-9713.

Agnew:1959:IRA

[AF59] R. P. Agnew and W. H. J. Fuchs. Inclusion relations among absolute Abel summation methods. Scripta Mathematica, 24(??):133–136, ???? 1959. ISSN 0036-9713.

Ahlfors:1964:KG

[Ahl64] Lars V. Ahlfors. Kleinian groups. Scripta Mathematica, 27(??):97– 103, ???? 1964. ISSN 0036-9713.

Albert:1963:IAD

[Alb63] A. A. Albert. On involutorial associative division algebras. Scripta Mathematica, 26(??):309–316, ???? 1963. ISSN 0036-9713.

Albert:1964:GSR

[Alb64] Eugene Albert. Generalized symmetric random walks. Scripta Mathematica, 27(??):185–187, ???? 1964. ISSN 0036-9713.

Albert:1967:CPS

[Alb67] A. A. Albert. On certain polynomial systems. Scripta Mathematica, 28(??):15–19, ???? 1967. ISSN 0036-9713.

Aliaga:1953:CN

[Ali53] C. A. Costa Aliaga. The cyclic number 142857. Scripta Mathematica, 19(??):181–184, ???? 1953. ISSN 0036-9713.

Altman:1960:FPT

[Alt60] M. Altman. A fixed point theorem in relative Banach spaces. Scripta Mathematica, 25(??):197–199, ???? 1960. ISSN 0036-9713.

Amir-Moez:1956:NME

[AM56] Ali R. Amir-Moéz. Numbers and music of East and West. Scripta Mathematica, 22(??):268–70, ???? 1956. ISSN 0036-9713.

Amir-Moez:1957:CMI

[AM57] A. R. Amir-Moéz. Comparison of the methods of Ibn Ezra and Karkhi. Scripta Mathematica, 23(??):173–178, ???? 1957. ISSN 0036-9713.

Amir-Moez:1958:CMI

[AM58a] A. R. Amir-Moéz. Comparison of the methods of Ibn Ezra and Karkhi. Scripta Mathematica, 23(??):173–178, ???? 1958. ISSN 0036-9713.

Amir-Moez:1958:CSP

[AM58b] Ali R. Amir-Moéz. Conic sections in perspective. Scripta Mathematica, 23(??):229–232, ???? 1958. ISSN 0036-9713.

Amir-Moez:1959:DDE

[AM59] A. R. Amir-Moéz. Discussion of difficulties in Euclid by Omar Ibn Abrahim al-Khayyami (Omar Khayyām). Scripta Mathematica, 24 (??):275–303, ???? 1959. ISSN 0036-9713. Translated by Ali R. Amir-Móez.

Amir-Moez:1963:POK

[AM63] Ali R. Amir-Moéz. A paper of Omar Khayyam. Scripta Mathematica, 26(4):323–337, ???? 1963. ISSN 0036-9713.

Amir-Moez:1968:KIM

[AM68] Ali R. Amir-Moéz. Khayyam and irrational magnitudes. Scripta Mathematica, 28(??):205–208, ???? 1968. ISSN 0036-9713.

Andree:1952:C

[And52] Richard V. Andree. Cryptanalysis. Scripta Mathematica, 18(??):5–15 (1 plate), ???? 1952. ISSN 0036-9713.

Andrews:1970:PIW

[And70] G. E. Andrews. A polynomial identity which implies the Rogers– Ramanujan identities. Scripta Mathematica, 28(??):297–305, ???? 1970. ISSN 0036-9713.

Anning:1956:C

[Ann56] N. Anning. Curiosa 444. Scripta Mathematica, 22(??):227–??, ????

1956. ISSN 0036-9713.

Anonymous:1932:NTJ

[Ano32] Anonymous. A note on Thomas Jefferson. Scripta Mathematica, 1 (??):87–90, ???? 1932. ISSN 0036-9713.

Anonymous:1937:DSF

[Ano37] Anonymous. Dinner of the Society of the Friends of Scripta Mathematica. Science, 85(2212):492, May 1937. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See corrections [Gin37].

Germain:1953:SG

[Ano53] Anonymous. Sophie Germain. Scripta Mathematica, 19(??):189, ????

1953. ISSN 0036-9713.

Anonymous:1963:POK

[Ano63] Anonymous. A paper of Omar Khayyam. Scripta Mathematica, 26: 323–337, 1963. ISSN 0036-9713.

Anonymous:1973:AAa

[Ano73a] Anonymous. A. Adrian Albert. Scripta Mathematica, 29(3–4):i–iii, ???? 1973. ISSN 0036-9713.

Anonymous:1973:AAb

[Ano73b] Anonymous. A. Adrian Albert. Scripta Mathematica, 29(3–4):iv–v (1 plate), ???? 1973. ISSN 0036-9713.

Archibald:1932:BM

[Arc32a] R. C. Archibald. Bibliographia de mathematicis. I. Scripta Mathematica, 1(??):173–181, ???? 1932. ISSN 0036-9713.

Archibald:1932:BMI

[Arc32b] R. C. Archibald. Bibliographia de mathematicis. II. Scripta Mathematica, 1(??):265–274, ???? 1932. ISSN 0036-9713.

Archibald:1932:BPD

[Arc32c] R. C. Archibald. Bibliographies of projective differential geometry. Scripta Mathematica, 1(??):171–172, ???? 1932. ISSN 0036-9713.

Archibald:1932:BB

[Arc32d] R. C. Archibald. Boncompagni’s Bullettino. Scripta Mathematica, 1 (??):87, ???? 1932. ISSN 0036-9713.

Archibald:1932:HMS

[Arc32e] R. C. Archibald. The Hamburger Mathematical Society publications. Scripta Mathematica, 1(??):85–86, ???? 1932. ISSN 0036-9713.

Archibald:1932:IAS

[Arc32f] R. C. Archibald. Institute for Advanced Study — School for Mathematics. Scripta Mathematica, 1(??):167–168, ???? 1932. ISSN 0036-9713.

Archibald:1932:LCB

[Arc32g] R. C. Archibald. Lewis Carroll bibliographies. Scripta Mathematica, 1(??):172–173, ???? 1932. ISSN 0036-9713.

Archibald:1932:MP

[Arc32h] R. C. Archibald. Malfatti’s problem. Scripta Mathematica, 1(??): 170–171, ???? 1932. ISSN 0036-9713.

Archibald:1933:QDE

[Arc33] Ralph G. Archibald. Quadratic diophantine equations. Scripta Mathematica, 2(??):27–33, ???? 1933. ISSN 0036-9713.

Archibald:1934:BMVa

[Arc34a] R. C. Archibald. Bibliographia de mathematicis. V. Scripta Mathe-

matica, 2(??):181–187, ???? 1934. ISSN 0036-9713.

Archibald:1934:BMVb

[Arc34b] R. C. Archibald. Bibliographia de mathematicis. VI. Scripta Mathe-

matica, 2(??):282–292, ???? 1934. ISSN 0036-9713.

Archibald:1934:BMVc

[Arc34c] R. C. Archibald. Bibliographia de mathematicis. VII. Scripta Math-

ematica, 2(??):363–373, ???? 1934. ISSN 0036-9713.

Archibald:1934:MTa

[Arc34d] R. C. Archibald. Mathematical tables. Scripta Mathematica, 2(??):

90–93, ???? 1934. ISSN 0036-9713.

Archibald:1934:MTb

[Arc34e] R. C. Archibald. Mathematical tables. Scripta Mathematica, 2(??):

192–197, ???? 1934. ISSN 0036-9713.

Archibald:1934:MTc

[Arc34f] R. C. Archibald. Mathematical tables. Scripta Mathematica, 2(??):

297–299, ???? 1934. ISSN 0036-9713.

Archibald:1934:MTd

[Arc34g] R. C. Archibald. Mathematical tables. Scripta Mathematica, 2(??):

378–380, ???? 1934. ISSN 0036-9713.

Archibald:1934:RSC

[Arc34h] R. C. Archibald. Robert Simson and Cantor’s Vorlesungen. Scripta

Mathematica, 2(??):73–75, ???? 1934. ISSN 0036-9713.

Archibald:1934:QDE

[Arc34i] Ralph G. Archibald. Quadratic diophantine equations. Scripta Math-

ematica, 2(??):27–33, ???? 1934. ISSN 0036-9713.

Archibald:1935:BMV

[Arc35a] R. C. Archibald. Bibliographia de mathematicis. VIII. Scripta Mathematica, 3(??):83–91, ???? 1935. ISSN 0036-9713.

Archibald:1935:BMX

[Arc35b] R. C. Archibald. Bibliographia de mathematicis X: Charles Babbage.

Scripta Mathematica, 3(??):266–267, ???? 1935. ISSN 0036-9713.

Archibald:1935:BGB

[Arc35c] R. C. Archibald. Bibliography of George Berkeley. Scripta Mathe-

matica, 3(??):81–83, ???? 1935. ISSN 0036-9713.

Archibald:1935:GAM

[Arc35d] R. C. Archibald. George Abram Miller. Scripta Mathematica, 3(??):

353–354, ???? 1935. ISSN 0036-9713.

Archibald:1935:GT

[Arc35e] Ralph G. Archibald. Goldbach’s theorem. I. Scripta Mathematica, 3

(??):44–50, ???? 1935. ISSN 0036-9713.

Archibald:1935:GTI

[Arc35f] Ralph G. Archibald. Goldbach’s theorem. II. Scripta Mathematica, 3

(??):153–161, ???? 1935. ISSN 0036-9713.

Archibald:1935:MN

[Arc35g] Raymond Clare Archibald. Mersenne’s numbers. Scripta Mathemat-

ica, 3(??):112–119, ???? 1935. ISSN 0036-9713.

Archibald:1936:Ba

[Arc36a] R. C. Archibald. Biography. Scripta Mathematica, 4(??):81–93, ????

1936. ISSN 0036-9713.

Archibald:1936:Bb

[Arc36b] R. C. Archibald. Biography. Scripta Mathematica, 4(??):175–195,

???? 1936. ISSN 0036-9713.

Archibald:1936:Bc

[Arc36c] R. C. Archibald. Biography. Scripta Mathematica, 4(??):273–289,

???? 1936. ISSN 0036-9713.

Archibald:1936:Bd

[Arc36d] R. C. Archibald. Biography. Scripta Mathematica, 4(??):317–334,

???? 1936. ISSN 0036-9713.

Archibald:1936:MTI

[Arc36e] Raymond Clare Archibald. Mathematical tables IX. Scripta Mathematica, 4(??):99–104, ???? 1936. ISSN 0036-9713.

Archibald:1936:MTXa

[Arc36f] Raymond Clare Archibald. Mathematical tables X. Scripta Mathematica, 4(??):198–201, ???? 1936. ISSN 0036-9713.

Archibald:1936:MTXb

[Arc36g] Raymond Clare Archibald. Mathematical tables XI. Scripta Mathematica, 4(??):292–295, ???? 1936. ISSN 0036-9713.

Archibald:1936:MTXc

[Arc36h] Raymond Clare Archibald. Mathematical tables XII. Scripta Mathematica, 4(??):336–390, ???? 1936. ISSN 0036-9713.

Archibald:1936:WT

[Arc36i] Raymond Clare Archibald. Wolfram s table. Scripta Mathematica, 4 (1):99–100, 1936. ISSN 0036-9713.

Archibald:1940:WPS

[Arc40] Ralph G. Archibald. Waring’s problem: squares. Scripta Mathematica, 7(1–4):33–48, ???? 1940. ISSN 0036-9713.

Archibald:1945:BP

[Arc45a] Ralph G. Archibald. Bertrand’s postulate. Scripta Mathematica, 11 (??):109–120, ???? 1945. ISSN 0036-9713.

Archibald:1945:MTM

[Arc45b] Raymond Clare Archibald. Mathematical table makers — portraits, paintings, busts, monuments, bio-bibliographical notes. I. Scripta Mathematica, 11(??):213–245, ???? 1945. ISSN 0036-9713. URL

https://doi.org/10.1016/0016-0032(45)90508-4.

Archibald:1946:MTMa

[Arc46a] Raymond Clare Archibald. Mathematical table makers — portraits, paintings, busts, monuments, bio-bibliographical notes. I. Scripta Mathematica, 11(??):213–245, ???? 1946. ISSN 0036-9713.

Archibald:1946:MTMb

[Arc46b] Raymond Clare Archibald. Mathematical table makers — portraits, paintings, busts, monuments, bio-bibliographical notes. II. Scripta Mathematica, 12(??):15–51, ???? 1946. ISSN 0036-9713.

Archibald:1948:MTM

[Arc48] Raymond Clare Archibald. Mathematical Table Makers: Portraits, Paintings, Busts, Monuments: Bio-Bibliographical Notes, volume 3 of The Scripta Mathematica Studies. Scripta Mathematica, New York, NY, USA, 1948. vii + 82 pp. URL https:

//ia802706.us.archive.org/6/items/mathematicaltabl00arch/
mathematicaltabl00arch.pdf.

Archibald:1948:JSG

[AS48] Raymond Clare Archibald and Marion Elizabeth Stark. Jacob Steiner’s geometrical constructions with a ruler, given a fixed circle with its center. Scripta Mathematica, 14(??):187–264, ???? 1948. ISSN 0036-9713.

Bankoff:1955:GA

[Ban55] Leon Bankoff. The golden arbelos. Scripta Mathematica, 21(??):70– 76, ???? 1955. ISSN 0036-9713.

Barbour:1940:ML

[Bar40] M. Murray Barbour. Musical logarithms. Scripta Mathematica, 7 (??):21–31, ???? 1940. ISSN 0036-9713.

Baravalle:1948:TCIa

[Bar48a] H. V. Baravalle. Transformation of curves by inversion. Scripta Mathematica, 14(??):113–125, ???? 1948. ISSN 0036-9713.

Baravalle:1948:TCIb

[Bar48b] Herman V. Baravalle. Transformation of curves by inversion. Scripta Mathematica, 14(??):266–272, ???? 1948. ISSN 0036-9713.

Baten:1938:RIP

[Bat38] W. D. Baten. Recent interests in probability and mathematical statistics. Scripta Mathematica, 5(??):165–170, ???? 1938. ISSN 0036-9713.

Bateman:1944:HWD

[Bat44] H. Bateman. Hamilton’s work in dynamics and its influence on modern thought. Scripta Mathematica, 10(??):51–63, ???? 1944. ISSN 0036-9713.

Bergman:1961:SRB

[BB61] S. Bergman and R. Bojanić. On some relations between integral operators. Scripta Mathematica, 25(??):317–321, ???? 1961. ISSN 0036-9713.

Briggs:1955:SIS

[BCKM55] W. E. Briggs, Sarvadaman Chowla, A. J. Kempner, and W. E. Mientka. On some infinite series. Scripta Mathematica, 21(??):28–30, ???? 1955. ISSN 0036-9713.

Beard:1950:GSF

[Bea50] Robert S. Beard. The golden section and Fibonacci numbers. Scripta Mathematica, 16(??):116–119, ???? 1950. ISSN 0036-9713.

Beard:1955:VAP

[Bea55] Robert S. Beard. A variation of the Apollonius problem. Scripta Mathematica, 21(??):46–48, ???? 1955. ISSN 0036-9713.

Beeger:1950:CNW

[Bee50] N. G. W. H. Beeger. On composite numbers n for which a^{n−1} ≡ (mod n) for every a prime to n. Scripta Mathematica, 16(??):133– 135, ???? 1950. ISSN 0036-9713.

Bell:1938:FSWa

[Bel38a] E. T. Bell. Father and son Wolfgang and Johann Bolyai. Scripta Mathematica, 5(??):37–44, ???? 1938. ISSN 0036-9713.

Bell:1938:FSWb

[Bel38b] E. T. Bell. Father and son Wolfgang and Johann Bolyai. Scripta Mathematica, 5(??):95–100, ???? 1938. ISSN 0036-9713.

Bell:1943:SS

[Bel43] Eric Temple Bell. Sixes and sevens. Scripta Mathematica, 9(??): 209–231, ???? 1943. ISSN 0036-9713.

Bell:1944:SS

[Bel44] Eric Temple Bell. Sixes and sevens. Scripta Mathematica, 10(??): 81–147, ???? 1944. ISSN 0036-9713.

Bell:1945:SSa

[Bel45a] Eric Temple Bell. Sixes and sevens. Scripta Mathematica, 11(??): 21–50, ???? 1945. ISSN 0036-9713.

Bell:1945:SSb

[Bel45b] Eric Temple Bell. Sixes and sevens. Scripta Mathematica, 11(??): 139–171, ???? 1945. ISSN 0036-9713.

Bell:1946:SS

[Bel46] Eric Temple Bell. Sixes and sevens. Scripta Mathematica, 12(??): 53–60, ???? 1946. ISSN 0036-9713.

Bell:1947:PLT

[Bel47a] E. T. Bell. The problem of Liouville’s theorems on arithmetical quadratic forms. Scripta Mathematica, 13(??):177–185, ???? 1947. ISSN 0036-9713. URL https://doi.org/10.1080/

03081088308817515.

Bell:1947:PV

[Bel47b] E. T. Bell. A Pythagorean variation. Scripta Mathematica, 13(??): 163–168, ???? 1947. ISSN 0036-9713.

Bell:1949:WF

[Bel49] E. T. Bell. Wallis on Fermat. Scripta Mathematica, 15(??):162–163, ???? 1949. ISSN 0036-9713.

Bellman:1956:SAT

[Bel56] Richard Bellman. Some aspects of the theory of dynamic programming. Scripta Mathematica, 21(??):273–277, ???? 1956. ISSN 0036-9713.

Belkin:1958:PJG

[Bel58] Samuel Belkin. Professor Jekuthiel Ginsburg. An appreciation. Scripta Mathematica, 23(??):7–9, ???? 1958. ISSN 0036-9713.

Bernstein:1950:DSD

[Ber50] B. A. Bernstein. A dual-symmetric definition of Boolean algebra free from postulated special elements. Scripta Mathematica, 16(??):157– 160, ???? 1950. ISSN 0036-9713.

Bernhart:1954:CP

[Ber54] Arthur Bernhart. Curves of pursuit. Scripta Mathematica, 20(??): 125–141 (1955), ???? 1954. ISSN 0036-9713.

Bergman:1957:OGS

[Ber57a] Stefan Bergman. Operators generating solutions of certain partial differential equations in three variables and their properties. I. Scripta Mathematica, 23(??):143–151, ???? 1957. ISSN 0036-9713.

Bernhart:1957:CPI

[Ber57b] Arthur Bernhart. Curves of pursuit. II. Scripta Mathematica, 23(??): 49–65, ???? 1957. ISSN 0036-9713.

Bergman:1958:OGS

[Ber58a] Stefan Bergman. Operators generating solutions of certain partial differential equations in three variables and their properties. I. Scripta Mathematica, 23(??):143–151, ???? 1958. ISSN 0036-9713.

Bernhart:1958:CPI

[Ber58b] Arthur Bernhart. Curves of pursuit. II. Scripta Mathematica, 23(??): 49–65, ???? 1958. ISSN 0036-9713.

Bernhart:1959:CGP

[Ber59a] Arthur Bernhart. Curves of general pursuit. Scripta Mathematica, 24(??):189–206, ???? 1959. ISSN 0036-9713. URL https://web.

stanford.edu/~jugander/rare/Bernhart1959_CurvesOfGeneralPursuit.
pdf.

Bernhart:1959:PP

[Ber59b] Arthur Bernhart. Polygons of pursuit. Scripta Mathematica, 24(1): 23–50, ???? 1959. ISSN 0036-9713.

Bergman:1961:OGS

[Ber61] Stefan Bergman. Operators generating solutions of certain differential equations in three variables and their properties. Scripta Mathematica, 26(??):5–31, ???? 1961. ISSN 0036-9713.

Beatty:1938:TAFa

[BF38a] S. Beatty and M. J. Fisher. The theory of algebraic functions based on a study of fundamental exponents. Scripta Mathematica, 5(??): 187–194, ???? 1938. ISSN 0036-9713.

Beatty:1938:TAFb

[BF38b] S. Beatty and M. J. Fisher. The theory of algebraic functions based on a study of fundamental exponents. Scripta Mathematica, 5(??): 239–245, ???? 1938. ISSN 0036-9713.

Blumenthal:1973:SSF

[BF73] Leonard M. Blumenthal and Raymond W. Freese. Solution of a spaceform problem. Scripta Mathematica, 29(3–4):273–277, ???? 1973. ISSN 0036-9713.

Blij:1967:VCA

[Bli67] F. Blij. Van der Combinatorial Aspects of the Hexagrams in the Chinese Book of Changes. Scripta Mathematica, 28(1):??, ???? 1967. ISSN 0036-9713.

Block:1953:CFS

[Blo53a] D. Block. Curiosum # 330: Fibonacci summations. Scripta Mathematica, 19(2–3):191, 1953. ISSN 0036-9713.

Block:1953:SFI

[Blo53b] Daniel Block. Symbolic formulae involving Fibonacci numbers. Scripta Mathematica, 18(??):306–307, ???? 1953. ISSN 0036-9713.

Block:1955:SPC

[Blo55] Daniel Block. Some properties of the convergents of \sqrt{2}. Scripta Mathematica, 21(??):208–213, ???? 1955. ISSN 0036-9713.

Butchart:1953:NCP

[BM53] J. H. Butchart and Leo Moser. No calculus, please. Scripta Mathematica, 18(??):221–236, ???? 1953. ISSN 0036-9713.

Bochner:1964:AMC

[Boc64] Salomon Bochner. Analytic measures on compact Bohr groups. Scripta Mathematica, 27(??):5–26, ???? 1964. ISSN 0036-9713.

Bowes:1932:DS

[Bow32] Julian Bowes. Dynamic symmetry. Scripta Mathematica, 1(??):236– 244, ???? 1932. ISSN 0036-9713.

Boyer:1941:CLD

[Boy41] Carl B. Boyer. Cavalieri, limits and discarded infinitesimals. Scripta Mathematica, 8(??):79–91, ???? 1941. ISSN 0036-9713.

Boyer:1943:PFS

[Boy43] Carl B. Boyer. Pascal’s formula for the sums of powers of the integers. Scripta Mathematica, 9(??):237–244, ???? 1943. ISSN 0036-9713.

Boyer:1945:EGS

[Boy45] Carl B. Boyer. Early graphical solutions of polynomial equations. Scripta Mathematica, 11(??):5–19, ???? 1945. ISSN 0036-9713.

Boyer:1946:PEF

[Boy46] Carl B. Boyer. Proportion, equation, function: three steps in the development of a concept. Scripta Mathematica, 12(1):5–13, ???? 1946. ISSN 0036-9713.

Boyer:1947:CGF

[Boy47] Carl B. Boyer. Cartesian geometry from Fermat to Lacroix. Scripta Mathematica, 13(??):133–153, ???? 1947. ISSN 0036-9713.

Boyd:1949:LD

[Boy49] Rutherford Boyd. In a laboratory of design. Scripta Mathematica, 15 (??):183–192, ???? 1949. ISSN 0036-9713.

Boyer:1950:NEa

[Boy50a] Carl B. Boyer. From Newton to Euler. Scripta Mathematica, 16(??): 141–157, ???? 1950. ISSN 0036-9713.

Boyer:1950:NEb

[Boy50b] Carl B. Boyer. From Newton to Euler. Scripta Mathematica, 16(??): 221–258, ???? 1950. ISSN 0036-9713.

Boyer:1951:GA

[Boy51a] Carl B. Boyer. The Golden Age. I. Scripta Mathematica, 17(??): 32–54, ???? 1951. ISSN 0036-9713.

Boyer:1951:GAI

[Boy51b] Carl B. Boyer. The Golden Age. II. Scripta Mathematica, 17(??): 209–230, ???? 1951. ISSN 0036-9713.

Boyer:1952:FD

[Boy52] Carl B. Boyer. Fermat and Descartes. Scripta Mathematica, 18(??): 189–217 (1953), ???? 1952. ISSN 0036-9713.

Boyer:1953:ECA

[Boy53a] Carl B. Boyer. Early contributions to analytic geometry. Scripta Mathematica, 19(??):230–238 (1954), ???? 1953. ISSN 0036-9713.

Boyer:1953:FD

[Boy53b] Carl B. Boyer. Fermat and Descartes. Scripta Mathematica, 18(??): 189–217, ???? 1953. ISSN 0036-9713.

Boyer:1954:AGA

[Boy54a] Carl B. Boyer. Analytic geometry in the Alexandrian age. I. Scripta Mathematica, 20(??):30–36, ???? 1954. ISSN 0036-9713.

Boyer:1954:AGAb

[Boy54b] Carl B. Boyer. Analytic geometry in the Alexandrian age. II. Scripta Mathematica, 20(??):143–154, ???? 1954. ISSN 0036-9713.

Boyer:1954:ECA

[Boy54c] Carl B. Boyer. Early contributions to analytic geometry. Scripta Mathematica, 19(??):97–108, ???? 1954. ISSN 0036-9713.

Boyer:1955:AGA

[Boy55a] Carl B. Boyer. Analytic geometry in the Alexandrian age. Scripta Mathematica, 20:143–154, 1955. ISSN 0036-9713.

Boyer:1955:PCA

[Boy55b] Carl B. Boyer. Post-Cartesian analytic geometry. Scripta Mathematica, 21(??):101–135, ???? 1955. ISSN 0036-9713.

Boyer:1958:JG

[Boy58] Carl B. Boyer. Jekuthiel Ginsburg (1889–1957). Isis, 49(3):335–336, September 1958. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

         (electronic). URL http://www.jstor.org/stable/226941.

Boyer:1960:MFR

[Boy60] Carl B. Boyer. Mathematicians of the French Revolution. Scripta Mathematica, 25(??):11–31, ???? 1960. ISSN 0036-9713.

Boyer:1963:PMM

[Boy63] Carl B. Boyer. Pascal: The man and the mathematician. Scripta Mathematica, 26(??):283–307, ???? 1963. ISSN 0036-9713.

Boyer:1966:MNC

[Boy66] C. B. Boyer. Miscellaneous notes. Colin MacLaurin and Cramer’s rule. Scripta Mathematica, 27(4):377–379, ???? 1966. ISSN 0036-9713.

Bradley:1934:MR

[Bra34] D. Bradley. A manuscript Rechenbüchlein of 1727. Scripta Mathematica, 2(??):235–242, ???? 1934. ISSN 0036-9713.

Bragdon:1936:FMS

[Bra36] C. Bragdon. The Franklin 16×16 magic square. Scripta Mathematica, 4(??):158–159, ???? 1936. ISSN 0036-9713.

Bradley:1938:PGA

[Bra38] A. D. Bradley. Pennsylvania German arithmetical books. Scripta Mathematica, 5(??):45–51, ???? 1938. ISSN 0036-9713.

Bradley:1940:MMS

[Bra40] A. Day Bradley. The mathematical manuscripts in the Schwenkfelder historical library. Scripta Mathematica, 7(??):49–58, ???? 1940. ISSN 0036-9713.

Bradley:1947:GCT

[Bra47] A. Day Bradley. Great circle trigonometry. Scripta Mathematica, 13 (??):64–70, ???? 1947. ISSN 0036-9713.

Bradley:1949:PVT

[Bra49] A. Day Bradley. Pieter Venema; teacher, textbook author, and freethinker. Scripta Mathematica, 15(??):13–16, ???? 1949. ISSN 0036-9713.

Brasch:1952:SNC

[Bra52] Frederick E. Brasch. A survey of the number of copies of Newton’s Principia in the United States, Canada and Mexico. Scripta Mathematica, 18(??):53–67, ???? 1952. ISSN 0036-9713.

Bradley:1955:DWG

[Bra55] A. Day Bradley. The day of the work for Gregorian dates. Scripta Mathematica, 21(??):82–87, ???? 1955. ISSN 0036-9713.

Brenman:1955:TD

[Bre55] Edwin Brenman. Testing for divisibility. Scripta Mathematica, 21 (??):88–90, ???? 1955. ISSN 0036-9713.

Bridgman:1934:PSRa

[Bri34a] P. W. Bridgman. A physicist’s second reaction to Mengenlehre. Scripta Mathematica, 2(??):101–117, ???? 1934. ISSN 0036-9713.

Bridgman:1934:PSRb

[Bri34b] P. W. Bridgman. A physicist’s second reaction to Mengenlehre. Scripta Mathematica, 2(??):224–234, ???? 1934. ISSN 0036-9713.

Britton:1960:E

[Bri60] C. E. Britton. The equation x^{3} + y^{3} = u^{3} + v^{3} = N. Scripta Mathematica, 25(??):165–166, ???? 1960. ISSN 0036-9713.

Bronstein:1939:SDE

[Bro39] J. Bronstein. A solution of the diophantine equation x^{2}−kxy+y^{2} = 1. Scripta Mathematica, 6(??):60, ???? 1939. ISSN 0036-9713.

Brown:1954:MN

[Bro54] Alan L. Brown. Multiperfect numbers. Scripta Mathematica, 20(??): 103–106, ???? 1954. ISSN 0036-9713.

Brown:1973:MST

[Bro73] Robert Bruce Brown. On maximal subalgebras of type G_{2} in Lie algebras of type E_{6}. Scripta Mathematica, 29(3–4):443–447, ???? 1973. ISSN 0036-9713.

Bruce:1938:SMAa

[Bru38a] R. E. Bruce. Sicily, and the march of ancient mathematics and science to the modern world. I. Scripta Mathematica, 5(??):117–121, ???? 1938. ISSN 0036-9713.

Bruce:1938:SMAb

[Bru38b] R. E. Bruce. Sicily, and the march of ancient mathematics and science to the modern world. II. Scripta Mathematica, 5(??):181–185, ???? 1938. ISSN 0036-9713.

Bruce:1938:SMAc

[Bru38c] R. E. Bruce. Sicily, and the march of ancient mathematics and science to the modern world. III. Scripta Mathematica, 5(??):245–250, ???? 1938. ISSN 0036-9713.

Candido:1951:RBF

[Can51] G. Candido. A relationship between the fourth powers of the terms of the Fibonacci series. Scripta Mathematica, 17(3-4):230, 1951. ISSN 0036-9713.

Carlitz:1952:NFS

[Car52] L. Carlitz. Note on a formula of Szily. Scripta Mathematica, 18(??): 249–253 (1953), ???? 1952. ISSN 0036-9713.

Carlitz:1953:NFS

[Car53] Leonard Carlitz. Note on a formula of Szily. Scripta Mathematica, 18(??):249–253, ???? 1953. ISSN 0036-9713.

Carlitz:1954:CPM

[Car54] Leonard Carlitz. Congruence properties of the ménage polynomials. Scripta Mathematica, 20(??):51–57, ???? 1954. ISSN 0036-9713.

Carlitz:1956:NBN

[Car56] L. Carlitz. A note on Bernoulli numbers of higher order. Scripta Mathematica, 22(??):217–221 (1957), ???? 1956. ISSN 0036-9713.

Carlitz:1957:NBN

[Car57] Leonard Carlitz. A note on Bernoulli numbers of higher order. Scripta Mathematica, 22(??):217–221, ???? 1957. ISSN 0036-9713.

Carlitz:1961:NBE

[Car61] Leonard Carlitz. A note on Bernoulli and Euler polynomials of the second kind. Scripta Mathematica, 25(??):323–330, ???? 1961. ISSN 0036-9713.

Carlitz:1963:BAM

[Car63] Leonard Carlitz. A basic analog of the multinomial theorem. Scripta Mathematica, 26(??):317–321, ???? 1963. ISSN 0036-9713.

Carlitz:1964:ECB

[Car64] L. Carlitz. An extension of the congruences of Bauer and Lubleski. Scripta Mathematica, 27(??):213–222, ???? 1964. ISSN 0036-9713.

Carlitz:1965:ECB

[Car65] Leonard Carlitz. An extension of the congruences of Bauer and Lubelski. Scripta Mathematica, 27(??):213–222, ???? 1965. ISSN 0036-9713.

Cauman:1968:IP

[Cau68] Leigh S. Cauman. On indirect proof. Scripta Mathematica, 28(2): 101–115, ???? 1968. ISSN 0036-9713.

Chern:1973:MVF

[Che73] Shiing-Shen Shen Chern. Meromorphic vector fields and characteristic numbers. Scripta Mathematica, 29(3–4):243–251, ???? 1973. ISSN

0036-9713. URL https://doi.org/10.1016/0370-1573(77)90076-

Clarke:1934:AEF

[Cla34] F. M. Clarke. On an anonymous essay on fluxions, published in 1741. Scripta Mathematica, 2(??):155–160, ???? 1934. ISSN 0036-9713.

Cohen:1957:GEF

[Coh57] Eckford Cohen. Generalizations of the Euler φ-function. Scripta Mathematica, 23(??):157–161, ???? 1957. ISSN 0036-9713.

Cohen:1958:GEF

[Coh58] Eckford Cohen. Generalizations of the Euler φ-function. Scripta Mathematica, 23(??):157–161, ???? 1958. ISSN 0036-9713.

Cohen:1960:EAF

[Coh60] Eckford Cohen. The elementary arithmetical functions. Scripta Mathematica, 25(??):221–228, ???? 1960. ISSN 0036-9713.

Cohen:1963:ANI

[Coh63] Eckford Cohen. Arithmetical notes. II. An estimate of Erdős and Szekeres. Scripta Mathematica, 26(??):353–356, ???? 1963. ISSN 0036-9713.

Cohen:1964:NCA

[Coh64] Eckford Cohen. Note on a class of affine transformations. Scripta Mathematica, 27(??):283–284, ???? 1964. ISSN 0036-9713.

Cohen:1965:NCA

[Coh65] E. Cohen. Note on a class of affine transformations. Scripta Mathematica, 27(??):283–284, ???? 1965. ISSN 0036-9713.

Cohen:1973:MIM

[Coh73] Leon W. Cohen. Measure and integration in the manner of Borel. Scripta Mathematica, 29(3–4):417–435, ???? 1973. ISSN 0036-9713.

Coolidge:1951:SFM

[Coo51] Julian L. Coolidge. Six female mathematicians. Scripta Mathematica, 17(??):20–31, ???? 1951. ISSN 0036-9713.

Court:1947:MA

[Cou47] Nathan Altshiller Court. Mathematical asides. Scripta Mathematica, 13(??):79–85, ???? 1947. ISSN 0036-9713.

Court:1948:TA

[Cou48] Nathan Altshiller Court. The tetrahedron and its altitudes. Scripta Mathematica, 14(??):85–97, ???? 1948. ISSN 0036-9713.

Court:1951:IEPa

[Cou51a] Nathan Altshiller Court. Imaginary elements in pure geometry — what they are and what they are not. Scripta Mathematica, 17(??): 55–64, ???? 1951. ISSN 0036-9713.

Court:1951:IEPb

[Cou51b] Nathan Altshiller Court. Imaginary elements in pure geometry — what they are and what they are not. Scripta Mathematica, 17(??): 190–201, ???? 1951. ISSN 0036-9713.

Court:1954:CPa

[Cou54a] Nathan Altshiller Court. Castillon’s problem. Scripta Mathematica, 20(??):118–120, ???? 1954. ISSN 0036-9713.

Court:1954:CPb

[Cou54b] Nathan Altshiller Court. Castillon’s problem. Scripta Mathematica, 20(??):232–235, ???? 1954. ISSN 0036-9713.

Court:1954:DHSa

[Cou54c] Nathan Altshiller Court. Desargues and his strange theorem. Scripta Mathematica, 20(??):5–13, ???? 1954. ISSN 0036-9713.

Court:1954:DHSb

[Cou54d] Nathan Altshiller Court. Desargues and his strange theorem. II. Scripta Mathematica, 20(??):155–164 (1955), ???? 1954. ISSN 0036-9713.

Court:1956:MPR

[Cou56] Nathan Altshiller Court. The might and plight of reasoning. Scripta Mathematica, 22(??):45–52, ???? 1956. ISSN 0036-9713.

Court:1957:CC

[Cou57] Nathan Altshiller Court. The Cevian chain. Scripta Mathematica, 22 (??):193–202, ???? 1957. ISSN 0036-9713.

Coxeter:1936:PSA

[Cox36] H. S. M. (Harold Scott Macdonald) Coxeter. The partition of a sphere according to the icosahedral group. Scripta Mathematica, 4(??):156– 157, ???? 1936. ISSN 0036-9713.

Coxeter:1939:RSS

[Cox39] H. S. M. (Harold Scott Macdonald) Coxeter. The regular sponges, or skew polyhedra. Scripta Mathematica, 6(??):240–244, ???? 1939. ISSN 0036-9713.

Coxeter:1952:IRS

[Cox52a] H. S. M. (Harold Scott Macdonald) Coxeter. Interlocked rings of spheres. Scripta Mathematica, 18(??):113–121, ???? 1952. ISSN 0036-9713.

Coxeter:1952:RBU

[Cox52b] H. S. M. (Harold Scott Macdonald) Coxeter. Rouse Ball’s unpublished notes on three fours. Scripta Mathematica, 18(??):85–86, ???? 1952. ISSN 0036-9713.

Coxeter:1953:GSP

[Cox53] H. S. M. (Harold Scott Macdonald) Coxeter. The golden section, Phyllotaxis and Wythoff’s game. Scripta Mathematica, 19(??):135– 143, ???? 1953. ISSN 0036-9713.

Coxeter:1954:SUP

[Cox54] H. S. M. (Harold Scott Macdonald) Coxeter. Six uniform polyhedra. Scripta Mathematica, 20(??):227, ???? 1954. ISSN 0036-9713.

Coxeter:1955:AP

[Cox55] H. S. M. (Harold Scott Macdonald) Coxeter. The affine plane. Scripta Mathematica, 21(??):5–14, ???? 1955. ISSN 0036-9713.

Coxeter:1956:HT

[Cox56] H. S. M. (Harold Scott Macdonald) Coxeter. Hyperbolic triangles. Scripta Mathematica, 22(??):5–13, ???? 1956. ISSN 0036-9713.

Coxeter:1957:MCP

[Cox57] H. S. M. (Harold Scott Macdonald) Coxeter. Map-coloring problems. Scripta Mathematica, 23(??):11–25, ???? 1957. ISSN 0036-9713.

Coxeter:1958:MCP

[Cox58] H. S. M. (Harold Scott Macdonald) Coxeter. Map-coloring problems. Scripta Mathematica, 23(??):11–25, ???? 1958. ISSN 0036-9713.

Cordes:1973:SOI

[CP73] Craig Cordes and Gordon Pall. Sequences of orthogonal integral vectors with a fixed norm. Scripta Mathematica, 29(3–4):293–295, ???? 1973. ISSN 0036-9713.

Charzynski:1960:GPB

[CS60] Z. Charzyński and M. Schiffer. A geometric proof of the Bieberbach conjecture for the fourth coefficient. Scripta Mathematica, 25(??): 173–181, ???? 1960. ISSN 0036-9713.

Cowan:1947:FCD

[CW47] Russel W. Cowan and Bernard C. Weber. Fermat’s contribution to the development of the differential calculus. Scripta Mathematica, 13 (??):123–127, ???? 1947. ISSN 0036-9713.

Davidson:1936:LBA

[Dav36] Israel Davidson. Levi ben Abraham ben Hayyim. A mathematician of the XIII-th century. Scripta Mathematica, 4(??):57–65, ???? 1936. ISSN 0036-9713.

Davis:1959:SNL

[Dav59] Chandler Davis. The set of non-linearity of a convex piecewise-linear function. Scripta Mathematica, 24(??):219–228, ???? 1959. ISSN 0036-9713.

Davis:1966:GBS

[Dav66] Constance Davis. A group of binary substitutions transitive on integral solutions of px^{2} − y^{2} − z^{2} = 0. Scripta Mathematica, 27(??): 331–337, ???? 1966. ISSN 0036-9713.

DeCicco:1945:SPF

[De 45] John De Cicco. Survey of polygenic functions. Scripta Mathematica, 11(??):51–56, ???? 1945. ISSN 0036-9713.

DeCicco:1947:EET

[De 47] John De Cicco. An extension of Euler’s theorem of homogeneous functions. Scripta Mathematica, 13(??):48–52, ???? 1947. ISSN 0036-9713.

DeVries:1948:HAG

[De 48] H. De Vries. How analytic geometry became a science. Scripta Mathematica, 14(??):5–15, ???? 1948. ISSN 0036-9713.

Delange:1957:TE

[Del57] Hubert Delange. On a theorem of Erdős. Scripta Mathematica, 23

(??):109–116, ???? 1957. ISSN 0036-9713.

Delange:1958:TE

[Del58] Hubert Delange. On a theorem of Erdős. Scripta Mathematica, 23

(??):109–116, ???? 1958. ISSN 0036-9713.

Delange:1963:CMA

[Del63] Hubert Delange. On a class of multiplicative arithmetical functions.

Scripta Mathematica, 26(??):121–141, ???? 1963. ISSN 0036-9713.

Dijksterhuis:1950:WHS

[Dij50] E. J. Dijksterhuis. Works on history of science published in The Netherlands in the years 1930–1947. Scripta Mathematica, 16(??): 43–59, ???? 1950. ISSN 0036-9713.

Divinsky:1973:DBR

[Div73] N. Divinsky. Duality between radical and semisimple classes of associative rings. Scripta Mathematica, 29(3–4):409–416, ???? 1973. ISSN 0036-9713.

Dorwart:1948:NFE

[Dor48] H. L. Dorwart. A normal form for the equation of the straight line in space. Scripta Mathematica, 14(??):181–185, ???? 1948. ISSN 0036-9713.

Dorwart:1950:EAU

[Dor50] H. L. Dorwart. An early American unpublished work in mathematics.

Scripta Mathematica, 16(??):181–185, ???? 1950. ISSN 0036-9713.

Douglas:1938:PP

[Dou38] J. Douglas. The problem of Plateau. Scripta Mathematica, 5(??):

159–164, ???? 1938. ISSN 0036-9713.

Douglas:1960:TSP

[Dou60] Jesse Douglas. A theorem on skew pentagons. Scripta Mathematica,

25(??):5–9, ???? 1960. ISSN 0036-9713.

Dresden:1950:SIO

[Dre50] Arnold Dresden. The Schwarz inequality and the order of operations. Scripta Mathematica, 16(??):259–260, ???? 1950. ISSN 0036-9713.

Dresden:1953:CI

[Dre53] Arnold Dresden. Complete independence. Scripta Mathematica, 19 (??):205–206, ???? 1953. ISSN 0036-9713.

Dunnington:1935:LS

[Dun35] G. Waldo Dunnington. Ludwig Schlesinger. Scripta Mathematica, 3 (??):67–68, ???? 1935. ISSN 0036-9713.

Dunnington:1954:NGT

[Dun54] G. Waldo Dunnington. Note on Gauss’ triangulations. Scripta Mathematica, 20(??):108–109, ???? 1954. ISSN 0036-9713.

Dutka:1953:STP

[Düt53] Jacques Dütka. Spinoza and the theory of probability. Scripta Mathematica, 19(1):24–33, ???? 1953. ISSN 0036-9713.

Easton:1966:TE

[Eas66] Joy B. Easton. A Tudor Euclid. Scripta Mathematica, 27(4):339–355, ???? 1966. ISSN 0036-9713.

Edrei:1956:CPCa

[Edr56a] Albert Edrei. On a conjecture of Pólya concerning the zeros of successive derivatives. Scripta Mathematica, 22:31–44, 1956. ISSN 0036-9713.

Edrei:1956:CPCb

[Edr56b] Albert Edrei. On a conjecture of Pólya concerning the zeros of successive derivatives. Scripta Mathematica, 22:106–121, 1956. ISSN 0036-9713.

Edrei:1957:GDT

[Edr57] Albert Edrei. Gap and density theorems for entire functions. Scripta Mathematica, 23:117–141, 1957. ISSN 0036-9713.

Edrei:1958:GDT

[Edr58] Albert Edrei. Gap and density theorems for entire functions. Scripta Mathematica, 23(??):117–141, ???? 1958. ISSN 0036-9713.

Eisele:1950:LGS

[Eis50] Carolyn Eisele. Lao Genevra Simons. Scripta Mathematica, 16(??): 22–30, ???? 1950. ISSN 0036-9713.

Eisele:1951:LAT

[Eis51] Carolyn Eisele. The Liber Abaci through the eyes of Charles S. Peirce. Scripta Mathematica, 17(??):236–259, ???? 1951. ISSN 0036-9713.

Eisele:1959:CPN

[Eis59] Carolyn Eisele. Charles S. Peirce nineteenth century man of science. Scripta Mathematica, 24(??):304–324, ???? 1959. ISSN 0036-9713.

Eisele:1960:CPN

[Eis60] Carolyn Eisele. Charles S. Peirce, nineteenth century man of science. Scripta Mathematica, 24(??):305–324, ???? 1960. ISSN 0036-9713.

Erdos:1946:SPM

[EK46] Paul Erdős and Irving Kaplansky. Sequences of plus and minus. Scripta Mathematica, 12(??):73–75, ???? 1946. ISSN 0036-9713.

Emch:1935:LDGa

[Emc35a] Arnold F. Emch. The Logica demonstrativa of Girolamo Saccheri. I. Scripta Mathematica, 3(??):51–60, ???? 1935. ISSN 0036-9713.

Emch:1935:LDGb

[Emc35b] Arnold F. Emch. The Logica demonstrativa of Girolamo Saccheri. II. Scripta Mathematica, 3(??):143–152, ???? 1935. ISSN 0036-9713.

Emch:1935:LDGc

[Emc35c] Arnold F. Emch. The Logica demonstrativa of Girolamo Saccheri. III. Scripta Mathematica, 3(??):221–233, ???? 1935. ISSN 0036-9713.

Emch:1950:RPP

[Emc50] Arnold Emch. Rare problems in plane geometry. Scripta Mathematica, 16(??):61–66, ???? 1950. ISSN 0036-9713.

Emch:1952:SPH

[Emc52] Arnold Emch. Some points in the history of the catacaustica and a clarifying representation of the curve. Scripta Mathematica, 18(??): 31–35, ???? 1952. ISSN 0036-9713.

Endrei:1956:CPCa

[End56a] Albert Endrei. On a conjecture of Pólya concerning the zeros of successive derivatives. Scripta Mathematica, 22(??):31–44, ???? 1956. ISSN 0036-9713.

Endrei:1956:CPCb

[End56b] Albert Endrei. On a conjecture of Polya concerning the zeros of successive derivatives. Scripta Mathematica, 22(??):106–121, ???? 1956. ISSN 0036-9713.

Endrei:1957:GDT

[End57] Albert Endrei. Gap and density theorems for entire functions. Scripta Mathematica, 23(??):117–141, ???? 1957. ISSN 0036-9713.

Escott:1946:AN

[Esc46] Edward Brind Escott. Amicable numbers. Scripta Mathematica, 12 (??):61–72, ???? 1946. ISSN 0036-9713.

Ettlinger:1956:RME

[Ett56] H. J. Ettlinger. The role of mathematics in every-day life. Scripta Mathematica, 22(??):132–136, ???? 1956. ISSN 0036-9713.

Eves:1959:PL

[Eve59] Howard Eves. Philo’s line. Scripta Mathematica, 24(??):141–148, ????

1959. ISSN 0036-9713.

Feld:1940:CAA

[Fel40] J. M. Feld. Counterpoints aud associated cubic curves. Scripta Mathematica, 7(??):113–118, ???? 1940. ISSN 0036-9713.

Freitag:1956:NPT

[FF56a] Herta Taussig Freitag and Arthur H. Freitag. Neo-Pythagorean triangles. Scripta Mathematica, 22(??):122–131, ???? 1956. ISSN 0036-9713.

Freitag:1956:NT

[FF56b] Herta Taussig Freitag and Arthur H. Freitag. Neopythagorean triangles. Scripta Mathematica, 22:122–131, 1956. ISSN 0036-9713.

Franqui:1954:NMP

[FG54] Benito Franqui and Mariano García. 57 new multiply perfect numbers. Scripta Mathematica, 20(??):169–171 (1955), ???? 1954. ISSN 0036-9713.

Finn:1963:PDE

[Fin63] Robert Finn. On partial differential equations (whose solutions admit no isolated singularities). Scripta Mathematica, 26(??):107–115, ???? 1963. ISSN 0036-9713.

Fadiman:1938:TTS

[FKCR38] Clifton Fadiman, E. J. Kahn, Jr., Charles Cooke, and Harold Ross. The talk of the town: Scripta Mathematica. The New Yorker, ?? (??):7, July 30, 1938. ISSN 0028-792X. URL https://archives.

newyorker.com/newyorker/1938-07-30/flipbook/006;   https:
//www.newyorker.com/magazine/1938/07/30/scripta-mathematica.

Fleisher:1943:SMC

[Fle43] David Fleisher. Scripta Mathematica: a contribution to American culture. The Synagogue Light, 10(6):3–4, 11–12, February 1943.

Fosdick:1943:HM

[Fos43] R. B. Fosdick. Hitler and mathematics. Scripta Mathematica, 9(??): 120–122, ???? 1943. ISSN 0036-9713.

Fraenkel:1932:MPF

[Fra32] Abraham Adolf Fraenkel. On modern problems in the foundations of mathematics. Scripta Mathematica, 1(??):222–227, ???? 1932. ISSN 0036-9713.

Fraenkel:1933:MPF

[Fra33] Abraham Adolf Fraenkel. On modern problems in the foundations of mathematics. Scripta Mathematica, 1(??):222–227, ???? 1933. ISSN 0036-9713.

Fraenkel:1938:AL

[Fra38] Abraham Adolf Fraenkel. Alfred Loewy (1873–1935). Scripta Mathematica, 5(??):17–22, ???? 1938. ISSN 0036-9713.

Franklin:1939:FCPa

[Fra39a] Philip Franklin. The four color problem. Scripta Mathematica, 6(??): 149–156, ???? 1939. ISSN 0036-9713.

Franklin:1939:FCPb

[Fra39b] Philip Franklin. The four color problem. Scripta Mathematica, 6(??): 197–210, ???? 1939. ISSN 0036-9713.

Fraenkel:1940:NNO

[Fra40] Abraham Adolf Fraenkel. Natural numbers as ordinals. Scripta Mathematica, 7(1–4):9–20, ???? 1940. ISSN 0036-9713.

Fraenkel:1943:PMMa

[Fra43a] Abraham Adolf Fraenkel. Problems and methods in modern mathematics. Scripta Mathematica, 9(??):81–84, ???? 1943. ISSN 0036-9713.

Fraenkel:1943:PMMb

[Fra43b] Abraham Adolf Fraenkel. Problems and methods in modern mathematics. Scripta Mathematica, 9(??):162–168, ???? 1943. ISSN 0036-9713.

Fraenkel:1943:PMMc

[Fra43c] Abraham Adolf Fraenkel. Problems and methods in modern mathematics. Scripta Mathematica, 9(??):245–255, ???? 1943. ISSN 0036-9713.

Fraenkel:1947:RCA

[Fra47] Abraham Adolf Fraenkel. The recent controversies about the foundation of mathematics. Scripta Mathematica, 13(1–2):17–36, ???? 1947. ISSN 0036-9713.

Fraenkel:1949:PMM

[Fra49] Abraham Adolf Fraenkel. Problems and methods in modern mathematics. II. Scripta Mathematica, 15(??):169–182, ???? 1949. ISSN 0036-9713.

Fraenkel:1951:CPE

[Fra51] Abraham Adolf Fraenkel. On the crisis of the principle of the excluded middle. Scripta Mathematica, 17(??):5–16, ???? 1951. ISSN 0036-9713.

Fraenkel:1960:JMA

[Fra60] Abraham Adolf Fraenkel. Jewish mathematics and astronomy. Scripta Mathematica, 25(??):33–47, ???? 1960. ISSN 0036-9713.

Franklin:1961:EOC

[Fra61] S. P. Franklin. Eudoxus, Omar, and continued fractions. Scripta Mathematica, 25(??):353–355, ???? 1961. ISSN 0036-9713.

Frank:1973:RSL

[Fra73] Marguerite Frank. Restrictedness and simplicity in Lie algebras of characteristic p. Scripta Mathematica, 29(3–4):395–407, ???? 1973. ISSN 0036-9713.

Freedman:1948:FNG

[Fre48] Benedict Freedman. The four number game. Scripta Mathematica, 14(??):35–47, ???? 1948. ISSN 0036-9713.

Freedman:1953:SHS

[Fre53] Benedict Freedman. On a symmetric homogeneous sum. Scripta Mathematica, 19(??):185–188, ???? 1953. ISSN 0036-9713.

Fry:1934:MTR

[Fry34] T. C. Fry. A mathematical theory of rational inference. Scripta Mathematica, 2(??):205–221, ???? 1934. ISSN 0036-9713.

Fryde:1964:WSM

[Fry64] Matthew M. Fryde. Wacław Sierpiński — mathematician. Scripta Mathematica, 27(??):105–111, ???? 1964. ISSN 0036-9713.

Furstenberg:1954:EMT

[Fur54] Harry Furstenberg. An extension of Menelaus’ theorem. Scripta Mathematica, 20(??):238–239, ???? 1954. ISSN 0036-9713.

Geppert:1935:GDA

[GA35] H. Geppert and R. C. Archibald. Gauß’s Disquisitiones arithmeticae and the French Academy of Sciences. Scripta Mathematica, 3(??): 285–286, ???? 1935. ISSN 0036-9713.

Ganguli:1932:IDI

[Gan32] Saradakanta Ganguli. On the Indian discovery of the irrational at the time of Śulva-Sûtras. Scripta Mathematica, 1(??):136–141, ???? 1932. ISSN 0036-9713.

Garver:1932:A

[Gar32] R. Garver. The Analyst, 1874–1883. Scripta Mathematica, 1(??): 247–251, ???? 1932. ISSN 0036-9713.

Garver:1935:BTM

[Gar35] Raymond Garver. Bieberbach’s trisection method. Scripta Mathematica, 3(??):251–255, ???? 1935. ISSN 0036-9713.

Gardner:1948:MCT

[Gar48] Martin Gardner. Mathematical card tricks. Scripta Mathematica, 14 (??):99–111, ???? 1948. ISSN 0036-9713. URL https://www.martin-

gardner.org/Scripta.html.

Gardner:1949:MTC

[Gar49] Martin Gardner. Mathematical tricks with common objects. Scripta Mathematica, 15(??):17–26, ???? 1949. ISSN 0036-9713. URL https:

//www.martin-gardner.org/Scripta.html.

Gardner:1951:TM

[Gar51] Martin Gardner. Topology and magic. Scripta Mathematica, 17(??): 75–83, ???? 1951. ISSN 0036-9713. URL https://www.martin-

gardner.org/Scripta.html.

Gardner:1952:MTS

[Gar52] Martin Gardner. Mathematical tricks with special equipment. Scripta Mathematica, 18(??):237–249, ???? 1952. ISSN 0036-9713. URL

https://www.martin-gardner.org/Scripta.html.

Garcia:1953:GMP

[Gar53] Mariano Garcia. A generalization of multiply perfect numbers. Scripta Mathematica, 19(??):209–210, ???? 1953. ISSN 0036-9713.

Garcia:1957:NAP

[Gar57] Mariano García. New amicable pairs. Scripta Mathematica, 23(??): 167–171, ???? 1957. ISSN 0036-9713.

Garcia:1958:NAP

[Gar58] Mariano García. New amicable pairs. Scripta Mathematica, 23(??): 167–171, ???? 1958. ISSN 0036-9713.

Gazis:1960:SRG

[GH60] Denos C. Gazis and Robert Herman. Square roots geometry and Archimedes. Scripta Mathematica, 25(??):228–241, ???? 1960. ISSN 0036-9713.

Gilbert:1956:KCM

[Gil56] E. N. Gilbert. Knots and classes of ménage permutations. Scripta Mathematica, 22(??):228–233 (1957), ???? 1956. ISSN 0036-9713.

Gilbert:1957:KCM

[Gil57] E. N. Gilbert. Knots and classes of menage permutations. Scripta Mathematica, 22(??):228–233, ???? 1957. ISSN 0036-9713.

Gilbert:1964:GAS

[Gil64a] R. P. Gilbert. On generalized axially symmetric potentials whose associates are distributions. Scripta Mathematica, 27(??):245–256, ???? 1964. ISSN 0036-9713.

Gilbert:1964:OWG

[Gil64b] R. P. Gilbert. Operators which generate harmonic functions in threevariables. Scripta Mathematica, 27(??):141–152, ???? 1964. ISSN 0036-9713.

Gilbert:1965:GAS

[Gil65] R. P. Gilbert. On generalized axially symmetric potentials whose associates are distributions. Scripta Mathematica, 27(??):245–256, ???? 1965. ISSN 0036-9713.

Gilmer:1973:NRO

[Gil73] Robert Gilmer. A note on rings with only finitely many subrings. Scripta Mathematica, 29(1–2):37–38, ???? 1973. ISSN 0036-9713.

Ginsburg:1932:JBMa

[Gin32a] Jekuthiel Ginsburg. Jacob ben Machir’s version of Menelaus’s work on spherical trigonometry. Scripta Mathematica, 1(??):72–78, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:JBMb

[Gin32b] Jekuthiel Ginsburg. Jacob ben Machir’s version of Menelaus’s work on spherical trigonometry. Scripta Mathematica, 1(??):153–155, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:EHD

[Gin32c] Jekuthiel Ginsburg. On the early history of the decimal point. Scripta Mathematica, 1(??):84–85, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:UTC

[Gin32d] Jekuthiel Ginsburg. On the use of the term cuento. Scripta Mathematica, 1(??):87, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:PSM

[Gin32e] Jekuthiel Ginsburg. The policy of Scripta Mathematica. Scripta Mathematica, 1(??):1–2, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:PM

[Gin32f] Jekuthiel Ginsburg. Predecessors of Magini. Scripta Mathematica, 1 (??):168–169, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:SST

[Gin32g] Jekuthiel Ginsburg. Spanish symbol for “thousand”. Scripta Mathematica, 1(??):264–265, ???? 1932. ISSN 0036-9713.

Ginsburg:1932:UMF

[Gin32h] Jekuthiel Ginsburg. An unknown mathematician of the fourteenth century. Scripta Mathematica, 1(??):60–62, ???? 1932. ISSN 0036-9713.

Ginsburg:1934:HUL

[Gin34] Jekuthiel Ginsburg. A hitherto unpublished letter by Benjamin Peirce. Scripta Mathematica, 2(??):278–282, ???? 1934. ISSN 0036-9713.

Ginsburg:1937:SM

[Gin37] Jekuthiel Ginsburg. Scripta Mathematica. Science, 86(2218):13, July 2, 1937. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://www.science.org/doi/10.1126/science.
86.2218.13-b. See [Ano37].

Ginsburg:1938:GAP

[Gin38a] Jekuthiel Ginsburg. Gauss’s arithmetization of the problem of 8 queens. Scripta Mathematica, 5(??):63–66, ???? 1938. ISSN 0036-9713.

Ginsburg:1938:GIB

[Gin38b] Jekuthiel Ginsburg. A geometrical interpretation of binomial coefficients. Scripta Mathematica, 5(??):141–142, ???? 1938. ISSN 0036-9713.

Ginsburg:1945:IE

[Gin45] Jekuthiel Ginsburg. Iterated exponentials. Scripta Mathematica, 11 (??):340–353, ???? 1945. ISSN 0036-9713.

Ginsburg:1954:RBC

[Gin54a] J. Ginsburg. A relationship between cubes of Fibonacci numbers. Scripta Mathematica, 19(??):242, December 1954. ISSN 0036-9713.

Ginsburg:1954:TER

[Gin54b] Jekuthiel Ginsburg. Triplets of equiareal rational triangles. Scripta Mathematica, 20(??):219–220, ???? 1954. ISSN 0036-9713.

Gloden:1948:BF

[Glo48a] A. Gloden. Binomial factorisations. Scripta Mathematica, 14(??):283, ???? 1948. ISSN 0036-9713.

Gloden:1948:SDE

[Glo48b] A. Gloden. A solution of a Diophantine equation. Scripta Mathematica, 14(??):185–186, ???? 1948. ISSN 0036-9713.

Gloden:1949:C

[Glo49a] A. Gloden. The congruence x^{8} − 1 ≡ 0 (mod p). Scripta Mathematica, 15(??):97–98, ???? 1949. ISSN 0036-9713.

Gloden:1949:NSD

[Glo49b] A. Gloden. Note on systems of diophantine equations. Scripta Mathematica, 15(??):163–164, ???? 1949. ISSN 0036-9713.

Gloden:1950:SC

[Glo50] A. Gloden. Solution of the congruence x^{4} + 1 ≡ 0(modp^{3}). Scripta Mathematica, 16(??):131–132, ???? 1950. ISSN 0036-9713.

Gloden:1951:RMI

[Glo51] A. Gloden. Remarkable multigrade identities. Scripta Mathematica, 17(??):151–152, ???? 1951. ISSN 0036-9713.

Gloden:1952:NDEa

[Glo52a] A. Gloden. Notes on Diophantine equations. Scripta Mathematica, 18(??):87–89, ???? 1952. ISSN 0036-9713.

Gloden:1952:NDEb

[Glo52b] A. Gloden. Notes on Diophantine equations. Scripta Mathematica, 18(??):177–179, ???? 1952. ISSN 0036-9713.

Gloden:1953:MA

[Glo53a] A. Gloden. Multigrade analysis. Scripta Mathematica, 18(??):312– 314, ???? 1953. ISSN 0036-9713.

Gloden:1953:NDE

[Glo53b] A. Gloden. Notes on diophantine equations. Scripta Mathematica, 18 (??):310–311, ???? 1953. ISSN 0036-9713.

Gloden:1953:SC

[Glo53c] A. Gloden. Solution of the congruences 2x^{2} + 1 ≡ 0 (mod p^{2}) and 2x^{2} + 1 ≡ 0 (mod p^{4}). Scripta Mathematica, 19(??):88–90, ???? 1953. ISSN 0036-9713.

Gloden:1955:TMS

[Glo55] Albert Gloden. Table of minimal solutions of the congruence x^{4} +1 ≡ 0 (mod p^{2}) for 4000 < p < 6000. Scripta Mathematica, 21(??):218, ???? 1955. ISSN 0036-9713.

Gloden:1960:CMS

[Glo60a] A. Gloden. Clasen’s method for the solution of a system of linear equations. Scripta Mathematica, 25(??):243–246, ???? 1960. ISSN 0036-9713.

Gloden:1960:DE

[Glo60b] A. Gloden. Diophantine equations. Scripta Mathematica, 25(??):167– 168, ???? 1960. ISSN 0036-9713.

Godfrey:1935:NBT

[God35] Edwin L. Godfrey. A note on Bieberbach’s trisection method. Scripta Mathematica, 3(??):326, ???? 1935. ISSN 0036-9713.

Golovensky:1932:MMR

[Gol32] David I. Golovensky. Maxima and minima in rabbinical literature. Scripta Mathematica, 1(??):53–55, ???? 1932. ISSN 0036-9713.

Goldberg:1952:SDS

[Gol52] Michael Goldberg. The squaring of developable surfaces. Scripta Mathematica, 18(??):17–24, ???? 1952. ISSN 0036-9713.

Goldberg:1956:CT

[Gol56] Michael Goldberg. Central tesselations. Scripta Mathematica, 21(??): 253–260, ???? 1956. ISSN 0036-9713.

Goldstein:1964:MTR

[Gol64] Bernard R. Goldstein. A medieval table for reckoning time from solar altitude. Scripta Mathematica, 27(??):61–66, ???? 1964. ISSN 0036-9713.

Goldstine:1973:BSC

[Gol73] Herman H. Goldstine. The beginnings of scientific computation. Scripta Mathematica, 29(3–4):379–389, ???? 1973. ISSN 0036-9713.

Goormaghtigh:1950:DFT

[Goo50] R. Goormaghtigh. Droz–Farny’s theorem. Scripta Mathematica, 16

(??):268–271, ???? 1950. ISSN 0036-9713.

Goormaghtigh:1952:LT

[Goo52] R. Goormaghtigh. Lemoyne’s theorem. Scripta Mathematica, 18(??):

182–184, ???? 1952. ISSN 0036-9713.

Good:1970:GBE

[Goo70] I. J. Good. A generalization of the Bernoulli–Euler partition formula.

Scripta Mathematica, 28(??):319–320, ???? 1970. ISSN 0036-9713.

Gorman:1973:DRM

[Gor73] Howard E. Gorman. Differential rings and modules. Scripta Mathe-

matica, 29(1–2):25–35, ???? 1973. ISSN 0036-9713.

Goudge:1943:SSL

[Gou43] Thomas A. Goudge. Science and symbolic logic. Scripta Mathematica,

9(2):69–80, ???? 1943. ISSN 0036-9713.

Gould:1963:SAF

[Gou63] H. W. Gould. The q-series analogue of a formula of Szily. Scripta

Mathematica, 26(??):155–157, ???? 1963. ISSN 0036-9713.

Grant:1954:PNT

[Gra54] Harold S. Grant. The prime number theorem. Scripta Mathematica,

20(??):235–236, ???? 1954. ISSN 0036-9713.

Gridgeman:1960:GPN

[Gri60] Norman T. Gridgeman. Geometric probability and the number π.

Scripta Mathematica, 25(??):183–195, ???? 1960. ISSN 0036-9713.

Gridgeman:1973:JNH

[Gri73a] Norman T. Gridgeman. John Napier and the history of logarithms.

Scripta Mathematica, 29(1–2):49–65, ???? 1973. ISSN 0036-9713.

Grilliot:1973:IDH

[Gri73b] Thomas J. Grilliot. Implicit definability and hyperprojectivity. Scripta Mathematica, 29(??):151–155, ???? 1973. ISSN 0036-9713.

Grossman:1945:TCP

[Gro45] Howard D. Grossman. The twelve-coin problem. Scripta Mathematica, 11(??):360–363, ???? 1945. ISSN 0036-9713.

Grossman:1950:FLP

[Gro50] Howard D. Grossman. Fun with lattice points. Scripta Mathematica, 16(??):207–212, ???? 1950. ISSN 0036-9713.

Grossman:1954:FLP

[Gro54] Howard D. Grossman. Fun with lattice points. Scripta Mathematica, 20(??):203–204, ???? 1954. ISSN 0036-9713.

Grossman:1956:FLP

[Gro56] Howard D. Grossman. Fun with lattice points. Scripta Mathematica, 22(??):153–158, ???? 1956. ISSN 0036-9713.

Grossman:1959:FLP

[Gro59] Howard D. Grossman. Fun with lattice points. Scripta Mathematica, 24(??):257–260, ???? 1959. ISSN 0036-9713.

Grunbaum:1955:DN

[Grü55] Hugo Grünbaum. On divisibility of numbers. Scripta Mathematica, 21(??):204–207, ???? 1955. ISSN 0036-9713.

Guillotin:1961:VT

[Gui61] M. R. Guillotin. Victor Thébault (1882–1960). Scripta Mathematica, 25(??):331–333, ???? 1961. ISSN 0036-9713.

Gupta:1953:GMF

[Gup53] Hansraj Gupta. A generalization of the Möbius function. Scripta Mathematica, 19(??):121–126, ???? 1953. ISSN 0036-9713.

Guttman:1947:UMS

[Gut47] Solomon Guttman. Universal magic squares and multigrade equations. Scripta Mathematica, 13(??):187–202, ???? 1947. ISSN 0036-9713.

Guttman:1948:NMO

[Gut48] Solomon Guttman. New magic in old magic squares. Scripta Mathematica, 14(??):284–286, ???? 1948. ISSN 0036-9713.

Hammerschmidt:1948:ANW

[Ham48a] William W. Hammerschmidt. Alfred North Whitehead [February 15, 1861 in Ramsgate — December 30, 1947 in Cambridge, MA]. Scripta Mathematica, 14(??):17–23 (1 plate), ???? 1948. ISSN 0036-9713.

Hammerschmidt:1948:OAN

[Ham48b] William W. Hammerschmidt. Obituary: Alfred North Whitehead. Scripta Mathematica, 14:17–23 (1 plate), 1948. ISSN 0036-9713.

Harary:1957:ATG

[Har57] Frank Harary. On arbitrarily traceable graphs and directed graphs. Scripta Mathematica, 23(??):37–41, ???? 1957. ISSN 0036-9713.

Harary:1958:ATG

[Har58] Frank Harary. On arbitrarily traceable graphs and directed graphs. Scripta Mathematica, 23(??):37–41, ???? 1958. ISSN 0036-9713.

Hargreave:1970:RPM

[Har70] David Hargreave. Reconstructing the planetary motions of the Eudoxean system. Scripta Mathematica, 28(4):335–345, ???? 1970. ISSN 0036-9713.

Hill:1935:ATGa

[HD35a] L. S. Hill and M. D. Darkow. On algebraic treatment of geometry on a spherical surface. I. Scripta Mathematica, 3(??):234–246, ???? 1935. ISSN 0036-9713.

Hill:1935:ATGb

[HD35b] L. S. Hill and M. D. Darkow. On algebraic treatment of geometry on a spherical surface. II. Scripta Mathematica, 3(??):329–336, ???? 1935. ISSN 0036-9713.

Heath:1935:CMSa

[Hea35a] R. V. Heath. Concentric magic squares. Scripta Mathematica, 3(??): 66–67, ???? 1935. ISSN 0036-9713.

Heath:1935:CMSb

[Hea35b] R. V. Heath. A curious magic square. Scripta Mathematica, 3(??): 250, ???? 1935. ISSN 0036-9713.

Heath:1935:MC

[Hea35c] R. V. Heath. A magic circle. Scripta Mathematica, 3(??):340, ???? 1935. ISSN 0036-9713.

Heath:1936:PMS

[Hea36] R. V. Heath. A panelled magic square. Scripta Mathematica, 4(??): 155, ???? 1936. ISSN 0036-9713.

Heath:1938:GMA

[Hea38] T. L. Heath. Greek mathematics and astronomy. Scripta Mathematica, 5(??):215–232, ???? 1938. ISSN 0036-9713.

Henderson:1955:LST

[Hen55] Archibald Henderson. The Lehmus–Steiner–Terquem problem in global survey. Scripta Mathematica, 21(??):223–232, ???? 1955. ISSN 0036-9713.

Henderson:1956:LST

[Hen56] Archibald Henderson. The Lehmus–Steiner–Terquem problem in global survey. Scripta Mathematica, 21(??):309–312, ???? 1956. ISSN 0036-9713.

Herstein:1973:TA

[Her73] I. N. Herstein. On a theorem of A. A. Albert. Scripta Mathematica, 29(3–4):391–394, ???? 1973. ISSN 0036-9713.

Hestenes:1973:TA

[Hes73] Magnus R. Hestenes. On a ternary algebra. Scripta Mathematica, 29 (3–4):253–272, ???? 1973. ISSN 0036-9713.

Higgins:1940:NHM

[Hig40] Thomas J. Higgins. A note on the history of mixed partial derivatives. Scripta Mathematica, 7(??):59–62, ???? 1940. ISSN 0036-9713.

Hill:1934:FBG

[Hil34] L. S. Hill. Fraenkel’s biography of Georg Cantor. Scripta Mathematica, 2(??):41–47, ???? 1934. ISSN 0036-9713.

Hill:1940:NRI

[Hil40] L. S. Hill. Notes on the regular icosahedron and the regular dodecahedron. Scripta Mathematica, 7(??):99–109, ???? 1940. ISSN 0036-9713.

Harary:1970:T

[HM70] Frank Harary and Bennet Manvel. Trees. Scripta Mathematica, 28 (??):327–333, ???? 1970. ISSN 0036-9713.

Hodges:1966:SPM

[Hod66] John H. Hodges. Some pairs of matrix equations over a finite field. Scripta Mathematica, 27(??):289–301, ???? 1966. ISSN 0036-9713.

Horner:1953:AMM

[Hor53] Walter W. Horner. Addition–multiplication magic squares. Scripta Mathematica, 18(??):300–303, ???? 1953. ISSN 0036-9713.

Horner:1955:AMM

[Hor55] Walter W. Horner. Addition–multiplication magic square of order 8. Scripta Mathematica, 21(??):23–27, ???? 1955. ISSN 0036-9713.

Huntington:1938:DLa

[Hun38a] E. V. Huntington. The duplicity of logic. Scripta Mathematica, 5(??): 149–157, ???? 1938. ISSN 0036-9713.

Huntington:1938:DLb

[Hun38b] E. V. Huntington. The duplicity of logic. Scripta Mathematica, 5(??): 233–238, ???? 1938. ISSN 0036-9713.

Huxley:1959:MWE

[Hux59] G. L. Huxley. The mathematical work of Edmond Halley. Scripta Mathematica, 24(??):265–273, ???? 1959. ISSN 0036-9713.

Huxley:1960:GWC

[Hux60a] G. L. Huxley. The geometrical work of Christopher Wren. Scripta Mathematica, 25(??):201–208, ???? 1960. ISSN 0036-9713.

Huxley:1960:MWE

[Hux60b] G. L. Huxley. The mathematical work of Edmond Halley. Scripta Mathematica, 24(??):265–273, ???? 1960. ISSN 0036-9713.

Huxley:1963:RCN

[Hux63] G. Huxley. Roger Cotes and natural philosophy. Scripta Mathematica, 26(??):231–238, ???? 1963. ISSN 0036-9713.

al-Khayyami:1960:DDE

[iAa60] Omar ibn Abrahim al-Khayyami (Omar Khayyam). Discussion of difficulties in Euclid. Tranl. by Ali R. Amir–Móez. Scripta Mathematica, 24(??):275–303, ???? 1960. ISSN 0036-9713.

Hano:1973:EPI

[iH73] Jun ichi Hano. Equivariant projective immersion of a complex coset space with non-degenerate canonical Hermitian form. Scripta Mathematica, 29(??):125–139, ???? 1973. ISSN 0036-9713.

Irani:1952:SMT

[Ira52] Rida A. K. Irani. A sexagesimal multiplication table in the Arabic alphabetical system. Scripta Mathematica, 18(??):92–93, ???? 1952. ISSN 0036-9713.

Ivins:1943:NGD

[Ivi43] William M. Ivins, Jr. A note on Gerard Desargues. Scripta Mathematica, 9(??):33–48, ???? 1943. ISSN 0036-9713.

Ivins:1947:NDT

[Ivi47] William M. Ivins, Jr. A note on Desargues’ theorem. Scripta Mathematica, 13(??):203–210, ???? 1947. ISSN 0036-9713.

Jarden:1954:NC

[Jar54] Dov Jarden. Nullifying coefficients. Scripta Mathematica, 19(??): 239–241, ???? 1954. ISSN 0036-9713.

Jacobson:1973:GAQ

[JK73] Nathan Jacobson and J. Katz. Generically algebraic quadratic Jordan algebras. Scripta Mathematica, 29(3–4):215–227, ???? 1973. ISSN 0036-9713.

Jones:1939:NF

[Jon39] H. Jones. Note on the factors of a^{10} + a^{5} + 1. Scripta Mathematica, 6(??):59–60, ???? 1939. ISSN 0036-9713.

Jones:1947:IAH

[Jon47] P. J. Jones. The identity of the author of a hitherto anonymous work. Scripta Mathematica, 13(??):119–120, ???? 1947. ISSN 0036-9713.

Jones:1973:QGQ

[Jon73] Burton W. Jones. Quasi-genera of quadratic forms. Scripta Mathematica, 29(3–4):359–370, ???? 1973. ISSN 0036-9713.

Juzuk:1945:EBN

[JT45] D. Juzuk and Z. Tuchman. Elementary bounds for the number of primes. Scripta Mathematica, 11(??):179–182, ???? 1945. ISSN 0036-

Jurkat:1959:SGPa

[Jur59a] W. B. Jurkat. On semi-groups of positive matrices. I. Scripta Mathematica, 24(??):123–131, ???? 1959. ISSN 0036-9713.

Jurkat:1959:SGPb

[Jur59b] W. B. Jurkat. On semi-groups of positive matrices. II. Scripta Mathematica, 24(??):207–218, ???? 1959. ISSN 0036-9713.

Juzuk:1939:CIO

[Juz39] Dov Juzuk. Curiosa 56: An interesting observation. Scripta Mathematica, 6(??):218–??, ???? 1939. ISSN 0036-9713.

Kalisch:1973:OLP

[Kal73] G. K. Kalisch. On operators with large point spectrum. Scripta Mathematica, 29(3–4):371–378, ???? 1973. ISSN 0036-9713. URL

https://doi.org/10.1016/s0016-0032(40)90875-4.

Kaniel:1956:EMB

[Kan56] Shmuel Kaniel. Extension of the method of Bonse Zermelo. Scripta Mathematica, 21(??):269–272, ???? 1956. ISSN 0036-9713.

Kaprekar:1949:ASG

[Kap49] D. R. Kaprekar. Another solitaire game. Scripta Mathematica, 15 (??):244–245, ???? 1949. ISSN 0036-9713.

Kaprekar:1953:PIR

[Kap53] D. R. Kaprekar. Problems involving reversal of digits. Scripta Mathematica, 19(??):81–82, ???? 1953. ISSN 0036-9713. URL https://

doi.org/10.1080/07468342.1988.11973098.

Kaprekar:1955:IPN

[Kap55] D. R. Kaprekar. An interesting property of the number 6174. Scripta Mathematica, 21(??):244–245, ???? 1955. ISSN 0036-9713. URL

https://en.wikipedia.org/wiki/6174.

Kaplansky:1973:IDL

[Kap73] Irving Kaplansky. Infinite-dimensional Lie algebras. Scripta Mathematica, 29(3–4):237–241, ???? 1973. ISSN 0036-9713.

Karpinski:1932:RMB

[Kar32] Louis C. Karpinski. Rare mathematical books in the University of Michigan Library. Scripta Mathematica, 1(??):63–65, ???? 1932. ISSN 0036-9713.

Karpinski:1933:ACT

[Kar33] Louis C. Karpinski. Arithmetic centenarians: Textbooks with a long life. Scripta Mathematica, 2(??):34–40, ???? 1933. ISSN 0036-9713.

Karpinski:1934:ACT

[Kar34] Louis C. Karpinski. Arithmetic centenarians; textbooks with a long life. Scripta Mathematica, 2(??):34–40, ???? 1934. ISSN 0036-9713.

Karpinski:1935:EGF

[Kar35] Louis C. Karpinski. The elusive George Fisher “accomptant” — writer or editor of three popular arithmetics. Scripta Mathematica, 3(??): 337–339, ???? 1935. ISSN 0036-9713.

Karpinski:1939:OMT

[Kar39] Louis Charles Karpinski. The origin of the mathematics as taught to freshmen. Scripta Mathematica, 6(??):??, ???? 1939. ISSN 0036-9713.

Karpinski:1941:SBM

[Kar41] Louis C. Karpinski. Supplement to the bibliography of mathematical works printed in America through 1850. Scripta Mathematica, 8(??): 233–236, ???? 1941. ISSN 0036-9713.

Karpinski:1943:PCT

[Kar43] Louis C. Karpinski. The progress of the Copernican theory. Scripta Mathematica, 9(??):139–154, ???? 1943. ISSN 0036-9713.

Karpinski:1944:AW

[Kar44] Louis C. Karpinski. Algebraic works to 1700. Scripta Mathematica, 10(??):149–169, ???? 1944. ISSN 0036-9713.

Karpinski:1945:PTD

[Kar45] Louis C. Karpinski. The place of trigonometry in the development of mathematical sciences. Scripta Mathematica, 11(3–4):268–272, ???? 1945. ISSN 0036-9713.

Karpinski:1946:BCL

[Kar46] Louis Charles Karpinski. Bibliographical check list of all works on trigonometry published up to 1700 A. D. Scripta Mathematica, 12 (??):267–283, ???? 1946. ISSN 0036-9713.

Karpinski:1947:MLA

[Kar47] Louis C. Karpinski. Mathematics in Latin America. A brief survey of their publications to 1850. Scripta Mathematica, 13(??):59–63, ???? 1947. ISSN 0036-9713.

Karpinski:1954:TSB

[Kar54] Louis C. Karpinski. Third supplement to the “Bibliography of mathematical works printed in America through 1850”. Scripta Mathematica, 20(??):197–202 (1955), ???? 1954. ISSN 0036-9713.

Karst:1970:LTM

[Kar70] E. Karst. Lists of ten or more primes in arithmetical progression. Scripta Mathematica, 28(??):313–317, ???? 1970. ISSN 0036-9713.

Kasner:1934:U

[Kas34] Edward Kasner. [unknown]. Scripta Mathematica, 2(??):131–138,

???? 1934. ISSN 0036-9713.

Kasner:1943:NNA

[Kas43] Edward Kasner. Note on non-Apollonian packing in space. Scripta Mathematica, 9(??):26, ???? 1943. ISSN 0036-9713.

Kasner:1945:RTH

[Kas45] Edward Kasner. The recent theory of the horn angle. Scripta Mathematica, 11(??):263–267, ???? 1945. ISSN 0036-9713.

Kasner:1947:NPT

[Kas47] Edward Kasner. Neo-Pythagorean triangles. Scripta Mathematica, 13(??):43–47, ???? 1947. ISSN 0036-9713.

Kaufman:1967:BNH

[Kau67] H. Kaufman. A bibliographical note on higher order sine functions. Scripta Mathematica, 28(??):29–36, ???? 1967. ISSN 0036-9713.

Kasner:1936:GMT

[KC36] Edward Kasner and George Comenetz. Groups of multipoint transformations with application to polygons. Scripta Mathematica, 4(??): 37–49, ???? 1936. ISSN 0036-9713.

Kasner:1943:CPC

[KCW43] Edward Kasner, George Comenetz, and John Wilkes. The covering of the plane by circles. Scripta Mathematica, 9(??):19–25, ???? 1943. ISSN 0036-9713.

Kasner:1947:GPT

[KD47] Edward Kasner and John De Cicco. General polar theory. Scripta Mathematica, 13(??):53–57, ???? 1947. ISSN 0036-9713.

Kasner:1949:SGP

[KD49] Edward Kasner and John De Cicco. Survey of the geometry of physical systems of curves. Scripta Mathematica, 15(??):193–199, ???? 1949. ISSN 0036-9713.

Keller:1932:NOH

[Kel32] Henry Keller. Numerics in old Hebrew medical lore. Scripta Mathematica, 1(??):66–67, ???? 1932. ISSN 0036-9713.

Kelisky:1957:FIB

[Kel57] Richard P. Kelisky. On formulas involving both the Bernoulli and Fibonacci numbers. Scripta Mathematica, 23(??):27–35, ???? 1957. ISSN 0036-9713.

Kelisky:1958:FIB

[Kel58] Richard P. Kelisky. On formulas involving both the Bernoulli and Fibonacci numbers. Scripta Mathematica, 23(??):27–35, ???? 1958. ISSN 0036-9713.

Kennedy:1951:FCL

[Ken51] Edward S. Kennedy. A fifteenth century lunar eclipse computer. Scripta Mathematica, 17(??):91–97, ???? 1951. ISSN 0036-9713.

Kennedy:1959:BGD

[Ken59] Edward S. Kennedy. Bīrūnī’s graphical determination of the local meridian. Scripta Mathematica, 24(??):251–255, ???? 1959. ISSN 0036-9713.

Kennedy:1964:AKJ

[Ken64] Edward S. Kennedy. Al-Khwārizmī on the Jewish calendar. Scripta Mathematica, 27(??):55–59, ???? 1964. ISSN 0036-9713.

Keyser:1932:BM

[Key32a] Cassius Jackson Keyser. The bearings of mathematics. Scripta Mathematica, 1(??):93–110, ???? 1932. ISSN 0036-9713.

Keyser:1932:MCC

[Key32b] Cassius Jackson Keyser. Mathematics as a culture clue. Scripta Math-

ematica, 1(??):185–203, ???? 1932. ISSN 0036-9713.

Keyser:1932:MM

[Key32c] Cassius Jackson Keyser. The meaning of mathematics. Scripta Math-

ematica, 1(??):15–28, ???? 1932. ISSN 0036-9713.

Keyser:1933:CSS

[Key33a] Cassius Jackson Keyser. The cultural significance of Scripta Mathe-

matica. Scripta Mathematica, ??(??):??, ???? 1933. ISSN 0036-9713.

Keyser:1933:MSS

[Key33b] Cassius Jackson Keyser. Mathematics and the science of semantics.

Scripta Mathematica, 2(??):29–31, 1933. ISSN 0036-9713.

Keyser:1934:MQC

[Key34a] Cassius Jackson Keyser. Mathematics and the question of cosmic mind. Scripta Mathematica, 2(??):125–129, ???? 1934. ISSN 0036-9713.

Keyser:1934:MSS

[Key34b] Cassius Jackson Keyser. Mathematics and the science of semantics.

Scripta Mathematica, 2(??):247–260, ???? 1934. ISSN 0036-9713.

Keyser:1934:NDFa

[Key34c] Cassius Jackson Keyser. The nature of the doctrinal function and its role in rational thought. Scripta Mathematica, 2(??):5–26, ???? 1934. ISSN 0036-9713.

Keyser:1934:NDFb

[Key34d] Cassius Jackson Keyser. The nature of the doctrinal function and its rôle in rational thought. Scripta Mathematica, 2(??):143–154, ???? 1934. ISSN 0036-9713.

Keyser:1934:WBS

[Key34e] Cassius Jackson Keyser. William Benjamin Smith. Scripta Mathe-

matica, 2(??):305–311, ???? 1934. ISSN 0036-9713.

Keyser:1935:GSI

[Key35a] Cassius Jackson Keyser. A glance at some of the ideas of Charles Sanders Peirce. Scripta Mathematica, 3(??):11–37, ???? 1935. ISSN 0036-9713.

Keyser:1935:TGS

[Key35b] Cassius Jackson Keyser. Three great synonyms: relation, transformation, function. Scripta Mathematica, 3(??):301–316, ???? 1935. ISSN 0036-9713.

Keyser:1936:RIC

[Key36a] Cassius Jackson Keyser. The role of infinity in the cosmology of Epicurus. Scripta Mathematica, 4(??):221–240, ???? 1936. ISSN 0036-9713.

Keyser:1936:VFD

[Key36b] Cassius Jackson Keyser. Vilfredo Federico Damaso Pareto. Mathematician, economist, sociologist. Scripta Mathematica, 4(??):5–23, ???? 1936. ISSN 0036-9713.

Keyser:1938:MPH

[Key38] Cassius Jackson Keyser. A mathematical prodigy: History and legend. Scripta Mathematica, 5(??):83–94, ???? 1938. ISSN 0036-9713.

Keyser:1939:P

[Key39] Cassius Jackson Keyser. Pythagoras. Scripta Mathematica, 6(??): 17–22, ???? 1939. ISSN 0036-9713.

Keyser:1941:CSP

[Key41] Cassius Jackson Keyser. Charles Sanders Peirce as a pioneer. Galois Lectures, Scripta Mathematica Library, 1941(5):87–112, 1941.

Keyser:1947:CWC

[Key47] Cassius Jackson Keyser. The Collected Works of Cassius Jackson Keyser …: Mathematics as a culture clue, and other essays., volume 1. Scripta Mathematica, New York, NY, USA, 1947. ???? pp.

Keyser:1949:MM

[Key49] Cassius Jackson Keyser. The meaning of mathematics. Scripta Mathematica, 15(??):??, 1949. ISSN 0036-9713.

Kasner:1950:VMH

[KH50] Edward Kasner and Irene Harrison. Voltaire on mathematics and horn angles. Scripta Mathematica, 16(??):13–21, ???? 1950. ISSN 0036-9713.

Kasner:1951:THA

[KH51] Edward Kasner and Irene Harrison. The trisection of horn angles. Scripta Mathematica, 17(??):231–235, ???? 1951. ISSN 0036-9713.

Khatri:1955:TNP

[Kha55] M. N. Khatri. Triangular numbers and Pythagorean triangles. Scripta Mathematica, 21(??):94–??, ???? 1955. ISSN 0036-9713.

Kasner:1934:CPG

[KJWC34] E. Kasner, E. Jomas, K. Way, and G. Comenetz. Centroidal polygons and groups. Scripta Mathematica, 2(??):131–138, ???? 1934. ISSN 0036-9713.

Klamkin:1954:RPC

[Kla54] Murray S. Klamkin. On rational points on a circle. Scripta Mathematica, 20(??):222–223, ???? 1954. ISSN 0036-9713.

Klamkin:1955:SIL

[Kla55] Murray S. Klamkin. On some identities of Lucas. Scripta Mathematica, 21(??):213–214, ???? 1955. ISSN 0036-9713.

Kleinfeld:1973:MGC

[Kle73] Erwin Kleinfeld. More on a generalization of commutative and associative rings. Scripta Mathematica, 29(3–4):351–357, ???? 1973. ISSN 0036-9713.

Karlin:1961:HPF

[KM61] Samuel Karlin and James L. McGregor. The Hahn polynomials, formulas and an application. Scripta Mathematica, 26(??):33–46, ???? 1961. ISSN 0036-9713.

Kaplansky:1946:PM

[KR46] Irving Kaplansky and John Riordan. The problème des ménages. Scripta Mathematica, 12(??):113–124, ???? 1946. ISSN 0036-9713.

Kraitchik:1945:CRC

[Kra45] Maurice Kraitchik. On certain rational cuboids. Scripta Mathematica, 11(??):317–326, ???? 1945. ISSN 0036-9713.

Kraitchik:1948:DF

[Kra48] Maurice Kraitchik. On the divisibility of factorials. Scripta Mathematica, 14(??):24–26, ???? 1948. ISSN 0036-9713.

Kraitchik:1952:F

[Kra52] Maurice Kraitchik. On the factorization of 2^{n} ± 1. Scripta Mathematica, 18(??):39–52, ???? 1952. ISSN 0036-9713.

Kramer:1957:SMF

[Kra57] Edna E. Kramer. Six more female mathematicians. Scripta Mathematica, 23(??):83–95, ???? 1957. ISSN 0036-9713.

Kramer:1958:SMF

[Kra58] Edna E. Kramer. Six more female mathematicians. Scripta Mathematica, 23(??):83–95, ???? 1958. ISSN 0036-9713.

Loria:1938:NGP

[LA38] Gino Loria and H. T. R. Aude. Notes on generalized Pythagorean numbers. Scripta Mathematica, 5(??):142–143, ???? 1938. ISSN 0036-9713.

Lafleur:1941:FIM

[Laf41] Laurence J. Lafleur. A functionalistic interpretation of mathematics. Scripta Mathematica, 8(1):27–33, March 1941. ISSN 0036-9713.

Lafleur:1947:ASB

[Laf47] Laurence J. Lafleur. Ambiguities in the Schröder–Bernstein theorem. Scripta Mathematica, 13(3–4):169–175, ???? 1947. ISSN 0036-9713.

Langer:1935:LLEa

[Lan35a] Rudolph Ernest Langer. The life of Léonard Euler. Scripta Mathematica, 3(??):61–66, ???? 1935. ISSN 0036-9713.

Langer:1935:LLEb

[Lan35b] Rudolph Ernest Langer. The life of Léonard Euler. Scripta Mathematica, 3(??):131–138, ???? 1935. ISSN 0036-9713.

Langer:1936:IN

[Lan36] R. E. Langer. Isaac Newton. Scripta Mathematica, 4(??):241–255, ???? 1936. ISSN 0036-9713.

Langman:1951:CDP

[Lan51] Harry Langman. Curiosa 261: A disc puzzle. Scripta Mathematica, 17(1–2):144, 1951. ISSN 0036-9713.

Langman:1953:CEO

[Lan53] Harry Langman. Curiosa 342: Easy but not obvious. Scripta Mathe-

matica, 19(4):242, 1953. ISSN 0036-9713.

Langman:1954:PC

[Lan54] Harry Langman. A problem in checkers. Scripta Mathematica, 20

(??):206–208, ???? 1954. ISSN 0036-9713.

Landin:1963:PPF

[Lan63] Joseph Landin. Polynomials and polynomial functions. Scripta Math-

ematica, 26(??):239–250, ???? 1963. ISSN 0036-9713.

Laposky:1952:EA

[Lap52] B. F. Laposky. Electronic abstractions. Scripta Mathematica, 18(??):

305–308, ???? 1952. ISSN 0036-9713.

Laposky:1953:EA

[Lap53] B. Laposky. Electronic abstractions. Scripta Mathematica, 19(??):

196–197, ???? 1953. ISSN 0036-9713.

Larguier:1940:BPM

[Lar40] E. H. Larguier. Brouwerian philosophy of mathematics. Scripta Math-

ematica, 7(??):69–78, ???? 1940. ISSN 0036-9713.

Larguier:1941:PM

[Lar41] Everett H. Larguier. Postulational methods. Scripta Mathematica, 8

(??):99–109, ???? 1941. ISSN 0036-9713.

Lasker:1938:NKD

[Las38] E. Lasker. Note on Keyser’s discussion of Epicurus. Scripta Mathe-

matica, 5(??):121–123, ???? 1938. ISSN 0036-9713.

Lehmer:1933:HBG

[Leh33] D. N. Lehmer. Hunting big game in the theory of numbers. Scripta

Mathematica, 1(??):229–235, ???? 1933. ISSN 0036-9713.

Lehmer:1947:TEP

[Leh47] D. H. Lehmer. The Tarry–Escott problem. Scripta Mathematica, 13 (??):37–41, ???? 1947. ISSN 0036-9713.

Lehmer:1951:TNF

[Leh51] D. H. Lehmer. A triangular number formula for the partition function. Scripta Mathematica, 17(??):17–19, ???? 1951. ISSN 0036-9713.

Lehmer:1963:NPR

[Leh63] D. H. Lehmer. A note on primitive roots. Scripta Mathematica, 26 (??):117–119, ???? 1963. ISSN 0036-9713.

Lehmer:1973:CFC

[Leh73] D. H. Lehmer. Continued fractions containing arithmetic progressions. Scripta Mathematica, 29(1–2):17–24, ???? 1973. ISSN 0036-9713.

Leonardi:1954:SBM

[Leo54] Raffaele Leonardi. Some bimagic matrices. Scripta Mathematica, 20 (??):165–167 (1955), ???? 1954. ISSN 0036-9713.

Levine:1960:NNP

[Lev60] Jack Levine. Note on the number of pairs of non-intersecting routes. Scripta Mathematica, 24(??):335–338, ???? 1960. ISSN 0036-9713.

Levin:1970:ISL

[Lev70] Mark Levin. Infinite sums of locally convex spaces with a projection property. Scripta Mathematica, 28(??):307–311, ???? 1970. ISSN 0036-9713.

Lewittes:1932:JMP

[Lew32] M. H. Lewittes. Joseph del Medigo on prosthaphaeresis. Scripta Mathematica, 1(??):56–59, ???? 1932. ISSN 0036-9713.

Loria:1932:ML

[LFA32] Gino Benedetto Loria, Abraham Adolf Fraenkel, and R. C. Archibald. Mathematics and law. Scripta Mathematica, 1(??):276, ???? 1932. ISSN 0036-9713.

Lightstone:1963:SAD

[Lig63] A. H. Lightstone. A simple alternative to Dedekind cuts. Scripta Mathematica, 26(??):347–351, ???? 1963. ISSN 0036-9713.

Lightstone:1964:CDL

[Lig64] A. H. Lightstone. A constructive definition of the limit of a sequence. Scripta Mathematica, 27(??):153–158, ???? 1964. ISSN 0036-9713.

Luchins:1964:L

[LL64] Edith H. Luchins and Abraham S. Luchins. Logicism. Scripta Mathematica, 27(??):223–243, ???? 1964. ISSN 0036-9713.

Luchins:1965:L

[LL65] E. H. Luchins and A. S. Luchins. Logicism. Scripta Mathematica, 27 (??):223–243, ???? 1965. ISSN 0036-9713.

Locke:1932:SA

[Loc32a] L. L. Locke. Synchronism and anachronism. Scripta Mathematica, 1 (??):147–152, ???? 1932. ISSN 0036-9713.

Locke:1932:APA

[Loc32b] L. Leland Locke. The ancient Peruvian abacus. Scripta Mathematica, 1(??):37–43, ???? 1932. ISSN 0036-9713.

Locke:1933:CLA

[Loc33] L. L. Locke. The contributions of Leibniz to the art of mechanical calculation. Scripta Mathematica, 1(??):315–321, ???? 1933. ISSN 0036-9713.

Locke:1941:RRG

[Loc41] L. Leland Locke. A royal road to geometry. Scripta Mathematica, 8 (1):34–42, March 1941. ISSN 0036-9713.

Loria:1932:CHA

[Lor32] Gino Loria. A. L. Cauchy in the history of analytic geometry. Scripta Mathematica, 1(??):123–128, ???? 1932. ISSN 0036-9713.

Loria:1934:MSM

[Lor34] Gino Loria. A. Mannheim — soldier and mathematician. Scripta Mathematica, 2(??):337–342, ???? 1934. ISSN 0036-9713.

Loria:1936:JLHa

[Lor36a] Gino Loria. J. Liouville and his work. Scripta Mathematica, 4(??): 147–154, ???? 1936. ISSN 0036-9713.

Loria:1936:JLHb

[Lor36b] Gino Loria. J. Liouville and his work. Scripta Mathematica, 4(??): 257–262, ???? 1936. ISSN 0036-9713.

Loria:1936:JLHc

[Lor36c] Gino Loria. J. Liouville and his work. Scripta Mathematica, 4(??): 301–306, ???? 1936. ISSN 0036-9713.

Loria:1947:EJS

[Lor47a] Gino Loria. Ernest de Jonquières, sailor and scientist. Scripta Mathematica, 13(??):5–15 (1 plate), ???? 1947. ISSN 0036-9713.

Loria:1947:PTE

[Lor47b] Gino Loria. Paul Tannery, engineer and historian. Scripta Mathematica, 13(??):155–162, ???? 1947. ISSN 0036-9713.

Lorch:1957:STC

[Lor57] Lee Lorch. Supplement to a theorem of Cesàro. Scripta Mathematica, 23(??):163–165, ???? 1957. ISSN 0036-9713.

Lorch:1958:STC

[Lor58] Lee Lorch. Supplement to a theorem of Cesàro. Scripta Mathematica, 23(??):163–165, ???? 1958. ISSN 0036-9713.

Lowenheim:1946:MIP

[Löw46] Leopold Löwenheim. On making indirect proofs direct. Scripta Mathematica, 12(??):125–139, ???? 1946. ISSN 0036-9713.

Lowan:1949:CLN

[Low49] Arnold N. Lowan. The Computation Laboratory of the National Bureau of Standards. Scripta Mathematica, 15(??):33–63, ???? 1949. ISSN 0036-9713.

Lowan:1950:VAT

[Low50] Arnold N. Lowan. On the vector-analytical treatment of certain problems. Scripta Mathematica, 16(??):129–131, ???? 1950. ISSN 0036-9713.

Lowan:1955:NEM

[Low55] Arnold N. Lowan. Note on an elementary method for generating inequalities. Scripta Mathematica, 21(2–3):218–220, ???? 1955. ISSN 0036-9713.

Lowan:1956:CVI

[Low56] Arnold N. Lowan. On the convergence of various iteration processes. Scripta Mathematica, 22(??):222–227, ???? 1956. ISSN 0036-9713.

Lowan:1957:CVI

[Low57] Arnold N. Lowan. On the convergence of various iteration processes. Scripta Mathematica, 22(??):222–227, ???? 1957. ISSN 0036-9713.

Lubin:1959:TLC

[Lub59] C. I. Lubin. The theorem of Lehmus and complex numbers. Scripta Mathematica, 24(??):137–140, ???? 1959. ISSN 0036-9713.

MacDuffee:1944:ADH

[Mac44] C. C. MacDuffee. Algebra’s debt to Hamilton. Scripta Mathematica, 10(??):25–35 (1 plate), ???? 1944. ISSN 0036-9713.

Malkin:1959:SA

[Mal59] I. Malkin. A sad anniversary. Scripta Mathematica, 24(??):79–81, ???? 1959. ISSN 0036-9713.

Malkin:1963:ABD

[Mal63] I. Malkin. On the 150th anniversary of the birth date of an immortal in mathematics. Scripta Mathematica, 26(3):197–200, ???? 1963. ISSN 0036-9713.

Malkin:1964:TDE

[Mal64] I. Malkin. Two dimensional elasticity and Muskhelishvili’s contribution to it. Scripta Mathematica, 27(??):49–53, ???? 1964. ISSN 0036-9713.

Mandan:1961:HCE

[Man61] Sahib Ram Mandan. Harmonic chains of equal circles. Scripta Mathematica, 26(??):47–64, ???? 1961. ISSN 0036-9713.

Mordukhai-Boltovskoy:1932:CIH

[MB32a] D. Mordukhai-Boltovskoy. The concept of infinity. Historical and critical notes. Scripta Mathematica, 1(??):132–134, ???? 1932. ISSN 0036-9713.

Mordukhai-Boltovskoy:1932:CII

[MB32b] D. Mordukhai-Boltovskoy. The concept of infinity. III: The dissolution of contradictions in infinity. Scripta Mathematica, 1(??):252–253, ???? 1932. ISSN 0036-9713.

McCoy:1938:AMS

[McC38] J. C. McCoy. The anatomy of magic squares. Scripta Mathematica, 5(??):137–141, ???? 1938. ISSN 0036-9713.

McCoy:1940:AMS

[McC40] J. C. McCoy. The anatomy of magic squares. Scripta Mathematica, 7(??):143–153, ???? 1940. ISSN 0036-9713.

McCoy:1941:MMT

[McC41] John Calvin McCoy. Manuel Moschopoulos’s treatise on magic squares. Scripta Mathematica, 8(1):15–26, March 1941. ISSN 0036-9713.

McCarthy:1957:OPN

[McC57] Paul J. McCarthy. Odd perfect numbers. Scripta Mathematica, 23 (??):43–47, ???? 1957. ISSN 0036-9713.

McCarthy:1958:OPN

[McC58] Paul J. McCarthy. Odd perfect numbers. Scripta Mathematica, 23 (??):43–47, ???? 1958. ISSN 0036-9713.

McCrimmon:1973:GNI

[McC73a] Kevin McCrimmon. The generic norm of an isotope of a Jordan algebra. Scripta Mathematica, 29(3–4):229–236, ???? 1973. ISSN 0036-9713.

McCrimmon:1973:SNQ

[McC73b] Kevin McCrimmon. Solvability and nilpotence for quadratic Jordan algebras. Scripta Mathematica, 29(3–4):467–483, ???? 1973. ISSN 0036-9713.

Meder:1940:RR

[Med40] A. E. Meder. Reseña de Russell, 1938. Scripta Mathematica, 7(??): 138–141, 1940. ISSN 0036-9713.

Mendelsohn:1939:CC

[Men39] Charles Jastrow Mendelsohn. Cardan on cryptography. Scripta Mathematica, 6(??):??, 1939. ISSN 0036-9713.

Menger:1956:CG

[Men56a] Karl Menger. Calculus 1950 — Geometry 1880. I. Scripta Mathematica, 22(??):89–96, ???? 1956. ISSN 0036-9713.

Menger:1956:CGI

[Men56b] Karl Menger. Calculus 1950 — Geometry 1880. II. Scripta Mathematica, 22(??):203–206 (1957), ???? 1956. ISSN 0036-9713.

Mikami:1932:MMM

[Mik32] Yoshio Mikami. Mamoru Mimori. A master teacher of mathematics in Japan. Scripta Mathematica, 1(??):254–255, ???? 1932. ISSN 0036-9713.

Mirsky:1948:RDH

[Mir48] L. Mirsky. A remark on D. H. Lehmer’s solution of the Tarry–Escott problem. Scripta Mathematica, 14(??):126–127, ???? 1948. ISSN 0036-9713.

Mirsky:1949:TRI

[Mir49] L. Mirsky. A theorem on representations of integers in the scale of r. Scripta Mathematica, 15(??):11–12, ???? 1949. ISSN 0036-9713.

Miura:1973:UAS

[Miu73] Robert M. Miura. Uniform asymptotic solution for linearized nearcontinuum plane Couette flow. Scripta Mathematica, 29(1–2):77–97, ???? 1973. ISSN 0036-9713.

Macon:1950:DFD

[MM50] N. Macon and L. Moser. On the distribution of first digits of powers. Scripta Mathematica, 16(??):290–291, ???? 1950. CODEN ???? ISSN 0036-9713.

Moessner:1947:CCM

[Moe47] Alfred Moessner. Curiosa #161: A curious magic square. Scripta Mathematica, 13(??):234–??, ???? 1947. ISSN 0036-9713.

Moessner:1952:MI

[Moe52] Alfred Moessner. On the multiple identity x_{1}^{n} + x_{2}^{n} + x_{3}^{n} + x_{4}^{n} + x_{5}^{n} = y_{1}^{n} + y_{2}^{n} + y_{3}^{n} + y_{4}^{n} + y_{5}^{n} for n = 1, 3, 5, 7. Scripta Mathematica, 18(??): 90–91, ???? 1952. ISSN 0036-9713.

Moessner:1953:ISC

[Moe53] Alfred Moessner. Identities with specia conditions. Scripta Mathematica, 18(??):311–312, ???? 1953. ISSN 0036-9713.

Moritz:1936:BGF

[Mor36] R. E. Moritz. On the beauty of geometrical forms. Scripta Mathematica, 4(??):25–35, ???? 1936. ISSN 0036-9713.

Moritz:1939:DNR

[Mor39] R. E. Moritz. The dual of Napier’s rules of the circular parts. Scripta Mathematica, 6(??):23–25, ???? 1939. ISSN 0036-9713.

Mordell:1960:RWN

[Mor60] L. J. Mordell. Recent work in number theory. Scripta Mathematica, 25(??):93–103, ???? 1960. ISSN 0036-9713.

Mordell:1963:RPC

[Mor63] L. J. Mordell. Rational points on cubic curves and surfaces. Scripta Mathematica, 26(??):147–153, ???? 1963. ISSN 0036-9713.

Mordell:1964:DE

[Mor64] L. J. Mordell. The Diophantine equation y^{2} = ax^{3} + bx^{2} + cx + d. Scripta Mathematica, 27(??):205–211, ???? 1964. ISSN 0036-9713.

Mordell:1965:DE

[Mor65] L. J. Mordell. The Diophantine equation y^{2} = ax^{3} + bx^{2} + cx + d. Scripta Mathematica, 27(??):205–211, ???? 1965. ISSN 0036-9713.

Moser:1953:DE

[Mos53] Leo Moser. On the Diophantine equation 1^{n}+2^{n}+3^{n}+⋯+(m−1)^{n} = m^{n}. Scripta Mathematica, 19(??):84–88, ???? 1953. ISSN 0036-9713.

Moser:1957:PC

[Mos57] Leo Moser. On a problem of Cayley. Scripta Mathematica, 22(??): 289–292, ???? 1957. ISSN 0036-9713.

Murnaghan:1944:EPT

[Mur44] F. D. Murnaghan. An elementary presentation of the theory of quaternions. Scripta Mathematica, 10(??):37–49, ???? 1944. ISSN 0036-9713.

Murdoch:1964:RNO

[Mur64] John E. Murdoch. Review of Nicole Oresme’s Quaestiones super geometriam Euclidis edited by H. L. L, Busard. Scripta Mathematica, 27(1964):67–91, 1964. ISSN 0036-9713.

Moser:1963:LPD

[MZ63] L. Moser and W. Zayachkowski. Lattice paths with diagonal steps. Scripta Mathematica, 26(??):223–229, ???? 1963. ISSN 0036-9713.

Nemitz:1973:ISL

[Nem73] William C. Nemitz. Implicative semi-lattices and ternary algebras.

Scripta Mathematica, 29(1–2):39–41, ???? 1973. ISSN 0036-9713.

Neugebauer:1934:BM

[Neu34a] O. Neugebauer. Babylonian mathematics. Scripta Mathematica, 2

(??):312–315, ???? 1934. ISSN 0036-9713.

Neumark:1934:NLC

[Neu34b] S. Neumark. Note on the life of Charles Gill. Scripta Mathematica, 2

(??):139–142, ???? 1934. ISSN 0036-9713.

Neugebauer:1953:HE

[Neu53] O. Neugebauer. On the “Hippopede” of Eudoxus. Scripta Mathemat-

ica, 19(4):225–229 (1954), ???? 1953. ISSN 0036-9713.

Neugebauer:1954:HE

[Neu54] O. Neugebauer. On the “Hippopede” of Eudoxus. Scripta Mathematica, 19(??):225–229, ???? 1954. ISSN 0036-9713. URL https://en.

wikipedia.org/wiki/Hippopede.

Neugebauer:1956:TPT

[Neu56] O. Neugebauer. The transmission of planetary theories in ancient and medieval astronomy. Scripta Mathematica, 22(??):165–192 (1957), ???? 1956. ISSN 0036-9713.

Neugebauer:1957:TPT

[Neu57] O. Neugebauer. The transmission of planetary theories in ancient and medieval astronomy. Scripta Mathematica, 22(??):165–192, ???? 1957. ISSN 0036-9713.

Neugebauer:1959:EEE

[Neu59] O. Neugebauer. The equivalence of eccentric and epicyclic motion according to Apollonius. Scripta Mathematica, 24(??):5–21, ???? 1959. ISSN 0036-9713.

Neville:1938:TM

[Nev38] E. H. Neville. Time and the mathematician. Scripta Mathematica, 5

(??):171–175, ???? 1938. ISSN 0036-9713.

Oehmke:1973:GPA

[Oeh73] Robert H. Oehmke. On the generic polynomial of an algebra. Scripta Mathematica, 29(3–4):331–336, ???? 1973. ISSN 0036-9713.

Osborn:1973:RRJ

[Osb73] J. Marshall Osborn. Representations and radicals of Jordan algebras. Scripta Mathematica, 29(3–4):297–329, ???? 1973. ISSN 0036-9713.

Ostrowski:1963:SIT

[Ost63] A. M. Ostrowski. On some inequalities in the theory of matrices. Scripta Mathematica, 26(??):201–222, ???? 1963. ISSN 0036-9713.

Palama:1953:DST

[Pal53a] G. Palamà. Diophantine systems of the type ∑_{i=1}^{p} a_{i}^{k} = ∑_{i=1}^{p} b_{i}^{k}, (k = 1, 2, …, n, n + 2, n + 4, …, n + 2r). Scripta Mathematica, 19(??):132– 134, ???? 1953. ISSN 0036-9713.

Palama:1953:TDH

[Pal53b] G. Palamà. On a theorem of D. H. Lehmer concerning the Tarry– Escott problem. Scripta Mathematica, 19(??):19–23, ???? 1953. ISSN 0036-9713.

Pauc:1954:LP

[Pau54] C. Y. Pauc. Leibniz in Paris. Scripta Mathematica, 20(??):37–50, ???? 1954. ISSN 0036-9713.

Pinl:1964:MLB

[Pin64] Max Pinl. In memory of Ludwig Berwald. Scripta Mathematica, 27 (??):193–203, ???? 1964. ISSN 0036-9713.

Pinl:1965:MLB

[Pin65] M. Pinl. In memory of Ludwig Berwald. Scripta Mathematica, 27 (??):193–203, ???? 1965. ISSN 0036-9713.

Polachek:1932:MMU

[Pol32] Harry Polachek. A method of multiplication used by Saadia Gaon in the 10th century. Scripta Mathematica, 1(??):245–246, ???? 1932. ISSN 0036-9713.

Polachek:1940:SH

[Pol40] Harry Polachek. The structure of the honeycomb. Scripta Mathematica, 7(??):87–98, ???? 1940. ISSN 0036-9713.

Poulet:1948:NCA

[Pou48] P. Poulet. 43 new couples of amicable numbers. Scripta Mathematica, 14(??):77, ???? 1948. ISSN 0036-9713.

Pollard:1956:ICD

[PS56] Harry Pollard and Charles Standish. Inversion of a class of discrete convolution transforms. Scripta Mathematica, 22(??):207–216 (1957), ???? 1956. ISSN 0036-9713.

Pollard:1957:ICD

[PS57] Harry Pollard and Charles Standish. Inversion of a class of discrete convolution transforms. Scripta Mathematica, 22(??):207–216, ???? 1957. ISSN 0036-9713.

Rajagopal:1951:HPG

[RA51] C. T. Rajagopal and T. V. Vedamurthi Aiyar. On the Hindu proof of Gregory’s series. Scripta Mathematica, 17(??):65–74, ???? 1951. ISSN 0036-9713.

Rajagopal:1952:HAP

[RA52] C. T. Rajagopal and T. V. Vedamurti Aiyar. A Hindu approximation to pi. Scripta Mathematica, 18(??):25–30, ???? 1952. ISSN 0036-9713.

Raine:1948:PTF

[Rai48] Charles W. Raine. Pythagorean triangles from the Fibonacci series 1, 1, 2, 3, 5, 8, . . .. Scripta Mathematica, 14(??):164–??, ???? 1948. ISSN 0036-9713.

Raine:1954:FET

[Rai54] Charles W. Raine. Fibonacci equiareal triangles. Scripta Mathematica, 20(??):96–98, ???? 1954. ISSN 0036-9713.

Rajagopal:1949:NCH

[Raj49] C. T. Rajagopal. A neglected chapter of Hindu mathematics. Scripta Mathematica, 15(??):201–209, ???? 1949. ISSN 0036-9713.

Rao:1955:SSF

[Rao55] K. Subba Rao. Some summation formulae involving Fibonacci numbers. Scripta Mathematica, 21(??):214–217, ???? 1955. ISSN 0036-9713.

Rao:1970:SUD

[Rao70] K. Nageswara Rao. On some unitary divisor functions. Scripta Mathematica, 28(??):347–352, ???? 1970. ISSN 0036-9713.

Rapaport:1959:CCG

[Rap59] Elvira Strasser Rapaport. Cayley color groups and Hamilton lines. Scripta Mathematica, 24(??):51–58, ???? 1959. ISSN 0036-9713.

Reeve:1935:TMS

[Ree35] W. D. Reeve. The teaching of mathematics in the schools. Scripta Mathematica, 3(??):138–142, ???? 1935. ISSN 0036-9713.

Reeve:1938:PMM

[Ree38] William David Reeve. The place of mathematics in modern education. Scripta Mathematica, 5(??):??, 1938. ISSN 0036-9713.

Reeve:1945:ODE

[Ree45] William David Reeve. Obituary: David Eugene Smith. Scripta Mathematica, 11(??):209–212, ???? 1945. ISSN 0036-9713.

Reichman:1954:SFI

[Rei54] Raphael I. Reichman. A summation formula involving Fibonacci numbers. Scripta Mathematica, 20(??):111–112, ???? 1954. ISSN 0036-9713.

Reid:1973:CPP

[Rei73] William T. Reid. A continuity property of principal solutions of linear Hamiltonian differential systems. Scripta Mathematica, 29(3–4):337– 350, ???? 1973. ISSN 0036-9713.

Resnikoff:1943:JCC

[Res43] Louis A. Resnikoff. Jewish calendar calculations. Scripta Mathematica, 9(??):191–195, ???? 1943. ISSN 0036-9713.

Richeson:1934:CHM

[Ric34] A. W. Richeson. Courses on the history of mathematics in the United States. Scripta Mathematica, 2(??):161–165, ???? 1934. ISSN 0036-9713.

Richeson:1939:NUM

[Ric39] A. W. Richeson. Notes on an unpublished manuscript of Valentin Engelhardt. Scripta Mathematica, 6(??):26–31, ???? 1939. ISSN 0036-9713.

Richeson:1941:MS

[Ric41] A. W. Richeson. Mary Somerville. Scripta Mathematica, 8(1):5–13,

March 1941. ISSN 0036-9713.

Richards:1943:NMAa

[Ric43a] John F. C. Richards. A new manuscript of arithmomachia. Scripta Mathematica, 9(??):87–99, ???? 1943. ISSN 0036-9713.

Richards:1943:NMAb

[Ric43b] John F. C. Richards. A new manuscript of arithmomachia. Scripta Mathematica, 9(??):169–183, ???? 1943. ISSN 0036-9713.

Richards:1943:NMAc

[Ric43c] John F. C. Richards. A new manuscript of arithmomachia. Scripta Mathematica, 9(??):256–264, ???? 1943. ISSN 0036-9713.

Richards:1946:BPG

[Ric46] John F. C. Richards. Boissière’s Pythagorean game, translated with notes on the text. Scripta Mathematica, 12(3):177–218, ???? 1946. ISSN 0036-9713.

Rice:1970:RAC

[Ric70] P. M. Rice. Regularity and the annulus conjecture. Scripta Mathematica, 28(??):293–295, ???? 1970. ISSN 0036-9713.

Riordan:1954:DP

[Rio54] John Riordan. Discordant permutations. Scripta Mathematica, 20 (??):14–23, ???? 1954. ISSN 0036-9713.

Roberts:1950:PTT

[Rob50] Harry C. Roberts, Jr. Pythagorean triangles and the Tarry–Escott problem. Scripta Mathematica, 16(??):132–133, ???? 1950. ISSN 0036-9713.

Rolen:2019:YCS

[Rol19] Rolen. Yeshiva College’s Scripta Mathematica encapsulated at the 1939 World s Fair. Yeshiva University News Story, Septem-

ber 24, 2019. URL https://www.yu.edu/library/2019/09/23/
yeshiva-colleges-scripta-mathematica-encapsulated-at-the-
1939-worlds-fair.

Rosen:1959:EMM

[Ros59] Edward Rosen. The editions of Maurolico’s mathematical works. Scripta Mathematica, 24(??):59–76, ???? 1959. ISSN 0036-9713.

Rosenfeld:1967:NTS

[Ros67] Azriel Rosenfeld. A note on two special types of rings. Scripta Mathematica, 28(1a):51–54, ???? 1967. ISSN 0036-9713.

Rosen:1970:JDC

[Ros70] E. Rosen. John Dee and Commandino. Scripta Mathematica, 28(??): 321–326, ???? 1970. ISSN 0036-9713.

Rahman:1966:EPP

[RU66] Q. I. Rahman and K. R. Unni. Extremal problems and polynomials of least deviation. Scripta Mathematica, 27(??):303–329, ???? 1966. ISSN 0036-9713.

Rufus:1941:DRM

[Ruf41] W. Carl Rufus. David Rittenhouse as a mathematical disciple of Newton. Scripta Mathematica, 8(??):228–231, ???? 1941. ISSN 0036-9713.

Rust:1934:OSC

[Rus34] W. M. Rust, Jr. An operational statement of Cantor’s Diagonalverfahren. Scripta Mathematica, 2(??):334–336, ???? 1934. ISSN 0036-9713.

Reich:1960:CCM

[RW60] Edgar Reich and S. E. Warschawski. Canonical conformal maps onto a circular slit annulus. Scripta Mathematica, 25(??):137–146, ???? 1960. ISSN 0036-9713.

Saccheri:1935:LD

[Sac35] Giovanni Saccheri. 1697, Logica demonstrativa. Scripta Mathematica, 3(1):??, ???? 1935. ISSN 0036-9713.

Sandler:1973:HIP

[San73] Reuben Sandler. On homomorphisms and images of projective planes and pseudo planes. Scripta Mathematica, 29(3–4):279–292, ???? 1973. ISSN 0036-9713.

Sawyer:1949:GO

[Saw49] W. W. Sawyer. The game of Oware. Scripta Mathematica, 15(??): 159–161, ???? 1949. ISSN 0036-9713.

Sawyer:1950:WKP

[Saw50] W. W. Sawyer. On a well-known puzzle. Scripta Mathematica, 16 (??):107–110, ???? 1950. ISSN 0036-9713.

Schub:1932:MGB

[Sch32a] P. Schub. A manuscript on geodesy from the Bastille. Scripta Mathematica, 1(??):142–146, ???? 1932. ISSN 0036-9713.

Schwartz:1932:TMM

[Sch32b] Joseph J. Schwartz. Two magical manuscripts. Scripta Mathematica, 1(??):44–52, ???? 1932. ISSN 0036-9713.

Schwartz:1934:MMS

[Sch34] Joseph J. Schwartz. A Menelaus manuscript in the Smith collection, Columbia University. Scripta Mathematica, 2(??):243–246, ???? 1934. ISSN 0036-9713.

Schrek:1950:PRPa

[Sch50a] D. J. E. Schrek. Prince Rupert’s problem and its extension by Pieter Nieuwland. Scripta Mathematica, 16(??):73–80, ???? 1950. ISSN 0036-9713.

Schrek:1950:PRPb

[Sch50b] D. J. E. Schrek. Prince Rupert’s problem and its extension by Pieter Nieuwland. Scripta Mathematica, 16(??):261–267, ???? 1950. ISSN 0036-9713.

Schub:1950:MFC

[Sch50c] P. Schub. A minor Fibonacci curiosity. Scripta Mathematica, 16(3): 214, ???? 1950. ISSN 0036-9713.

Schrek:1955:DBH

[Sch55] D. J. E. Schrek. David Bierens de Haan. Scripta Mathematica, 21 (??):31–41, ???? 1955. ISSN 0036-9713.

Schumaker:1958:PMH

[Sch58] John A. Schumaker. Pierre de Maupertuis and the history of comets. Scripta Mathematica, 23(??):97–108, ???? 1958. ISSN 0036-9713.

Schuster:1959:LDP

[Sch59] Carl N. Schuster. Locating the decimal point. Scripta Mathematica, 24(??):149–153, ???? 1959. ISSN 0036-9713.

Schafer:1973:CTC

[Sch73a] R. D. Schafer. A coordinatization theorem for commutative powerassociative algebras. Scripta Mathematica, 29(3–4):437–442, ???? 1973. ISSN 0036-9713.

Schechter:1973:ESP

[Sch73b] Martin Schechter. On the essential spectrum of partial differential operators bounded from below. Scripta Mathematica, 29(1–2):5–16, ???? 1973. ISSN 0036-9713.

Schechter:1973:OOW

[Sch73c] Martin Schechter. Operators obeying Weyl’s theorem. Scripta Mathematica, 29(1–2):67–75, ???? 1973. ISSN 0036-9713.

Seidenberg:1959:PCA

[Sei59] A. Seidenberg. Peg and cord in ancient Greek geometry. Scripta Mathematica, 24(2):107–122, ???? 1959. ISSN 0036-9713.

Selfridge:1955:RDT

[Sel55] John L. Selfridge. Reviews and descriptions of tables and books. Scripta Mathematica, 21(??):261–268, ???? 1955. ISSN 0036-9713.

Seligman:1973:TDA

[Sel73] George B. Seligman. On two-dimensional algebraic groups. Scripta Mathematica, 29(3–4):453–465, ???? 1973. ISSN 0036-9713.

Shapiro:1958:LTD

[Sha58] Irving Shapiro. The last ten digits of 9^{9^{9^{9}}} . Scripta Mathematica, 23 (??):235–236, ???? 1958. ISSN 0036-9713.

Shoen:1934:ARPa

[Sho34a] Harriet H. Shoen. Archimedes. The reconstruction of a personality. Scripta Mathematica, 2(??):261–264, ???? 1934. ISSN 0036-9713.

Shoen:1934:ARPb

[Sho34b] Harriet H. Shoen. Archimedes. The reconstruction of a personality. Scripta Mathematica, 2(??):342–347, ???? 1934. ISSN 0036-9713.

Sierpinski:1960:SUP

[Sie60] Wacław Sierpiński. On some unsolved problems of arithmetics. Scripta Mathematica, 25(??):125–136, ???? 1960. ISSN 0036-9713.

Sierpinski:1963:PAS

[Sie63] Wacław Sierpiński. Projective and analytic sets. Scripta Mathematica, 26(3):187–195, ???? 1963. ISSN 0036-9713.

Sierpinski:1967:LIT

[Sie67] Wacław Sierpiński. L’induction incomplète dans la théorie des nom-

bres. Scripta Mathematica, 28(??):5–14, ???? 1967. ISSN 0036-9713.

Simons:1932:GAA

[Sim32a] L. G. (Lao Genevra) Simons. A German–American algebra of 1837.

Scripta Mathematica, 1(??):29–36, ???? 1932. ISSN 0036-9713.

Simons:1932:MP

[Sim32b] L. G. (Lao Genevra) Simons. A mathematical publication of 1834.

Scripta Mathematica, 1(??):129–131, ???? 1932. ISSN 0036-9713.

Simons:1932:IGA

[Sim32c] T. E. Simons. Isaac Greenwood’s arithmetic. Scripta Mathematica, 1

(??):262–264, ???? 1932. ISSN 0036-9713.

Simons:1934:BML

[Sim34a] L. G. (Lao Genevra) Simons. Books on mathematics in the library of a New England minister of the seventeenth century. Scripta Mathematica, 2(??):294–295, ???? 1934. ISSN 0036-9713.

Simons:1934:CW

[Sim34b] L. G. (Lao Genevra) Simons. Christopher Wren. Scripta Mathemat-

ica, 2(??):362, ???? 1934. ISSN 0036-9713.

Simons:1934:IGF

[Sim34c] L. G. (Lao Genevra) Simons. Isaac Greenwood, first Hollis professor.

Scripta Mathematica, 2(??):117–124, ???? 1934. ISSN 0036-9713.

Simons:1934:NGA

[Sim34d] L. G. (Lao Genevra) Simons. Note on “A German–American algebra”.

Scripta Mathematica, 2(??):165, ???? 1934. ISSN 0036-9713.

Simons:1935:EEC

[Sim35a] L. G. (Lao Genevra) Simons. An early eighteenth century American readyreckoner. Scripta Mathematica, 3(??):94–96, ???? 1935. ISSN 0036-9713.

Simons:1935:MTS

[Sim35b] L. G. (Lao Genevra) Simons. Mathematical terms in the supplement to the Oxford English dictionary — 1933. Scripta Mathematica, 3 (??):38–43, ???? 1935. ISSN 0036-9713.

Simons:1936:AMF

[Sim36] L. G. (Lao Genevra) Simons. The adoption of the method of fluxions in American schools. Scripta Mathematica, 4(??):207–220, ???? 1936. ISSN 0036-9713.

Simons:1941:UPM

[Sim41] L. G. (Lao Genevra) Simons. Undergraduate publications in mathematics. Scripta Mathematica, 8(??):165–175, ???? 1941. ISSN 0036-9713.

Simons:1945:AAL

[Sim45] L. G. (Lao Genevra) Simons. Among the autograph letters in the David Eugene Smith Collection. Scripta Mathematica, 13(??):??, ???? 1945. ISSN 0036-9713.

Singer:1967:SSF

[Sin67] F. Richard Singer. Some Sheffer functions for M-valued logics. Scripta Mathematica, 28(??):21–27, ???? 1967. ISSN 0036-9713.

Slaught:1933:RPS

[Sla33] H. E. Slaught. Recent publications: Scripta Mathematica. American Mathematical Monthly, 40(4):230–232, April 1933. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Smith:1932:GMP

[Smi32a] David Eugene Smith. Gaspard Monge, politician. Scripta Mathematica, 1(??):111–122, ???? 1932. ISSN 0036-9713.

Smith:1932:MC

[Smi32b] David Eugene Smith. Moritz Cantor. Scripta Mathematica, 1(3): 204–207, ???? 1932. ISSN 0036-9713.

Smith:1932:TJM

[Smi32c] David Eugene Smith. Thomas Jefferson and mathematics. Scripta Mathematica, 1(??):3–14, ???? 1932. ISSN 0036-9713.

Smith:1934:IDR

[Smi34] David Eugene Smith. An interesting document relating to American mathematics. Scripta Mathematica, 2(??):221–223, ???? 1934. ISSN 0036-9713.

Smith:1935:CEM

[Smi35a] David Eugene Smith. Changes in elementary mathematical terms in the last three centuries. Scripta Mathematica, 3(??):291–300, ???? 1935. ISSN 0036-9713.

Smith:1935:EOK

[Smi35b] David Eugene Smith. Euclid, Omar Khayyâm, and Saccheri. Scripta Mathematica, 3(??):5–10, ???? 1935. ISSN 0036-9713.

Smith:1936:AFT

[Smi36] David Eugene Smith. Algebra of four thousand years ago. Scripta Mathematica, 4(??):111–125, ???? 1936. ISSN 0036-9713.

Smith:1938:AYA

[Smi38] David Eugene Smith. Algebra of 4000 years ago. II. Scripta Mathematica, 5(??):15–16, ???? 1938. ISSN 0036-9713.

Smith:1944:SWR

[Smi44] David Eugene Smith. Sir William Rowan Hamilton. Scripta Mathematica, 10(??):9–11, ???? 1944. ISSN 0036-9713.

Spencer:1963:TFE

[Spe63] D. C. Spencer. A type of formal exterior differentiation associated with pseudogroups. Scripta Mathematica, 26(??):101–106, ???? 1963. ISSN 0036-9713.

Schaeffer:1950:MIT

[SS50] A. C. Schaeffer and D. C. Spencer. Models illustrating the third coefficient region for Schlicht functions. Scripta Mathematica, 16(??): 67–71 (2 plates), ???? 1950. ISSN 0036-9713.

Staniland:1938:AM

[Sta38] A. E. Staniland. Art in mathematics. Scripta Mathematica, 5(??): 101–109, ???? 1938. ISSN 0036-9713.

Stancliff:1953:NDD

[Sta53] Fenton Stancliff. Nine-digit determinants. Scripta Mathematica, 19 (??):278, ???? 1953. ISSN 0036-9713.

Steiner:1948:GCR

[Ste48] Jacob Steiner. Geometrical constructions with a ruler, given a fixed circle with its center. Scripta Mathematica, 14:187–264, 1948. ISSN 0036-9713.

Steiner:1950:GCR

[Ste50] Jacob Steiner. Geometrical Constructions with a Ruler (a translation of his 1833 book). Scripta Mathematica, New York, NY, USA, 1950. ???? pp.

Steele:1951:MRC

[Ste51] Domhnall A. Steele. A mathematical reappraisal of the Corpus Platonicum. Scripta Mathematica, 17(??):173–189, ???? 1951. ISSN 0036-9713.

Steinhaus:1963:SB

[Ste63] Hugo Steinhaus. Stefan Banach, 1892–1945. Scripta Mathematica, 26 (??):93–100, ???? 1963. ISSN 0036-9713.

Struik:1949:SLM

[Str49] Dirk J. Struik. A selected list of mathematical books and articles published after 1200 and translated into English. Scripta Mathematica, 15(??):115–131, ???? 1949. ISSN 0036-9713.

Synge:1944:LEW

[Syn44] John Lighton Synge. The life and early work of Sir William Rowan Hamilton. Scripta Mathematica, 10(??):13–24, ???? 1944. ISSN 0036-9713.

Synge:1945:LEW

[Syn45] John Lighton Synge. The life and early work of Sir William Rowan Hamilton: a Collection of Papers in memory of Sir William Rowan Hamilton. Scripta Mathematica, 13(??):??, ???? 1945. ISSN 0036-9713.

Szabo:1964:TMDa

[Sza64a] Árpád Szabó. The transformation of mathematics into deductive science and the beginnings of its foundation on definitions and axioms. I. Scripta Mathematica, 27(??):27–48A, June 1964. ISSN 0036-9713.

Szabo:1964:TMDb

[Sza64b] Árpád Szabó. The transformation of mathematics into deductive science and the beginnings of its foundation on definitions and axioms. II. Scripta Mathematica, 27(??):113–139, August 1964. ISSN 0036-

Szechtman:1960:NDZ

[Sze60] Joshua Szechtman. Notes on Dr. Zuckermann’s introduction to his “Mathematical concepts in the Talmud”. Scripta Mathematica, 25 (??):49–62, ???? 1960. ISSN 0036-9713.

Tabatchnik:1938:SSS

[Tab38] J. Tabatchnik. A spiral scale of square roots. Scripta Mathematica, 5(??):260–262, ???? 1938. ISSN 0036-9713.

Taibleson:1965:MPC

[Tai65] Mitchell H. Taibleson. Mean and pointwise convergence of logarithmic means of Laplace series. Scripta Mathematica, 28(??):197–203, ???? 1965. ISSN 0036-9713.

Taibleson:1968:MPC

[Tai68] M. H. Taibleson. Mean and pointwise convergence of logarithmic means of Laplace series. Scripta Mathematica, 28(??):197–203, ???? 1968. ISSN 0036-9713.

Talmey:1932:PRE

[Tal32] Max Talmey. Personal recollections of Einstein’s boyhood and youth. Scripta Mathematica, 1(??):68–71, September 1932. ISSN 0036-9713.

URL https://en.wikipedia.org/wiki/Max_Talmey;   https://
en.wikipedia.org/wiki/Scripta_Mathematica.

Thebault:1949:RGa

[Thé49a] Victor Thébault. Recreational geometry. Scripta Mathematica, 15 (??):82–88, ???? 1949. ISSN 0036-9713.

Thebault:1949:RGb

[Thé49b] Victor Thébault. Recreational geometry. Scripta Mathematica, 15 (??):149–155, ???? 1949. ISSN 0036-9713.

Thebault:1952:ATF

[Thé52a] Victor Thébault. The area of a triangle as a function of its sides. Scripta Mathematica, 18:151–161, 1952. ISSN 0036-9713.

Thebault:1952:RGA

[Thé52b] Victor Thébault. Recreational geometry: The area of a triangle as a function of its sides. Scripta Mathematica, 18(??):151–161, ???? 1952. ISSN 0036-9713.

Thebault:1953:RGC

[Thé53] Victor Thébault. Recreational geometry: Curvilinear and mixtilinear figures. Scripta Mathematica, 19(??):69–77, ???? 1953. ISSN 0036-9713.

Thebault:1955:RGG

[Thé55] Victor Thébault. Recreational geometry. Geodesics. Scripta Mathe-

matica, 21(??):146–158, ???? 1955. ISSN 0036-9713.

Thebault:1956:RGTa

[Thé56a] Victor Thébault. Recreational geometry: the triangle. I. Scripta

Mathematica, 22(??):14–30, ???? 1956. ISSN 0036-9713.

Thebault:1956:RGTb

[Thé56b] Victor Thébault. Recreational geometry: the triangle. II. Scripta

Mathematica, 22(??):97–105, ???? 1956. ISSN 0036-9713.

Thorndike:1950:GBPa

[Tho50a] Lynn Thorndike. Giovanni Bianchini in Paris manuscripts. Scripta

Mathematica, 16:5–12, 1950. ISSN 0036-9713.

Thorndike:1950:GBPb

[Tho50b] Lynn Thorndike. Giovanni Bianchini in Paris manuscripts. Scripta

Mathematica, 16:169–180, 1950. ISSN 0036-9713.

Thorndike:1953:GBI

[Tho53] Lynn Thorndike. Giovanni Bianchini in Italian manuscripts. Scripta

Mathematica, 19(??):5–17, ???? 1953. ISSN 0036-9713.

Thorndike:1955:GBA

[Tho55] Lynn Thorndike. Giovanni Bianchini’s astronomical instrument.

Scripta Mathematica, 21(??):136–137, ???? 1955. ISSN 0036-9713.

Thorndike:1957:SMA

[Tho57] Lynn Thorndike. The study of mathematics and astronomy in the thirteenth and fourteenth centuries as illustrated by three manuscripts. Scripta Mathematica, 23(??):67–76, ???? 1957. ISSN 0036-9713.

Thorndike:1958:SMA

[Tho58] Lynn Thorndike. The study of mathematics and astronomy in the thirteenth and fourteenth centuries as illustrated by three manuscripts. Scripta Mathematica, 23(??):67–76, ???? 1958. ISSN 0036-9713.

Thompson:1973:WAA

[Tho73a] John G. Thompson. Weighted averages associated to some codes. Scripta Mathematica, 29(3–4):449–452, ???? 1973. ISSN 0036-9713.

Thompson:1973:SVMb

[Tho73b] R. C. Thompson. On the singular values of matrix products. II. Scripta Mathematica, 29(1–2):111–114, ???? 1973. ISSN 0036-9713.

Thomas:19xx:LBN

[Thoxx] J. W. Thomas. On a lower bound to the number of solutions to a nonlinear operator equation. Scripta Mathematica, ??(??):??, ???? 19xx. ISSN 0036-9713.

Touchard:1953:PNP

[Tou53a] Jacques Touchard. On prime numbers and perfect numbers. Scripta Mathematica, 19(??):35–39, ???? 1953. ISSN 0036-9713.

Touchard:1953:PDT

[Tou53b] Jacques Touchard. Permutations discordant with two given permutations. Scripta Mathematica, 19(??):109–119, ???? 1953. ISSN 0036-9713.

Trigg:1955:CGF

[Tri55] Charles W. Trigg. Configuration generated by folding a square. Scripta Mathematica, 21(??):77–80, ???? 1955. ISSN 0036-9713.

ThomasaKempis:1940:WP

[TT40] Sister Mary Thomas a Kempis and Sister Mary Thomas. The walking polyglot. Scripta Mathematica, 6(??):211–217, ???? 1940. ISSN 0036-9713.

Thompson:1973:SVMa

[TT73a] R. C. Thompson and S. Therianos. On the singular values of matrix products. I. Scripta Mathematica, 29(1–2):99–110, ???? 1973. ISSN 0036-9713.

Thompson:1973:SVMc

[TT73b] R. C. Thompson and S. Therianos. On the singular values of matrix products. III. Scripta Mathematica, 29(1–2):115–123, ???? 1973. ISSN 0036-9713.

Tutte:1961:SGC

[Tut61] W. T. Tutte. Symmetrical graphs and coloring problems. Scripta Mathematica, 25(??):305–316, ???? 1961. ISSN 0036-9713.

Uhler:1949:ANF

[Uhl49a] Horace Scudder Uhler. The Arabian Nights’ factorial and the weighted-mean factorial. Scripta Mathematica, 15(??):94–96, ???? 1949. ISSN 0036-9713.

Uhler:1949:TEV

[Uhl49b] Horace Scudder Uhler. Table of exact values of high powers of 2. Scripta Mathematica, 15(??):247–251, ???? 1949. ISSN 0036-9713.

Uhler:1950:MHE

[Uhl50a] Horace Scudder Uhler. Miscellaneous hints for and experiences in computation. Scripta Mathematica, 16(??):31–42, ???? 1950. ISSN 0036-9713.

Uhler:1950:NCP

[Uhl50b] Horace Scudder Uhler. A new constant property of the parabola. Scripta Mathematica, 16(??):161–167, ???? 1950. ISSN 0036-9713.

Uhler:1951:MFV

[Uhl51] Horace Scudder Uhler. Many-figure values of the logarithms of the year of destiny and other constants. Scripta Mathematica, 17(??): 202–208, ???? 1951. ISSN 0036-9713.

Uhler:1952:BHI

[Uhl52a] Horace S. Uhler. A brief history of the investigations on Mersenne numbers and the latest immense primes. Scripta Mathematica, 18 (??):122–131, ???? 1952. ISSN 0036-9713. URL https://t5k.org/

mersenne/LukeMirror/lit/lit_019.htm.

Uhler:1952:MFA

[Uhl52b] Horace S. Uhler. Many-figure approximations for ^{3} \sqrt{2}, ^{3} \sqrt{3}, ^{3} \sqrt{4}, and 3 \sqrt{9} with χ^{2} data. Scripta Mathematica, 18(??):173–176, ???? 1952. ISSN 0036-9713.

Uhler:1952:SNP

[Uhl52c] Horace S. Uhler. Sequel to the note on the parabola. Scripta Mathematica, 18(??):35–38, ???? 1952. ISSN 0036-9713.

Uhler:1953:ELT

[Uhl53a] Horace S. Uhler. Exact location of the 10^{m} th digit in the consecutively written sequence of the natural numbers 1, 2, 3, 4, · · ·. Scripta Mathematica, 19:201–204, 1953. ISSN 0036-9713.

Uhler:1953:ELD

[Uhl53b] Horace S. Uhler. Exact location of the 10^{m}-th digit in the consecutively written sequence of the natural numbers 1, 2, 3, 4, . . .. Scripta Mathematica, 19(??):201–204, ???? 1953. ISSN 0036-9713.

Uhler:1953:MCC

[Uhl53c] Horace S. Uhler. A mathematical copying camera. Scripta Mathematica, 19(??):40–44, ???? 1953. ISSN 0036-9713.

Uhler:1953:PN

[Uhl53d] Horace S. Uhler. On the 16th and 17th perfect numbers. Scripta Mathematica, 19(??):128–131, ???? 1953. ISSN 0036-9713.

Uhler:1954:FVF

[Uhl54] H. S. Uhler. Full values of the first seventeen perfect numbers. Scripta Mathematica, 20(??):240, ???? 1954. ISSN 0036-9713.

Uhler:1955:NEF

[Uhl55] Horace S. Uhler. Nine exact factorials between 449! and 751!. Scripta Mathematica, 21(??):138–145, ???? 1955. ISSN 0036-9713.

Uhler:1956:EVS

[Uhl56] Horace S. Uhler. Exact values of 996!, and 1000!, with skeleton tables of antecedent constants. Scripta Mathematica, 21(??):261–268, ???? 1956. ISSN 0036-9713.

Vandiver:1955:DPN

[Van55] H. S. Vandiver. Divisibility problems in number theory. Scripta Mathematica, 21(??):15–19, ???? 1955. ISSN 0036-9713.

Vandiver:1956:TIR

[Van56] H. S. Vandiver. Is there an infinity of regular primes? Scripta Mathematica, 21(??):306–309, ???? 1956. ISSN 0036-9713.

Vandiver:1961:DAT

[Van61] H. S. Vandiver. On developments in an arithmetic theory of the Bernoulli and allied numbers. Scripta Mathematica, 25(??):273–303, ???? 1961. ISSN 0036-9713.

vanderBlij:1967:CAH

[vdB67] F. van der Blij. Combinatorial aspects of the hexagrams in the Chinese Book of Changes. Scripta Mathematica, 28(1):37–49, ???? 1967. ISSN 0036-9713.

VenkachalamIyer:1954:HA

[Ven54a] R. Venkachalam Iyer. The Hindu abacus. Scripta Mathematica, 20: 58–63, 1954. ISSN 0036-9713.

VenkatachalamIyer:1954:HA

[Ven54b] R. Venkatachalam Iyer. The Hindu abacus. Scripta Mathematica, 20 (??):58–63, ???? 1954. ISSN 0036-9713.

VenkatachalamIyer:1954:MPN

[Ven54c] R. Venkatachalam Iyer. Multigrades with palindromic numbers as elements. Scripta Mathematica, 20(??):220–222, ???? 1954. ISSN 0036-9713.

VenkachatalamIyer:1957:SSC

[Ven57] R. Venkachatalam Iyer. On sums of squares of consecutive numbers. Scripta Mathematica, 22(??):270–273, ???? 1957. ISSN 0036-9713.

VenkatachalamIyer:1958:PE

[Ven58] R. Venkatachalam Iyer. Pell’s equation and u^{2} + v^{2} = w^{2} + 1. Scripta Mathematica, 23(??):233–234, ???? 1958. ISSN 0036-9713.

Vogel:1935:ASS

[Vog35] K. Vogel. Anania of Shirak (Shiragooni). Scripta Mathematica, 3(??): 283–284, ???? 1935. ISSN 0036-9713.

Walker:1934:AM

[Wal34] Helen M. Walker. Abraham de Moivre. Scripta Mathematica, 2(??): 316–333, ???? 1934. ISSN 0036-9713.

Watson:1939:PPA

[Wat39] E. C. Watson. A possible portrait of Arthur Cayley as an undergraduate at Cambridge. Scripta Mathematica, 6(??):32–36, ???? 1939. ISSN 0036-9713.

Wayne:1946:IIO

[Way46] Alan Wayne. Inequalities and inversions of order. Scripta Mathematica, 12(2):164–169, 1946. ISSN 0036-9713.

Wells:1960:MFS

[WC60] Franklin B. Wells and Edward H. Cutler. A modification of a figure suggested by Boyd. Scripta Mathematica, 24(??):339–343, ???? 1960. ISSN 0036-9713.

Weinberg:1935:DBL

[Wei35] J. Weinberg. The disputation between Leonardo of Pisa and John of Palermo. Scripta Mathematica, 3(??):279–281, ???? 1935. ISSN 0036-9713.

Weiner:1955:FTC

[Wei55] L. M. Weiner. A factorization theorem concerning sums of homogeneous functions. Scripta Mathematica, 21(??):20–23, ???? 1955. ISSN 0036-9713.

Weyl:1935:EN

[Wey35] H. Weyl. Emmy Noether. Scripta Mathematica, 3(??):201–220, ???? 1935. ISSN 0036-9713.

White:1932:AMH

[Whi32] M. W. G. White. “Art mindful of him” (a poem). Scripta Mathematica, 1(??):228, ???? 1932. ISSN 0036-9713.

Whitlock:1943:RRTa

[Whi43a] W. P. Whitlock, Jr. Rational right triangles with equal areas. Scripta Mathematica, 9(??):155–161, ???? 1943. ISSN 0036-9713.

Whitlock:1943:RRTb

[Whi43b] W. P. Whitlock, Jr. Rational right triangles with equal areas. Scripta Mathematica, 9(??):265–267, ???? 1943. ISSN 0036-9713.

Whitlock:1946:PV

[Whi46] W. P. Whitlock, Jr. Pythagorean variations. Scripta Mathematica, 12(??):259–265, ???? 1946. ISSN 0036-9713.

Whitlock:1951:DE

[Whi51] W. P. Whitlock, Jr. The Diophantine equation A^{2} + 2B^{2} = C^{2} + D^{2}. Scripta Mathematica, 17(??):84–89, ???? 1951. ISSN 0036-9713.

Whitlock:1953:NPT

[Whi53] W. P. Whitlock, Jr. Nests of Pythagorean triangles. Scripta Mathematica, 19(??):66–68, ???? 1953. ISSN 0036-9713.

Willerding:1960:MTL

[Wil60] Margaret F. Willerding. Mathematics through a looking glass. Scripta Mathematica, 25(??):209–219, ???? 1960. ISSN 0036-9713.

Wintner:1957:SAC

[Win57] Aurel Wintner. On the sine approximations to convex arches. Scripta Mathematica, 23(??):153–156, ???? 1957. ISSN 0036-9713.

Wintner:1958:SAC

[Win58] Aurel Wintner. On the sine approximations to convex arches. Scripta Mathematica, 23(??):153–156, ???? 1958. ISSN 0036-9713.

Wright:1960:SHS

[WW60] E. M. Wright and E. P. Wright. Sir Henry Savile’s lectures on Euclid. Scripta Mathematica, 25(??):63–65, ???? 1960. ISSN 0036-9713.

Wyman:1973:HCF

[Wym73] Bostwick F. Wyman. Hilbert class fields and group extensions. Scripta Mathematica, 29(??):141–149, ???? 1973. ISSN 0036-9713.

Yalden:1932:SEE

[Yal32] J. E. G. Yalden. The sundial as an emblem on early American coins. Scripta Mathematica, 1(??):169–170, ???? 1932. ISSN 0036-9713.

Zeitlin:1973:NCG

[Zei73] David Zeitlin. A new class of generating functions for hypergeometric functions. Scripta Mathematica, 29(1–2):43–48, ???? 1973. ISSN 0036-9713.

Zlot:1960:PCP

[Zlo60] William Leonard Zlot. The principle of choice in pre-axiomatic set theory. Scripta Mathematica, 25(??):105–123, ???? 1960. ISSN 0036-